A Selected Bibliography of Publications by, and about, Sir Rudolf Peierls

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

09 July 2025 Version 1.21

Title word cross-reference

$35.00 [Wal15]. $50.00 [Wal15]. $75.00 [Wal15]. $87.00 [Wal15]. _{2}^{+} [MP74]. 3 [DPS70, DPS76]. ^{4} [DPS76]. _{2}^{+} [MP97]. b [Pei40l]. β [ECPR37, GP58, Pei37a, Pei37b]. γ [BPW54, BPW97].

-Decay [GP58, Pei37b]. -rays [BPW54, BPW97]. -Type [ECPR37, Pei37a].

0 [Pei80c, Pei82b, Pei81c]. 0-06-011108-9 [Pei80c]. 0-14-052341-3 [Pei81c]. 0-333-30511-6 [Pei81c]. 0-7043-2254-4 [Pei81c]. 0-7091-8592-8 [Pei81c]. 0-9594164-0-4 [Pei82b].

1 [FP64, Pei35c]. 100th [Enz00, KRW05]. 1023 [Pei86g]. 11th [AP77]. 12th [AP77]. 17.25 [Pei87b]. 1930s [Pei90c, Stu79]. 1937 [ECPR37]. 1940 [Arn03]. 1942 [KW93]. 1945 [Pei81b, HBP94]. 1947 [Ano47b]. 1958 [Bey59, Pei60b, Uhl59]. 1959 [Wal15]. 1972 [Meh73]. 1976 [MP77]. 1979 [Pei81e]. 1984 [DP86]. 1995 [Lee07b]. 1s [Ano47b].

2 [Ano47b]. 2016 [Sch17]. 20th [Meh73].

3 [Pei81c]. 356 [Pei83a]. 387 [Pei43a].

4 [ECPR37, Pei82b]. 40th [MKR87]. 42 [Pei42h].

56 [Pei42m]. 59 [The12].

6 [Tho99]. 6th [Pei53a].

9 [Pei80c]. 978-0-226-61063-4 [Wal15]. 978-0-230-35487-6 [The12, Wal15]. 978-0-674-06587-1 [Wal15]. 978-981-4307-81-9 [Wal15]. 98 [Pei92b]. 981 [Tho99]. 981-02-2692-6 [Tho99]. 981-02-2876-7 [Tho99].

A-387 [Pei43a]. A. [Pei33a]. AB4 [Pei42h]. AB4/42 [Pei42h]. Abhängigkeit [Pei29a]. Abraham [Pei83d]. Absolute [DPS70]. Absorption [Pei97l, BPS77, FP43, Pei32b, Pei32d]. Absorptionsspektren [Pei32b, Pei32d]. Academic [Pei54a, Pei55a, Pei56a]. Acceleration [PE17]. according [dB32]. Accordion [Pei42a]. Addison [Bey59, Uhl59]. Addison-Wesley [Bey59, Uhl59]. Address [ECPR37, Pei87a]. Admissible [Pei40c]. Adrien [DP86, KW87, Pei85e]. Adsorbed [DPS70, Pei97-27, DSSP76, DPS76, DSSP78, DP82, Pei36d, Pei78b]. Adsorption [Pei97-27, Pei36d]. Advancement [BPM84]. Affairs [Pei69b]. Against [Pei46a]. Age [AM86, BPM84, Pei48c, Pei84a, Stu18, Pei84b, Pei70d]. Agonising [Pei61b]. Air [FKP97, FKP42]. Alamos [HBP94, McC24, MKR87, Pei81b, Pei92b]. Alan [Pei95a]. Albert [Pei83d, Pei55g]. Alice [Pei80f]. alloys [PE17]. Alphen [Sho39]. always [Seg93]. am [Pei50a]. American [ALW^{+}88]. amplitude [Pei40u]. Amsterdam [Pei57b]. analyzer [ZLZ19]. Angular [PMB^{+}58]. anharmonicity [GP78]. Anja [Wal15]. Anniversary [MKR87, OJCP^{+}47, ALW^{+}88]. Answer [Pei34a]. anthology [KK88, KK97]. Antwort [Pei34a]. any [Pei86f]. apparatus [Pei40j, Pei40k, Pei40-30, Pei41v]. Application [DPS70, DFPP43, GP31, PFD43, Pei77b]. Applications [Bro77]. Appreciation [Pei63a, Pei88a]. Approach [PT97, BP68b, Pei42c, PT62]. Approximate [PFD43]. Approximation [PM97, PM80]. approximations [Pei87g]. Arbeit [Pei33a]. arbitrary [PFD43]. Archer [Pei94b]. area [Pei42n, Pei42m]. Arena [Pei69b]. arms [Pei82c, Pei86f]. Arnold [Ano38, Pei51, Pei52a, Pei54a, Pei55a, Pei56a]. Arrangement [Pei40d]. Arrow [Pei94b]. Aspects [SW72, Pei71b]. Assembly [HBP94, Pei41b, Pei41c, Pei41a, Pei42b]. Assessing [AM86]. Assistant [Pei85j, Enz00]. Association [Pei50c]. asymptotic [Pei42c]. Atom [Pei46a, Pei55k, Pei81a, Gol87, Pei92b, Rey03]. atome [Gol87]. Atomic [AM86, BBD^{+}46, BPM84, CP50, FR13, Gol90, Gow64, GA74, Gre20, HBP94, Kel04, PE47, Pei47c, Pei50c, Pei53a, Pei59a, Pei71a, Pei84a, Pei97-45, Tho95, Arn03, Dir36, FP39, Kel07, Kel09, OPM38, Pei29a, Pei35c, Pei47d, Pei47e, Pei84b, Pei87e, Ano47b, Pei70d, Ano47a, Sha50]. Atoms [BPW54, Pei97-27, BPW97, Pei36d]. attached [Pei42h]. attempts [Nav06]. Auflage [Pei51]. August [DP86]. autobiography [Seg93]. automated [ZLZ19]. aux [Per39]. Avery [Pei81c]. Award [Pei62b]. axial [PS49].

B [Pei61a, Pei83a, Pei86g]. Back [Sal02]. background [Pei55e]. backgrounds [Pei50b]. Badash [Pei81b]. ban [Pei66b, Pei70a, Pei89f]. Band [Pei52a, Pei51]. barrier [Pei45]. Barriers [KP37, KP97b, Pei44b]. based [FP40b, FP64, FP65]. Basic [Pei53b]. Basingstoke [The12]. Basis [Pei74c, Pei97o, Pei97-44]. be [Pei91k]. become [ALW^{+}88]. behavior [Pei30c]. Beiträge [KRW05]. Bell [Pei91b]. Bemerkungen [Pei30b, Pei34a, Pei34d]. Berichte [KRW05]. Beta [BHP39, BHP97]. Beta-Disintegration [BHP39, BHP97]. Bethe [Pei57a, Tho99, Bet97, Lee07a, Wal15]. Better [Pei87b]. Between [Pei69b, Pei97-27, Stu18, Pei36d]. Bhabha [Pei86g, Pei86g]. Binding [NTP67]. Binding-energy [NTP67]. Biographies [Wal15]. biography [Pei92g]. Bird [Bas10, Pei85k, Cas86b, Wei86, Bad86, Cas86a, Fel86]. Birmingham [Pei53e, Pei59c]. birth [Kel07, Kel09]. Birthday [KRW05, Enz00]. bismuth [Sho39]. black [FP43, Pei66b]. Blast [BFvN^{+}44, BFH^{+}47, PPvNF44]. bodies [Pei32b, Pei32d, Pei35d]. Body [HP42]. Boer [MP37]. Bohr [BHK^{+}08, FK85, Hal12, Pei40e, Pei55k, Pei63a, Pei85i, Pei88f, Pei97a, Pei97-43, Pei08, Pei10, Whi96]. Boltzmann [Pei94b]. Bomb [Cha92, Gre00, Kel04, Pea24, Pei46a, Pei93b, Pei95b, FP40b, FP40c, FP64, FP65, FP04, Kel07, Kel09, Pei87e, Pei92b, Pei92c, Pei92d, Ree24a, Wel92, Pei92b]. Bombay [Pei86g]. bombs [Arn03, McC25]. Book [Ano47a, Bad86, Bar92, Bel80, Bey59, BPM84, Cas86a, Cas86b, Fel86, Fre56, HBP94, Koe55, Pei47a, Pei51, Pei52a, Pei55a, Pei56a, Pei57a, Pei57b, Pei67b, Pei79a, Pei80e, Pei80f, Pei81b, Pei82b, Pei83a, Pei84a, Pei87b, Pei92b, Pei94a, Pei95a, Rob56, Sal92, Sha50, The12, Tho81, Tho99, Tur56, Uhl59, Wal15, Wei81, Wei86]. Books [Ano47b, Lan57, Pei54a, Pei55a, Pei56a, Pei67b]. Bopp [Pei56a]. Born [Pei71d]. bound [BP68c]. boxes [Pei66b]. Branching [PD44b, Pei40-37, Pei43b, PD44a]. Breaking [Pei84d]. Briefe [KRW05]. Britain [Pei81c, Arn00, Gow64, GA74, HPPR95a, HPPR95b, Pei70d, Pei81c]. British [The12, Wal15, Lau12, Pei50c, Pei52d, Pei85h, Pei86h, Pei92b]. Broida [Pei81b]. Broken [Pei89a, PKA91, Pei92a, Pei97b, Pei91i, Pei91j]. Brum [Gre00]. Build [Pei92b]. Built [Pei55k]. Bulletin [AM86]. bushel [Pei85c].

C [Enz00]. C.B.E. [AP77]. Calcul [Per39]. Calculation [BP67a, BP67b, McC24, Per39]. Calculations [BP67c, NTP67]. California [Pei92b]. Cambridge [Pei79a, Pei92a, Wal15, Hen84, Pei84c, Pei85a, Pei87a]. capillary [Pei40-31]. caricature [Pei92f]. Carnegie [BPM84]. Cascade [Pei41d, Pei42d]. Cascades [PF43b, PF43a, Pei41d, Pei42d]. case [Pei87b, Pei88b, Pei07]. Casimir [Pei83a]. Causes [BPM84]. Cavendish [Nav06]. celebrates [ALW^{+}88]. Celebration [KRW05]. centenary [FK85]. Central [HP42]. Centre [Meh73]. centrifuge [Whi84a, Whi84b]. Century [Pei83a, Meh73, Pei76b, Wal15]. certain [Pei41v]. CF338 [Pei43a]. chaîne [Per39]. Chadwick [Pei94d]. Chain [BPM84, FP40b, FP64, FP65, Pei40f, Pei40g, Pei40h, Pei40i, Pei84a, Pei40-27, Pei40-29, Pei40-33, Pei41f, Per39]. change [Pei31]. changed [Ber11]. Characteristics [DP82, Pei43d]. charge [DP73, DP97c]. Charles [Pei80f]. Chicago [Wal15]. Christoph [The12, Wal15]. city [Pei81b]. claimed [Pei91k]. Clarity [Pei86a]. class [Pei41v]. Classically [GPS86, GPS97, GPS02]. Clerk [Dom63]. Closing [Pei70b, Pei85b]. Clusius [Pei41e]. Clustering [Pei78b]. Cockburn [Pei82b]. Coherent [BPW54, BPW97]. Colaba [Pei86g]. Cold [Pei53b]. Collapse [Pei97-41, Pei85f]. Collected [BRH^{+}99, Pei86g, tH65, BHK^{+}08, Pei08]. Collective [PY57, PY97, PT97, PT62, Pei64f]. College [Tho99]. Collier [Pei87b]. come [Pei91c]. commemorative [Dom63]. Comment [WP83]. Commentary [Pei68b, Bet97, DP97a, Tho99]. Comments [Pea24, Pei43a, POZ^{+}77, Pei78a, Ree22, Pei34a]. Commission [BBD^{+}46]. Communist [Pei95b]. Commutation [Pei52b, Pei52c, Pei97-31]. Company [Bey59, Pei57a, Pei57b, Tho99, Uhl59, Pei67b]. Complex [Pei59b, Pei97-32, Pei97d, Pei48a]. Components [Pei36e, Pei97-28, Pei40-32, Pei41s, Pei41t]. composed [PF43a]. composite [Pei41k, Pei44b]. compound [PF41]. comprehensive [Pei66b]. compression [Pei42g]. compressor [PF43a]. Concentrations [Pei36e, Pei97-28]. Concept [Pei65b]. conception [Meh73]. Conceptual [Pei67b]. Concluding [Pei58a]. conclusions [Pei64d]. Conditions [Pei97c, Pei39b, Pei41f, Pei42c, Per39]. Conduction [APU37, Pei97f, Pei97n, Pei97s, Pei97r, Pei29f, Pei29c, Pei32e, Pei33b, Pei33c]. Conductivity [Pei97p, Pei97-29, Pei30a, Pei30b, Pei34b, Pei64g]. conductors [Pei30c]. conference [Pei53e]. Configuration [Pei97v, LP30]. connection [Pei66b]. conservation [Dir36, Pei58b, Pei90a]. Conservative [Pei86b]. Considerations [Pei40f, Pei40g, Pei40h, Pei40i]. Consolidation [Pei65a]. constant [Pei41f]. Constants [PPG83]. Constituents [Pei53a]. Construction [FP40b, FP64, FP65, Wal15, Arn03]. Continuous [BPP39, BPP97]. Contribution [Pei37a]. Contributions [KRW05]. control [Pei82c, BPM84]. conventional [PE17]. Convergence [AMP97, AMP73]. Cook [Pei95a]. cooking [Pei89b]. Copley [Ano24]. core [ZLZ19]. corps [Pei35d]. Correlations [PMB^{+}58]. correspondence [Lee07a, Lee09, Pei71d]. Cost [Pei42e]. Countdown [Pei81c]. Counting [Pei35e, Pei81c, Pei97z]. Countries [Arm66b]. couple [Pei82d]. Course [Bey59, Uhl59]. courses [Pei52e]. crackling [KK88, KK97]. Crammer [Sha50]. creators [Kel07, Kel09]. crisis [Pei71c]. criteria [AMP73, AMP97]. Critical [Pea24, Pei39b, Pei41f, Pei97c, McC24, McC25, Pei40z, Pei40o, Pei40p, HBP94]. criticism [Pei34a]. cross [Pei40-34]. Crystal [Pei97-34, Pei54b]. crystalline [DSSP76]. Crystals [Pei97n, Pei29c, Pei64g]. cubic [Nab47]. Cul [Pei41h, Pei41i, Pei41q]. Culture [The12, Wal15, Lau12, Pei81d]. Curves [Pei36a, Pei97g].

D [Bey59, BPM84, Kha89, Pei62a, Pei68a, Uhl59, tH65]. D. [PF43b]. Dalitz [Tho99]. Damping [FP47]. Dan [Pei81c]. dans [NBP^{+}34]. d’après [dB32]. Dark [Gre20]. Data [BHP39, BHP97]. David [Pei82b]. day [HHW99]. Days [Pei82a, Pei91a, Pei75, Pei83e, Pei90b, Pei92e]. deadlock [Pei84d]. Decay [GP58, WP83, YBL^{+}58, Pei37b]. Decaying [Pei76a]. Decays [PSP^{+}58]. December [MP77]. deep [Pei87d]. Defence [Pei46a, Pei86i, Pei87b, Pei91d, Pei97-37]. Defense [Pei87b]. Definition [BP49]. Degenerate [Pei97i, Pei36b]. dell’uranio [Pei47a]. Density [DP43, PFD43]. Dependence [PPG83, Pei29a]. dependent [Pei61c, Pei97-30]. derformierbaren [Pei51]. Derivation [Pei97i, Pei36b]. design [Pei40j, Pei40k]. Desorption [DPS70]. detained [Pei85a]. detect [Nav06]. Deterrence [GA74]. Development [BPM84, Meh73, Pei65a, Pei67b, Pei64a, Pei70a, Pei73a, Pei79b, Pei09]. developments [Pei56b]. device [Pei41e, Pei41m, Pei42a]. Diamagnetism [Pei97s, Pei97r, Dat10, Pei32e, Pei33b, Pei33c, Pei53d]. Diamagnetismus [Pei32e, Pei33b, Pei33c]. Dickel [Pei41e]. Did [Pei83d, Pei89d]. Dies [Tho95]. Dieterich’sche [Pei51, Pei52a]. Differentialgleichungen [Pei51]. diffusing [Pei40l]. Diffusion [Arm66a, Pei40j, Pei40k, Pei40r, Pei40-30, Pei42n, Pei44a]. dilemma [Whi96]. Dimensions [KP37, KP97b]. d’incertitude [dB32]. Diplon [BP34c, BP35a]. dipole [BP68a]. Dirac [DP86, Pei92a, KW87, Paĭ90, Pei34c, Pei35a, Pei42h, Pei72, Pei85e, Pei87a, Pei87c, Tay87]. diracsche [Pei35a]. Direct [DSSP76]. Directions [MKR87]. directly [BP67b]. discoveries [Pei94d]. Discovery [BHT86, Pei89e]. Discussion [CP58, ECPR37, GP58, MCS^{+}35, MP37, Pei40j, Pei40k, PMB^{+}58, Pei58c, PSP^{+}58, PPG83, YBL^{+}58, Pei37a, Pei58b]. Disintegration [BP34c, BHP39, BHP97]. Dislocation [Pei97y, Pei40a, ZLZ19]. Dislocations [Nab47, PE17]. Dispersion [KP38, KP97a]. Dispersive [Pei83c, Pei88e, Pei97-39]. distribution [Pei42f, Pei42l, DP43]. Disturbing [Pei80c]. do [Pei63d]. Does [Dir36, HPPR95a, HPPR95b, Pei91c]. double [Pei42g]. double-stage [Pei42g]. Dr. [Pei55c]. Drawing [Ber03]. drink [KK88, KK97]. drop [Pei41g]. due [Pei43b, PD44a]. During [HBP94]. Dyson [Pei80c].

E. [Pei34a]. E.P [Pei81c]. Early [Nav06, Pei80a, Pei82a, Pei91b, Pei93c, Pei83e, Pei90b, Pei92e]. Earth [Pei80e]. Ecology [BPM84, BPM84]. Edited [Ano47b, Pei56a, Pei80f, Pei81b, Pei86g, Pei81c, Sha50]. edition [Pei54a]. Editor [PKA91, Pei94b]. editorial [Pei53b]. education [Pei92f]. Edward [Pei80e, Pei87b]. Effect [Pei40o, Pei40p, Pei40l, PF41, PF43c, Pei40m, Pei40n, Pei40-31, Pei41o, Sho39, Pei29d, Pei97t]. Effective [BP49]. Effects [Pei97q, Pei29e, PF40, Pei40-29]. Effekte [Pei29e]. Effekts [Pei29d]. Efficiency [Pei40r, Pei40q, Pei41g, Pei43b, PD44a, Pei40-37]. Eigenvalues [Pei59b, Pei97d, Pei48a, Pei97-32]. Einführung [dB29]. Einstein [Pei55g, Pei71d, Pei83d, Whi96, Pei83d]. Elected [Ano93, DP86]. Electric [Pei34b, Pei97p]. electrical [Pei30a, Pei32c]. Electrodynamics [Pei97v, LP30]. Electrolytic [Pei40s, Pei46b]. electromigration [DP75, DP97b]. Electron [CP58, Pei34e, PMB^{+}58, Pei58c, Pei97o, DP73, DP97c, Pei34c, Pei32a]. Electrons [BPW54, Pei48b, Pei97s, Pei97r, BPW97, NBP^{+}34, Pei32e, Pei33b, Pei33c, NBP^{+}34]. elektrischen [Pei30a, Pei32c]. Elektrodynamik [Pei51]. Elektronentheorie [Pei34e, Pei32a]. Elemental [Lau12, Wal15, The12]. Elements [PSW52, PSW97, BP67a]. Ellen [BPM84]. Ellyard [Pei82b]. elucidation [Dat10]. Émigré [Wal15]. Emilio [Seg93]. Encounter [Pei88c, Pei93d]. end [Pei91c]. Energy [Ano47a, Ano47b, BBD^{+}46, BPP39, BPP97, CP50, Gow64, GA74, Pei38, Pei42f, PE47, Pei90a, Pei97j, Seg85, BPT67, Dir36, NTP67, Pei47c, Pei47e, Sha50, Pei80e]. England [Tho95]. English [Pei88c, Pei93d]. enjoys [Pei79a]. Enogat [Ano47b, Ano47a]. enough [Pei87d]. enquiry [Pei52e]. Entropy [DPS70, Pei94b, DPS76, PE17]. entry [Pei52e]. Environment [BPM84]. Enz [Enz00]. Equation [FKP97, Pei41j, Pei97i, FKP42, Pei36b]. Equations [EP54, EP97]. Equilibrium [FP42, Pei40t, Pei41k, Pei41l, Pei42c]. Erinnerungen [Pei90b]. Ernests [Oli66a, Oli66b, Oli85]. Ernst [Gre11, Lee07b, Pei60b, Cro68, Dal08]. Error [Pei35e, Pei97z]. Erweiterung [LP31]. Erwiderung [Pei33a]. Essay [Pei65b]. Estimate [Pea24]. ethics [Pei92c]. Evaluation [BPT67]. Evanston [Pei57a]. evaporation [Pei41r]. éventuelles [Per39]. ever [Pei91c]. examination [Pei67b]. Existence [Pei97m, Pei29b]. Existenz [Pei29b]. Expansion [GPS86, GPS97, GPS02, Pei40m]. Expansions [Pei48a, Pei97d]. Expelled [MP01]. Experiences [Ano55, Pei52d]. Experiment [Pei36f]. Experimental [GP58]. Experiments [Pei35e, Pei97z, Pei47b]. Extension [LP65, Pei65a, Pei97e, LP31]. extraordinary [Pei82b]. eyewitnesses [Kel07, Kel09].

F [Bey59, Pei56a, Uhl59]. F. [Arm66b]. F.R.S [AP77]. F.R.S. [DP86, Pei61a]. Fact [Pei82e, Pei86c]. factor [Pei41j]. Factors [CP97a, CP53]. family [Pei80c]. fancy [Pei82e]. Far [Pei83b]. Far-flung [Pei83b]. Farm [McC24]. fast [Pei40n, PF40]. February [MP77, Arn03]. Feld [Pei43a]. Fellows [KK88, KK97]. Ference [Pei92b]. Fermi [BP68b, BP58, BP97e, GP31, Pei72]. Ferromagnetism [Pei97k, Pei36c]. fester [Pei32b, Pei32d]. Festschrift [KRW05]. fiction [Pei84d]. Field [EP54, EP97, Pei52c, Pei53c, Pei54c, Pei57a, Pei57b, Pei63b, Pei52b, PSMF55, Pei55e, Pei55f, Pei73a, Pei97-31]. Fields [Pei33c, Pei57a, Pei57b, Pei30c, Pei97s, Pei57a]. fifth [ALW^{+}88]. film [DSSP78]. Films [DPS70, DSSP76, DP82, Pei78b]. Final [Pei55f]. Finite [Pei40u, Pei40v, Pei40w, BP68b, NTP67]. First [Pei84d, Pei88c, Pei92b, Pei93d, Pei82c, Stu18]. Fisica [Pei47a]. Fission [BHT86, BW39, NBDG40, Pea24, Tur40, Pei42f, Pei86c, Pei89e]. five [ALW^{+}88]. fizike [Paĭ88]. flow [Pei40-34, Pei41j]. fluctuations [Pei44c]. flung [Pei83b]. flushing [Pei42i, Pei42j, Pei42k]. Food [Pei88c, Pei93d, KK88, KK97]. Forbidden [GPS86, GPS97, GPS02]. force [DP73, DP75, DP97b, DP97c, NVP52, Pei68b]. Forces [BP49, HP42, Pei40b, Pei80h, Pei97-30, Wal15, BP67c, BP67d, Pei40-36, Pei61c, Pei79b, Pei81c]. Foreign [Cla23, KK88, KK97]. Foreword [Boh60, Pei86d]. Form [CP97a, EP54, CP53, EP97, dB32]. formalism [PU68, PU97]. forme [dB32]. Formula [KP38, KP97a]. Formulation [Pei64a]. Forty [BPM84, Pei84a, Pei84b, Pei84a]. Foundation [BPM84]. Foundations [BPM84, Pei34e, Pei95a]. four [Bas10]. fourth [Pei54a]. foxes [Pei89b]. fractionating [Pei40-30, Pei41v]. Frage [Pei32c]. Frederic [Pei57a]. Free [Pei38, Pei87b, Pei97j]. Freedman [Pei81c]. Freedom [PD49]. Freeman [Pei80c]. French [Gol87, NBP^{+}34, Pei35d, Per39, dB32]. Frenkel [Pei94c]. Frisch [Pei79a, Pei81e, Arn03, FP40a, FP97a, McC25, Pea24, Pei80d, PR04, Ree22, Ree24b]. front [Pei40l]. frontier [Pei55h]. Frühzeit [Pei90b]. Fuchs [The12, Wal15, Gre20, Lau12, Pei07]. fuel [Pei47d]. Function [Pei97-41, Pei42e, Pei85f]. Functional [EP54, EP97]. Functions [Pei97d, Pei48a]. Fundamental [Pei35b, Pei46c, PPG83, Pei47b]. Further [Sho39].

G [BPM84, Pei55c, Pei83a, Pei84a, Pei94a]. G.m.b.H [Pei52a]. Galvano [Pei97q]. Galvano-magnetic [Pei97q]. galvanomagnetic [Pei29e]. galvanomagnetischen [Pei29e]. Gas [Pei97i, DP73, DP97c, Pei36b, Whi84a, Whi84b]. Gauge [CP54, CP97b]. Gauge-Invariant [CP54, CP97b]. Geburtstag [KRW05]. General [Pei64d, Pei55e]. Generalist [Sal02]. Generalization [Pei77b]. Generation [Wal15]. George [Pei91h]. German [Pei54a, Pei93b, KRW05, LP30, LP31, Pei29a, Pei29b, Pei29f, Pei29c, Pei29e, Pei29d, Pei30a, Pei30c, Pei30b, Pei31, Pei32b, Pei32a, Pei32c, Pei32d, Pei32e, Pei33b, Pei33c, Pei33a, Pei34a, Pei34d, Pei34e, Pei35a, Pei63d, Pei64g, Pei66a, Pei90b, Pei92c, Pei93a, dB29]. Germans [The12, Wal15, Lau12, Pei71a]. get [Pei89d]. Getting [Pei88d]. Gift [MP01]. Gland [PF42, Pei43c, Pei41e]. glands [Pei42i, Pei42j, Pei42k]. glorious [Pei75]. going [Pei89d]. Good [BPM84]. Graduate [Pei59c]. graphite [DPS76]. great [KW87]. growing [Pei80f]. growth [Pei67b]. Grundlagen [Pei34e]. Guest [Pei53b].

H [BPM84, Pei33a, Pei57b, Pei61a, Tho99, MP74, MP97]. H. [Tho99]. Haas [Sho39]. Hackinsack [Wal15]. Hale [Pei81c]. Half [Pei83a]. Hall [McC24, Pei29d, Pei97t]. Hall-Effekts [Pei29d]. Hans [Pei57a, Tho99, Wal15, Bet97]. Haphazard [Pei83a]. hardback [Tho99]. hardcover [Wal15]. Harmondsworth [Ano47b]. Harper [Pei80c, Pei83a].

Harrison [BPM84]. Harvard [Pei80f, Wal15]. hbk [The12]. Heat [APU37, Pei97f, Pei29f, Pei29c]. Heaven [Pei80e]. Heavy [BPW54, BPW97]. Heisenberg [KW93, KRW05, MP77, Cha92, Cla23, KRW05, McC24, Pei71a, Pei92d, Pei92g, Pei93a, Ree24a, Wel92, Pei93b]. Held [Meh73, AP77]. Helium [APU37, Pei97f]. hemisphere [DFPP43, PFD43]. hemispheres [FP43]. hemispheroid [DFPP43, PFD43]. Hendrik [Pei83a]. Herbert [Pei81b]. hexagonal [NVP52]. Hiding [Pei85c]. High [FKP97, FKP42, PE17, Whi84b]. high-entropy [PE17]. high-speed [Whi84b]. Hill [Pei67b]. Hiroshima [OJCP^{+}47]. Hirschfelder [Pei81b]. historians [Kel07, Kel09]. Historical [Sch17, Seg85]. Histories [Pei97-45]. History [Ber03, BPM84, HBP94, Pei93b, WP85, Arn03]. Hitler [MP01]. Hoffmann [Pei57a]. hold [Dir36]. Holders [Ano24]. hole [Pei35a, Pei41m, PF43c]. Holland [Pei57b]. home [Pei85a]. Homi [Pei86g, Pei86g, Pei86g]. Honorary [Ano93]. Humor [Hal12]. Hydrodynamics [BFvN^{+}44, PPvNF44]. hypotheses [Pei87g].

idealised [Pei44c]. ideas [Pei79b]. II [Pei33c, AMP73, AMP97, Oli66b, Pei33c, Pei40x, Pei41h, Pei42j, PFD43, PF43a, Pei43c, Pei65a, Pei77b, Pei97s, Whi84b]. III [Pei41i, Pei42k, Sho39]. Il’ich [Pei94c]. Ill [Pei57a, Ano47b]. illus [Wal15]. im [LP30, Pei35a]. Imperial [Tho99]. including [McC25]. Independence [GA74]. index [Wal15]. India [Pei86g]. Influence [Pei41m, Pei41n]. information [Pei40-35]. Inh [Pei52a]. Initiative [Pei86i]. Innocence [Stu18]. Institute [Pei86g, Hal12]. Interaction [Pei97-27, Pei36d]. Interactions [BP97e, CP54, BP58, CP97b, Pei53c, Pei54c, Pei71d]. Intern [Koe55]. International [Meh73, Pei71b]. Interpretation [BHP39, BHP97, Pei36f, Pei55j, Pei97-33, Pei64a]. Interscience [Pei57b]. Interview [Nai88]. Introduction [Pei58b, Pei77a, Pei78c, Pei86e, dB29]. Introductory [Pei64b]. Invariance [DPP97, DPP42]. Invariant [CP54, CP97b]. Ionisierungsspannung [Pei29a]. ionization [Pei29a]. ions [GP31]. Isaac [Pei67a]. ISBN [Pei80c, Pei81c, Pei82b, The12, Tho99, Wal15]. Ising [Pei36c, Pei97k]. Isobaric [BP68a]. Isotope [Pei41o, Pei40r, Pei40q, Pei47d]. Isotopes [FP97b, NBDG40, FP41, PF41]. Italy [Meh73]. Ivory [Pei69b].

J [Pei56a, Sha50]. J. [Pei55i, Pei74b]. Jacobsen [Wal15]. James [Pei81c, Pei94d]. Jammer [Pei67b]. Jehangir [Pei86g]. Jitterbug [Pei40u, Pei40x]. Jo [Wal15]. John [Ano47a, Ano47b]. Joseph [Pei81b]. Journey [FR13]. Joyce [Pei81c]. July [AP77]. jumps [Pei67b]. June [Lee07b, Pei92a].

Kapitza [Pei85a]. Kapur [Bro77]. Kenntniss [Pei63d]. Kept [Cha92]. Kernmaterie [Pei66a]. Kestin [Pei56a]. key [Arn00, Pei82c]. Kimball [Pei80f]. Kinetic [Pei97n, Pei29f, Pei29c]. kinetischen [Pei29f, Pei29c].

Klaus [The12, Wal15, Gre20, Lau12]. Klemm [Pei52a]. Knopf [Pei93b]. knowledge [Pei63d]. Konfigurationsraum [LP30]. Körper [Pei32b, Pei32d]. Kretschmann [Pei34a]. Kristallen [Pei29c, Pei64g]. Kritik [Pei34a].

L [Bey59, Sha50, Uhl59]. L. [Kha89, Pei62a, tH65]. Laboratory [Nav06]. Lagemann [BPM84]. Landau [Bey59, Uhl59, dB32, Kha89, Pei62a, Pei68a, Pei89c, Pei90c, tH65]. Laporte [Pei55a]. last [Enz00]. Later [BPM84, Pei84a]. lateral [Pei43b, PD44a]. lattice [Nab47]. lattices [NVP52]. Laucht [The12, Wal15]. Law [Pei70c, Pei97-36, NVP52, Pei32c]. Lawrence [Pei81b, Pei81c]. Laws [Fre56, Lan57, Pei52c, Pei55b, Pei59d, Tur56, Pei52b, Pei97-31, Pei63d]. Laxen [BPM84]. layer [Pei40l]. layers [Pei40d]. Lead [BPM84]. leaks [Pei40c]. Lecture [BCC^{+}46, Pei92a, Pei53a]. Lectures [BFvN^{+}44, Pei51, Pei54a, Pei55a, Pei56a, Pei92b, Dom63]. Legacy [Kel04]. Leipzig [KW93]. Leiter [Pei30c]. Leitfähigkeit [Pei30a, Pei30b]. Leitungselektronen [Pei33c, Pei32e, Pei33b]. Léon [Wal15]. Lesson [Pei07]. Letter [Pei46a, PKA91, Pei94b]. Letters [KRW05, Pei80f, Pei92d, Wel92]. Lev [Pei68a]. Lichtgeschwindigkeit [Pei35a]. Life [Arm66b, Pei82b, Pei83d]. Lifshitz [Bey59, Uhl59]. Light [BP73, FP47, FP43, Pei97h, BP97d, Pei35a, Pei76c, Pei77b, Pei85d, Pei88e, Pei91a, Pei97-40]. lights [Pei85c]. Limited [Pei82f]. lines [BPS77]. Liquid [APU37, Pei97f]. Little [Pei79a]. Lives [Gre20]. Living [Pei48c]. Local [CP54, CP97a, PM80, PM97, CP53, CP97b, Pei53c, Pei54c]. Löchertheorie [Pei35a]. logic [Pei86f]. London [Bey59, Pei54a, Pei55a, Pei56a, Pei87b, Uhl59, Wal15]. Longitudinal [Pei44a]. Looks [Sal02]. Lord [Pei83d]. Lorentz [DPP42, DPP97, GPS86, GPS97, GPS02, Pei62b]. loss [Pei43b, PD44a]. Low [MCS^{+}35, GP78, Pei32c, Pei34b, Pei41r]. low-temperature [GP78]. Ltd [Pei54a, Pei55a, Pei56a].

M [BGM91, Bey59, BPM84, Pei86g, Uhl59]. MA [Wal15]. machine [PF43d, Pei81c]. Macmillan [Pei87b, Pei92b, The12, Wal15]. Macmillan/St [Pei92b]. Macrophysical [PPG83]. Made [Pei92h]. Magnetfelder [Pei33c]. Magnetfeldern [Pei30c]. Magnetic [Pei33c, Pei36a, Pei46b, Pei97g, Pei30c, Pei31, PS49, Pei97s, Pei97q, Sho39]. magnetischen [Pei31]. mainly [Pei93c]. Makes [Pei94a]. Making [Kel04, Pei80g, Pei87e, Pei97-38, The12, Wal15, Lau12, Pei92b]. man [Kha89]. Manhattan [Pei92b, HHW99, Kel04, Kel07, Kel09]. March [ECPR37]. Mark [Pei82b, AP77]. Martin [Pei54a, Pei92b]. Marvin [BGM91]. Mary [Wal15]. Mass [Pea24, McC24]. Mathematical [Pei59c]. mathematicians [Pei82d]. Mathematics [Ber03, Pei52e]. matrix [BP67a]. matter [Pei55i, Pei63c, Pei66a]. Maurice [DP86, KW87, Pei85e]. Max [Pei67b]. Maxwell [Dom63, Pei63b]. McGraw [Pei67b]. McGraw-Hill [Pei67b]. Me [Ber09]. means [Pei41o, Pei41r]. Measurable [WP83]. Measurement [CP58, PMB^{+}58, Pei58c, DSSP76, Pei91d, Pei97-37]. measurements [Sho39]. Mechanics [Pei97-41, Pei67b, Pei85f, Pei90b, Pei92e, Pei84e, dB29, Pei56a, Pei67b]. Mechanics. [Pei54a]. Mechanik [Pei51, Pei51]. Mechanism [BW39]. Medal [Ano24, Pei62b]. Medien [Pei51]. Medium [BP73, Pei83c, BP97d, Pei76c, Pei77b, Pei88e, Pei97-40, Pei97-39]. Meeting [Pei87a]. Meixner [Pei56a]. Mellanby [BPM84]. Members [Ano93, KK88, KK97]. Membership [Cla23]. Membrane [Pei40y, Pei43d, Pei40l, Pei40v, Pei40w, Pei40-28, Pei41g, Pei41n, Pei41u, Pei42e, Pei42n, Pei42m, Pei43b, PD44a, Pei44a]. membranes [Pei40d]. Memo [BBD^{+}46]. Memorandum [FP40a, FP97a, FP04, Ree22, Ree24b, Arn03, Ber11, McC25]. Memorial [Pei87a, Pei92a]. Memories [Pei81b]. men [Pei81a]. Menual [Pei81c]. Meson [Pei57a, Pei39a]. Mesons [Pei57a, Pei57a]. mesotron [Pei39c]. Metalle [Pei33a, Pei34a, Pei32a, Pei34e]. Metallen [Pei30a]. metallic [Pei30c]. metallischer [Pei30c]. Metals [Pei34e, Pei97w, Pei97o, Pei97p, NBP^{+}34, Pei30a, Pei32a, Pei33a, Pei34a, Pei53d]. métaux [NBP^{+}34]. method [Pei40s, Pei41h, Pei41i, Pei41q, BPW54, BPW97]. MeV [PP47]. Mexico [Pei56c]. Michael [Wal15]. Microwave [Pei89b]. Middlesex [Ano47b]. Miller [BGM91]. mind [Seg93]. Minimum [Pei38, Pei97j]. Miramare [Meh73]. misappraisal [Pei61b]. Misconceptions [Pei80b]. Mixing [DPS76]. mixture [Pei40-32]. MM. [dB32]. Model [BP97e, MP97, PY57, Pei80g, PY97, Pei97k, Pei97-38, AMP73, AMP97, BP58, GP78, GP31, MP74, Pei36c, PS49]. Model-Making [Pei80g, Pei97-38]. modeling [PE17]. Models [Pei87g]. Modern [POZ^{+}77, Pei78a, Tho99, Dom63, Pei48b]. Moldauer [Pei55a]. molecular [Pei41m]. molecules [PF41]. Momentum [BP73, Pei83c, Pei85d, Pei88e, Pei91a, Pei97h, BP97d, Pei76c, Pei77b, Pei97-40, Pei97-39]. Monogr [Koe55]. monolayers [DPS76]. Mordechai [Pei88b]. Morton [Pei92b]. Motion [PY57, PY97, PT97, PT62, Seg93]. moving [DP73, DP97c]. MS. [Pei42m]. Multiplication [Pei97c, Pei39b, PFD43]. Muon [YBL^{+}58]. My [Bha10, Pei89c]. Mystery [FR13].

Nabarro [Pei68b, PE17, ZLZ19]. National [BPM84]. Nature [Pei55b, Pei59d, Pei94a, Meh73, Fre56, Lan57, Tur56]. Naturgesetze [Pei63d]. Nazi [MP01]. Nazis [Cha92]. near [DP43]. necessary [Pei91k]. need [HPPR95a, HPPR95b]. needed [Pei47b]. net [Rey03]. neubearbeitete [Pei51]. Neutrino [BP34a, BP34b, BP97a, BP97b, PMB^{+}58, Pei82a, Pei83e, Nav06]. Neutron [Pei97c, Pei39b, Pei40-27, PFD43]. Neutrons [BP35b, BP97c, Pei40n, PF40, Pei40-33, Pei42f]. Never [Pei93b]. News [Ano47b]. Newton [Pei94a]. nichtmetallischen [Pei64g]. Niels [BHK^{+}08, FK85, Pei63a, Pei08]. Nobel [Pei62a, Sch17]. Non [CP54, CP97a, HP42, PM97, CP53, CP97b, Pei43b, PF43c, Pei53c, Pei54c, Pei58b, PM80]. Non-Central [HP42]. non-conservation [Pei58b]. Non-Local [CP54, CP97a, PM97, CP53, CP97b, Pei53c, Pei54c, PM80]. non-rectangular [Pei43b]. non-uniform [PF43c]. nonmetallic [Pei64g]. nonrectangular [PD44a]. North [Pei57b]. North-Holland [Pei57b]. Note [DSSP78, Pei36b, Pei42g, Pei54b, Pei97i, Pei97-34, Sch17]. Notes [Pei42h, Pei44b]. November [Pei80c]. Nuclear [AM86, BHT86, BP49, BGM91, BCC^{+}46, BW39, BPP39, BPR86, BPP97, BP67d, BPM84, ECPR37, FP40b, FP64, FP65, HP42, KP38, NBDG40, Pei40e, Pei40b, Pei48c, Pei48d, PY57, Pei63c, Pei80c, Pei81c, Pei84a, Pei89d, Pei91f, Pei91g, Pei97a, PY97, Pei97-35, Pei97-30, Pei08, Seg85, Stu18, The12, Tur40, Wal15, Arn00, Arn03, BP67a, BP67b, HPPR95a, HPPR95b, KP97a, Lau12, Pei37a, Pei40-36, Pei56b, Pei59e, Pei61c, PU68, Pei79b, Pei82f, Pei84d, Pei86f, Pei91k, PU97, Pei66a, Stu79, Wal15]. nucleare [Pei47a]. nuclei [BP68b, NTP67, Pei35c]. nucleon [BP67b, BP68c]. Nucleus [FR13, Pei53a, Pei59a]. Number [PE47, Pei29a, Ano47b, Ano47a]. Nye [Wal15].

O [Pei54a, Pei81b]. O. [Pei80d]. Obituary [Cat95, DP86, Pei55c, Pei67a, Pei80d, Pei85e, Pei91h]. oblate [DFPP43, PFD43]. oblique [Pei77b]. Observational [Pei95a]. Observations [Pei85f, Pei97-41]. occasion [AP77]. October [DP86, Pei81c, Pei81e]. Odd [Pei82d, Paĭ90]. Oliphant [Pei82b, Pei82b]. One [Pei81d]. Open [Pei60c]. Opening [ECPR37]. Oppenheimer [HBP94, Pei80f, Pei55i, Pei74b, Pei80f]. Optics [Pei55a]. Optik [Pei52a]. Optimization [PE17]. Optimum [Pei41p]. order [BPT67]. ordinary [Pei40n, PF40]. Ordnungszahl [Pei29a]. Origin [PSW52, PSW97]. Origins [Wal15]. Orion [Pei80c]. Other [MCS^{+}35]. Otto [Pei55a, Pei79a, PR04, Pei81e]. our [Pei63d, Pei79b, Pei91a]. Output [Pei44c]. Oxford [AP77].

P [BPM84, Enz00, Pei81b]. Packing [Pei40-28, Pei40o, Pei40p]. pains [Pei80f]. pair [Pei40v, Pei41g, Pei42e, Pei43b, Pei43d, PD44a]. pairs [Pei40y, Pei40-28, Pei41u, Pei44a]. Pais [Pei83d]. Palgrave [The12, Wal15]. paper [Ano47b, MP37, Pei42h]. Papermac [Pei81c]. papers [DP97a, Tho99, tH65, Pei86g]. Paradox [Pei35b]. Parity [PSP^{+}58, Pei58b]. Part [FP40a, FP65, Pei40x, FP64, Pei35c, PFD43, PF43a, Pei43c, Pei64a, Pei65a, Whi84a, Whi84b]. participant [Pei85h, Pei86h]. Particle [PSP^{+}58, PE17]. Particles [Pei46c, Pei80h, Pei55f]. Partielle [Pei51]. Passage [Bad86, Cas86a, Cas86b, Pei85k, Wei86, Bas10, Pei40-31, Fel86]. Paul [DP86, KW87, Pei85e, Tay87]. Pauli [Pei60b, Enz00, Pei85j, Pei92h]. Peierls [AP77, Ano47a, Ano47b, Bel80, Bey59, Cas86b, Fre56, Koe55, Lee07b, Rob56, Sal92, Sha50, The12, Tho99, Tur56, Uhl59, Wal15, Wei86, Arn03, Bad86, Bar92, BGM91, Bha10, Bro77, Cas86a, Cat95, Cro68, DP97a, Dal08, Dat10, Fel86, FP40a, FP97a, Gre11, Kni95, Lan57, Lau12, Lee07a, Lee09, McC25, Nai88, Pea24, PE17, Pei68b, Pei97u, Ree22, Ree24b, Tho95, Tho81, Wei81, ZLZ19, dB32]. Penetration [KP37, KP97b]. Penguin [Ano47b, Pei81c]. Penney [Pei91h]. People [Pei92e]. performance [Pei41j]. Pergamon [Bey59, Uhl59]. personal [Enz00]. Perspective [Seg85, POZ^{+}77, Pei78a]. Perspectives [Kel04, Pei87b]. Perturbation [AMP73, AMP97, MP97, Pei73b, Pei97-42, MP74]. perturbed [Pei50a]. Peter [Pei55a]. Peterson [Pei57a]. phase [BP67b]. Phenomena [MCS^{+}35, PPG83, FP39, OPM38]. phenomenological [BP67c]. Philanthropic [BPM84]. Philosophy [Wal15]. Photoelectric [BP34c]. Physicist [Pei82b, Pei85k, Tho95, Wal15, Kha89, KW87, Meh73, Pei64e, Pei79a, Pei83b, PR04, Bad86, Cas86a, Cas86b, Wei86]. Physicists [Pei69a, Pei83a]. Physics [Bey59, BCC^{+}46, BPR86, Hal12, Hen84, LL58, Meh73, MKR87, Pei48d, Pei51, Pei54a, Pei55a, Pei56a, Pei57b, Pei59c, Pei62a, Pei64b, Pei79c, Pei80g, Pei80a, Pei82a, Pei83e, Pei86g, Pei91e, Pei91a, Pei93e, Pei97-38, Ree22, Sch17, Stu18, Tho99, Uhl59, Wal15, WP85, AP77, Paĭ88, Pei47b, Pei48b, Pei55h, Pei75, Pei76b, POZ^{+}77, Pei78a, Pei82e, Pei91c, Pei08, Stu79, Whi84a, Hal12, Koe55, Pei56c, Pei63d, Bar92, Bel80, Bey59, Pei95a, Sal92, Tho81, Uhl59, Wei81]. Physik [Pei52a, Pei51, Pei51]. Pion [YBL^{+}58]. Pioneers [Tho99, Gol87]. Pionniers [Gol87]. pipe [Pei40-34]. Placzek [Pei55c]. Planning [BPM84]. plant [FP42, Pei40c, Pei40t, Pei41b, Pei41c, Pei41a, Pei41d, Pei41j, Pei41k, Pei41l, Pei41p, Pei41s, Pei41t, Pei42b, Pei42e, Pei42l, Pei42n, Pei44c]. plates [Ano47b, Pei51]. Playful [Hal12]. plus [dB32]. PNADIS [ZLZ19]. poem [Pei85g]. point [DP43]. Polanyi [Wal15]. Polarization [CP58, Pei58c]. Politics [Wal15, Pei69a]. Pollution [BPM84, Pei71b]. Polyneutron [PSW52, PSW97]. Pomeranchuk [Pei67a]. porosity [Pei42e, Pei45]. positive [GP31, Pei34c]. Positron [Pei65b]. possibility [Pei40-27, Pei82c]. possible [Arn03, Pei66b, Per39]. Potential [KP37, KP97b, PM97, PM80]. Power [BPM84]. Powers [Pei93b]. Pp [Bel80, Bey59, Pei51, Pei52a, Pei54a, Pei55a, Pei56a, Pei57a, Pei57b, Pei80f, Pei82b, Pei83a, Pei86g, Pei92b, Tho99, Uhl59, Ano47b, BPM84, Pei79a, Pei80c, Pei81b, Pei81c, Pei84a, Pei87b, The12, Wal15]. Practice [Pei57a]. Praemia [Pei62a]. Predictions [McC25]. present [Pei37b]. Press [Bel80, Bey59, Pei54a, Pei55a, Pei56a, Pei79a, Pei80f, Pei87b, Pei92b, Tho99, Uhl59, Wal15]. Pressure [DPS70, Pei42l, DSSP78, Pei41g, Pei42e, Pei42n, Pei42m, Pei88e]. pressures [DSSP76, Pei41r]. Primer [Pei92b, McC24]. Princeton [Bel80]. Principle [LP31, LP65, Pei97e]. Principles [Cha92, Whi84a, Whi84b]. private [Lee09, BPM84]. Prize [Pei62a]. Problem [HP42, Pei40-35, Pei71b, Pei82c]. Problems [Pei46b, Pei48d, Pei64f, Pei41b, Pei41c, PF42, Pei43c, PS49, Pei60c]. proceedings [AP77, Stu79]. processes [Dir36]. Production [Pei47d]. Prof [Pei51, Pei52a, Pei57a, Pei57b]. Prof. [Enz00]. Professor [AP77, Ano38, Pei61a, Pei88a]. Project [Arm66a, Pei92b, HHW99, Kel04, Kel07, Kel09, Pei93b]. Projected [MP97, Pei73b, AMP73, AMP97, MP74, Pei97-42]. promise [Pei47c]. Propagation [FP47]. propagators [Pei55j, Pei97-33]. Properties [CP53, CP97a, FP40c, FP04, Pei35d, Pei41n, Pei55j, Pei97-33, Sho39]. Property [Pei38, Pei97j]. Prophet [Arm66b]. Proposal [Pei91g]. propriétés [Pei35d]. Protection [BPM84]. Protest [Pei81c]. Proton [PP47, WP83]. Protons [BP35b, BP97c]. Prowling [Pei84e]. Pseudomomentum [Pei97h, Pei85d, Pei91a]. Public [BPM84, Pei69b]. Publishers [Pei57b]. Publishing [Bey59, Pei57b, Tho99, Uhl59]. Pugwash [PP68]. Pulse [Pei83c, Pei97-39].

Quantenelektrodynamik [LP30]. Quantenmechanik [Pei90b]. Quantentheorie [LP31]. Quantum [BP35a, DPP97, Hal12, LP65, Pei35b, Pei55d, Pei57b, Pei60a, Pei64c, Pei65a, Pei65c, Pei74a, Pei82e, Pei96, Pei97e, Pei97-41, Pei97v, Pei01, DPP42, LP31, Pei35c, Pei64a, Pei67b, Pei73a, Pei85f, Pei90b, Pei92e, Pei09, Pei84e, SW72, Whi96, Koe55, LP30, Pei35c, Rob56, Pei57b]. Quartet [Pei81c]. Quelques [Pei35d]. question [Pei32c].

R [Ano47b, Bel80, Bey59, BPM84, Lan57, Pei79a, Pei80d, Sha50, Tho99, Uhl59]. R. [Fre56, Koe55, Rob56, Tur56]. rabbit [PF43d]. race [Pei86f]. Radiation [FP47]. radii [McC25]. Radioactive [FP40c, FP04, Pei40-29]. Radioactivity [FP39, OPM38]. radius [Pei40p]. Range [BP49]. rate [PFD43]. Ratio [PD44b, Pei40-37, Pei43b, PD44a]. ray [BPS77]. rays [BPW54, BPW97]. Reaction [BHT86, BPM84, FP40b, FP64, FP65, Pei84a, BP67a, Pei40-27, Pei40-29, Pei40-33, Pei41f]. Reactions [BPP39, BPP97, KP38, Pei40e, Pei40f, Pei40g, Pei40h, Pei40i, Pei97a, KP97a]. reactor [McC24]. reactors [McC25]. Reading [Bey59, Uhl59]. realistic [BP67d, BP68b, Pei82c]. Reality [Pei83a]. reasons [ALW^{+}88]. Recalling [Arn00]. Recollections [Bad86, Cas86a, Cas86b, Pei80a, Pei85k, Pei94d, Wei86, Kha89, Pei71a, Pei85i, Pei89c, Pei80f]. recovery [Pei41c]. rectangular [Pei40-34, Pei43b]. Reflections [Pei85h, Pei86h, Pei89e, PKA91, Pei91i]. reflexion [Pei77b]. Refracting [BP73, BP97d, Pei76c, Pei77b, Pei97-40]. Refugee [Seg85]. Regarding [Pei40f, Pei40g, Pei40h, Pei40i]. Regime [MP01]. Region [BPP39, BPP97]. Reidel [Pei81b]. relatif [Per39]. relations [dB32]. Relativistic [LP65, Pei52c, Pei97e, Pei97i, LP31, Pei36b, Pei52b, Pei97-31]. relativistische [LP31]. release [Pei47e]. Remark [Pei92h]. Remarks [Pei58a, Pei64b, Pei74c, Pei97w, Pei97x, Pei97-44, Pei97-29, Pei30b, Pei55f, Pei34d]. Remember [Pei79a]. Remembering [Kel04]. Reminiscences [KW87, Pei84c, Pei94d, Pei90b, Pei81b]. Reply [BGM91, Pei87f, Pei88d]. Report [Pei53e, Pei43a]. Reports [KRW05]. resistance [Pei31, Pei32c, ZLZ19]. Resonance [Bro77]. Resonant [GCP76]. Response [Pei62b, Pei33a]. restrictive [dB32]. results [Pei41d]. retirement [AP77]. retrospect [Stu79]. reversal [Pei70c, Pei97-36]. Review [Ano47a, Bad86, Bar92, Bel80, Bey59, BPM84, Cas86a, Cas86b, Fel86, Fre56, Pei47a, Pei51, Pei52a, Pei54a, Pei55a, Pei56a, Pei57a, Pei57b, Pei65b, Pei80e, Pei80f, Pei81b, Pei82b, Pei83a, Pei94a, Pei95a, Rob56, Sha50, The12, Tho81, Uhl59, Wal15, Wei81, Wei86, Whi84a, Whi84b, Pei79a, Pei87b, Pei93b, Sal92, Tho99]. Reviews [HBP94, Koe55, Pei92b, Tur56, BPM84, Pei84a]. Revisiting [Ree24b]. revolutionary [Pei86b]. Rivals [Gol90]. Road [Pei86g]. Robert [BPM84, Pei55i, Pei74b, Pei81e, Pei84a, Pei92b, Pei80f, PR04, Pei80f]. Roberts [BPM84]. Roger [Pei94a]. role [Arn00]. Rosenfeld [Wal15]. rotation [PU68, PU97, Whi84b]. Row [Pei57a, Pei80c, Pei83a]. Royal [KK88, KK97, Cla23]. Rudolf [AP77, Ano47a, Bad86, Bar92, Bha10, Cas86a, Cat95, Cro68, Dal08, Fel86, Gre11, Lee09, The12, Tho81, Tho99, Wal15, Wei81, AP77, BGM91, Cas86b, Kni95, Lau12, Lee07b, Nai88, Sal92, Tho95, Wei86]. Russian [Bey59, Uhl59, Paĭ88, Pei85a]. Rutherford [Pei53a, Pei88f, Pei97-43, Pei10].

S [Wal15]. S. [PF43b]. sac} [Pei41q, Pei41h, Pei41i]. Sachs [BPM84, Pei84a]. scale [Pei41u]. Scattering [BP35b, BP97c, BPW54, PP47, Pei59b, BPW97, Pei97-32]. School [Pei56c, Pei59c]. Schrödinger [Cla23]. Schweber [Pei57a, Wal15]. Science [HBP94, PD49, Pei53b, Pei83d, Dom63, Pei71c, Wal15, Pei83a, Ano47b]. Scientific [BPM84, Pei86g, Tho99, Wal15, DP97a, Lee09, Pei50b]. Scientist [Pei50c, Pei69b, ALW^{+}88, Pei92g, ALW^{+}88]. Scientists [MP01, Pei92b, Seg85, AM86]. scissione [Pei47a]. Second [OJCP^{+}47, BPT67, Pei42m, Pei70c, Pei97-36, Stu18]. second-order [BPT67]. Secret [Pei93b, Pei81b]. section [Pei40-34]. security [Pei64f, Rey03]. Segrè [Seg93]. seinem [KRW05]. Selected [Bet97, DP97a, Pei59e, Pei97-35, Lee09, Tho99]. Selections [AM86, POZ^{+}77, Pei78a]. self [Pei40-30, Pei41v]. self-fractionating [Pei40-30, Pei41v]. Separated [NBDG40]. Separating [Pei40-31]. Separation [FP41, FP97b, Pei40-32, Pei41q, Pei41r, FP42, Pei40r, Pei40q, Pei40t, Pei41a, PF41, Pei41l, Pei41o, Pei41p, Pei41s, Pei41t, Pei41u, Pei42b, Pei42e, PF43c, Pei47d, Whi84a]. September [Lee07b, Meh73, Pei81c, Pei81e]. Ser [Koe55]. Serber [McC24]. Series [BCC^{+}46, Pei57b]. Sets [Pei97d, Pei48a]. Seventy [ALW^{+}88]. seventy-fifth [ALW^{+}88]. Seventy-five [ALW^{+}88]. Several [KP37, KP97b]. Shankland [Pei36f]. Shape [GPS86, GPS97, GPS02, PFD43]. Shield [Pei87b]. shift [BPT67]. shifts [BP67b]. Shock [BFvN^{+}44, PPvNF44]. Short [Pei54a]. Side [Hal12]. Silvan [Pei57a, Wal15]. Simon [Arm66b, Pei40c, Pei41b, Pei41d, Pei41j, Pei41k, Pei41s, Pei41t]. Simple [Pei74c, Pei97-44, Nab47, NVP52]. Simplified [MP97, MP74]. since [Pei63b]. Singh [Pei86g]. single [Pei40w, Pei42g]. single-stage [Pei42g]. Sir [AP77, Dal08, Pei82b, Tho99, Bha10, Cat95, Gre11, Lee09]. Situation [GP58]. six [Dom63]. Sixty [FR13]. Size [Pei97y, Pei40z, Pei40o, Pei40a, Pei41c, Pei41p, PF43c]. sizes [PF43a]. Skaar [Wal15]. Skinner [Pei61a]. Skyrme [Pei88a]. Slightly [Pei83c, Pei97-39]. slip [ZLZ19]. slipped [Rey03]. slow [Pei40-27, Pei40-33]. slow-neutron [Pei40-27]. Small [Pei41u, Pei41g]. Smith [Pei80f, Pei81c]. Social [Wal15]. Society [BPM84, Cla23, KK88, KK97]. Solid [Pei64b, Pei80a, Pei32b, Pei32d, Pei35d, PFD43]. Solid-State [Pei64b]. solides [Pei35d]. Solids [Pei55d, Pei64c, Pei65c, Pei74a, Pei96, Pei97l, Pei01, Pei60a, Pei93c, Koe55, Rob56]. Solitons [GPS86, GPS97, GPS02]. soluble [AMP73, AMP97]. Some [Ano55, Bro77, Paĭ90, Pei52d, Pei74c, Pei85i, Pei97-44, Pei35d]. Sommerfeld [Pei51, Pei52a, Pei54a, Pei55a, Pei56a, Ano38]. Sound [Pei83c, Pei97h, Pei85d, Pei91a, Pei97-39]. source [DP43]. Space [Pei97v, LP30]. spacing [Pei41m]. specific [Pei40j, Pei40k]. Spectra [Pei97l, Pei32b, Pei32d]. Spectrum [Pei97-34, Pei54b]. speculation [Pei86c]. Speed [Pei40-33, Pei35a, Whi84b]. Sperber [Pei92b]. sphere [Pei40z, Pei40m, Pei40p]. spheroid [DFPP43, PFD43]. spin [BP68a, Pei88d]. Spontaneously [Pei91j]. spreading [DSSP76, DSSP78]. Spy [Gre20, Rey03]. squabbles [Pei82e]. Sreekantan [Pei86g]. St [Pei92b]. stability [NVP52]. Stage [BP97e, BP58, Pei42g]. stand [Pei63d]. Starke [Pei33c]. starken [Pei30c]. State [FKP97, Pei64b, Pei80a, Pei97i, BP68c, FKP42, Pei36b, Pei37b]. Statements [OJCP^{+}47]. States [BPM84, GPS86, GPS97, GPS02, Pei73b, Pei97m, AMP73, AMP97, BP68a, GCP76, Pei29b, Pei76a, Pei97-42]. stationärer [Pei29b]. Stationary [Pei97m, Pei29b, Pei40l]. Statistical [LL58, Pei34e, Pei35e, Pei36d, Pei36e, Pei56a, Pei97o, Pei97z, Pei97-27, Pei97-28, Bey59, Uhl59]. statistics [Pei72]. statistischen [Pei34e]. stehen [Pei63d]. Stern [Pei54a]. Stewart [Pei81c, Pei82b]. Stone [Pei92f]. Story [MP01]. Strange [PSP^{+}58]. Strange-Particle [PSP^{+}58]. Strategic [Pei81c, Pei86i]. Streamline [Pei40-34]. strike [Pei82c, Pei84d]. Strong [Pei71d, Pei97s, Pei30c, Pei33c]. structure [BP67b, BP67d, ZLZ19]. Struggling [Pei84e]. Student [Pei93a]. Study [CP54, CP97b]. sub [FP39, OPM38, PF43a]. sub-atomic [FP39, OPM38]. sub-cascades [PF43a]. substrates [DSSP76]. Subtle [Pei83d]. Summary [Pei40-35, Pei70b]. Summer [Pei56c]. sums [Pei64e]. sun [HHW99]. Super [FP40b, FP40c, FP64, FP65, FP04]. Super-bomb [FP40b, FP40c, FP64, FP65, FP04]. superb [KK88, KK97]. Superlattices [Pei36e, Pei97-28]. supplement [Pei42m]. Supraconductivity [MCS^{+}35]. Supraconductors [Pei36a, Pei97g]. Surprises [Bar92, Pei79c, Pei91e, Sal92, Paĭ88, Bel80, Tho81, Wei81]. Survey [GP58, PSMF55, Pei55e, Pei55f]. Survive [Pei81c]. Swarm [PE17]. Sword [Pei87b]. SWU [Ber09]. symmetries [Pei89a, Pei91j, Pei92a, Pei97b]. Symmetry [PKA91, PS49, Pei91i]. Symposium [Meh73, AP77, Stu79]. system [PF43b]. Syurprizy [Paĭ88]. Szasz [Pei92b]. Szilard [Pei43a].

T [BPM84]. Tank [Pei46b, PS49]. Tata [Pei86g]. Teaching [BPM84]. Technical [HBP94, Pei50b]. Technology [Pei87b]. Teller [Pei87b, Pei80e]. Temperature [MCS^{+}35, GP78]. Temperaturen [Pei32c]. Temperatures [FKP97, Pei97x, FKP42, Pei32c, Pei34d, Pei34b]. teoreticheskoĭ [Paĭ88]. Terms [Pei97d, Pei48a]. Test [MP74, MP97, Pei89f, Pei66b]. tests [Pei91k]. text [Ano47b]. Their [HHW99, GCP76, Pei85c]. Theoretical [Bar92, Bel80, Bey59, Hal12, Meh73, Pei51, Pei54a, Pei55a, Pei56a, Pei79c, Pei91e, Sal92, Tho81, Tho99, Uhl59, Wei81, AP77, Paĭ88, Pei48b]. theoretische [Pei51, Pei52a]. Theorie [Pei29f, Pei29c, Pei29e, Pei29d, Pei30a, Pei30b, Pei31, Pei32d, Pei32e, Pei33b, Pei33c, Pei33a, Pei34a, NBP^{+}34]. Theories [CP97a, CP53, Pei53c, Pei54c]. Theory [BP35a, Bro77, DPP97, FP47, Koe55, LP65, MP97, Pei29d, Pei34e, Pei35b, Pei36e, Pei40e, Pei40b, Pei40-36, Pei41v, Pei52c, PSW52, Pei55d, Pei57a, Pei57b, Pei59b, Pei64c, Pei65a, Pei65c, Pei73b, Pei74a, Pei74c, Pei96, Pei97a, Pei97e, Pei97n, PSW97, Pei97w, Pei97o, Pei97-44, Pei97-27, Pei97-35, Pei97-28, Pei97s, Pei97r, Pei97p, Pei97q, Pei97t, Pei97-29, Pei01, Rob56, AMP73, AMP97, DPP42, LP31, MP74, McC24, NBP^{+}34, Pei29f, Pei29c, Pei29e, Pei30a, Pei30b, Pei31, Pei32a, Pei32d, Pei32e, Pei33b, Pei33c, Pei33a, Pei34a, Pei34c, Pei35a, Pei35c, Pei36d, Pei37b, Pei52b, PSMF55, Pei55e, Pei55f, Pei56b, Pei59e, Pei60a, Pei63b, Pei64a, Pei73a, Pei97-32, Pei97-31, Pei97-42, Pei09, SW72, Pei57b]. there [Pei71c, Pei86f, Pei89d]. Thermal [Pei97n, Pei97p, Pei30a, Pei40r, Pei64g]. thermischen [Pei30a]. thermo [Pei41m]. thermo-molecular [Pei41m]. Thermodynamics [Pei70c, Pei97-36, Pei56a]. thermomolecular [Pei41o]. thickness [Pei40l]. Thirties [Hen84, Pei93c, Pei84c]. Thomas [Pei93b, BP68b, GP31]. Thompson [Pei81c]. thought [Paĭ90]. Threat [BGM91, Pei91f, Pei91g, Pei47c]. three [BP68c, Pei40-32]. three-nucleon [BP68c]. Tick [Pei94a]. tiefe [Pei32c]. Time [Bha10, Pei70c, Pei94b, Pei97-36, FP42, Pei40t, Pei41f, Pei41k, Pei41l]. Times [Pei82b]. Tizard [Cla65]. Tony [Pei88a]. top [Pei42l]. Topics [Pei97-35, Pei59e]. Tower [Pei69b]. Transition [Pei36a, Pei97g, Pei97u, Pei97x, Pei34d]. Translated [Bey59, Pei54a, Pei55a, Pei56a, Uhl59]. transmutation [Per39]. Transmutations [ECPR37, Pei37a]. Transport [Pei74c, Pei97-44]. Tributes [Tay87]. Trieste [Meh73]. True [MP01]. Truth [Pei86a]. Twentieth [Pei76b, Wal15]. Twentieth-century [Pei76b]. Two [Arm66b, BP58, BP97e, HP42, Pei82d, Pei97-29, Tho99, BP67b, FP43, Oli66a, Oli66b, Oli85, PF43a, Pei30b]. Two-Body [HP42]. two-nucleon [BP67b]. Two-Stage [BP97e, BP58]. Type [ECPR37, Pei37a]. typical [Pei35d]. typiques [Pei35d].

U [Ber09]. U235 [Pei40z]. Udgaonkar [Pei86g]. Umezawa [Pei57b]. Umlaut [Pei87f]. Umwandlungstemperaturen [Pei34d]. Unbestimmtheitsprinzips [LP31]. uncertain [Pei92g]. uncertainty [dB32, LP31, LP65, Pei97e]. Unequal [Pei36e, Pei97-28]. uniform [PFD43, PF43c]. United [BPM84]. Universe [Pei80c]. universities [Bas10]. University [Bel80, Pei79a, Pei80f, Pei92b, Wal15, Pei56c]. Unravelling [FR13]. until [Whi84a, Whi84b]. Uranium [FP40b, FP64, FP65, NBDG40, Pea24, Pei40n, PF40, Pei40-29, Pei40-35, Per39]. US$19.95 [Pei87b]. Use [Pei46b, Pei66b]. uses [GCP76].

V [Pei86g, Paĭ88]. vacuum [Pei34c, Pei35a]. Vakuum [Pei35a]. Values [PPG83]. Vanunu [Pei88b]. Vapor [DPS70]. Variation [Pei40-37, Pei42n, Pei42m, PD44b, Pei45, Pei43b, PD44a]. Variational [PT62, PT97]. Velocity [Pei61c, Pei97-30]. Velocity-dependent [Pei61c, Pei97-30]. Verhalten [Pei30c]. verification [Pei89f]. Verlagsbuchhandlung [Pei51, Pei52a]. Verwey [MP37]. VI [Pei55f]. Vibration [Pei97-34, Pei54b]. Vierte [Pei51]. view [Enz00]. viii [Ano47b, Tho99, Wal15]. Villars [PU68, PU97]. Violation [PSP^{+}58]. Virendra [Pei86g]. Visa [Ano55]. Vol [Pei54a, Pei56a, Pei57a, Pei57a]. voltage [Pei29a]. Volume [Bey59, Pei55a, POZ^{+}77, Pei78a, Uhl59, FK85, MKR87, Pei08]. Vorlesungen [Pei51, Pei52a]. vs [Pei42g].

W [Pei52a, Pei61a]. W. [Enz00]. War [Pei81c, Pei93b, Pei82f, Pei53b]. Wärmeleitung [Pei29c, Pei64g, Pei29f]. Wars [Stu18]. Was [Pei93b]. Wave [BFH^{+}47, BP73, Pei97-41, BP97d, Pei85f, dB29]. Waves [BFvN^{+}44, PPvNF44]. way [Pei87c]. weapons [Arn03, HPPR95a, HPPR95b, Pei70a, Pei81c, Pei89d, Pei91k, Pei81c]. Weiner [Pei80f]. Wellenmechanik [dB29]. went [Pei85a]. Werner [KW93, KRW05, MP77]. Wesley [Bey59, Uhl59]. Where [Pei92h, Pei89d, Pei63d]. White [BPS77]. who [Pei79a, Pei85a]. Widerstandsänderung [Pei31]. Widerstandsgesetzes [Pei32c]. width [Pei40v, Pei40w, Pei42e]. Wiesbaden [Pei51, Pei52a]. Wild [Pei86c]. William [Pei91h]. Wilson [Pei33a]. Winchester [Pei80c]. wir [Pei63d]. Wo [Pei63d]. Wolfgang [Enz00, Pei60b]. women [HHW99]. words [Kel09, Pei85b]. Work [Pei91b, Pei33a, Pei93c]. works [Bet97, BRH^{+}99, BHK^{+}08, Pei08, Tho99]. World [Stu18, Tho99, Wal15, Ber11, PP68]. Wrong [Pei92h].

X [BPS77]. X-ray [BPS77]. xi [Wal15]. xii [Pei51, Wal15]. xiii [Pei51, Pei55a, Pei57a]. xiv [Pei54a, The12, Wal15]. xv [Pei51, Pei57b]. xvi [Pei51, Pei52a, Pei57a]. xviii [Pei56a]. xxiii [Tho99].

Yakov [Pei94c]. Year [FR13]. Years [BPM84, HBP94, Pei84a, Pei84b, Pei92b]. York [Ano47b, Pei54a, Pei55a, Pei56a, Pei57b, Pei87b, Pei93b]. York/Collier [Pei87b].

zur [Pei30b, Pei34a, Pei29f, Pei29c, Pei29e, Pei29d, Pei30a, Pei31, Pei32c, Pei32d, Pei32e, Pei33b, Pei33c, Pei33a]. Zustände [Pei29b]. Zwei [Pei30b].

References

Arnheim:1988:SFR

[ALW^{+}88] Rudolf Arnheim, Thomas E. Lovejoy, David Gordon Wilson, Freeman Dyson, Jane Goodall, Ian Shelton, Kenneth H. Olsen, Irene C. Peden, Richard W. Hamming, Thomas Eisner, Preston Cloud, Matt Cartmill, Samuel C. Florman, Jeremy Bernstein, George A. Miller, Robert M. May, G. Evelyn Hutchinson, Jerome Bruner, Priscilla C. Grew, William Bevan, Elisabeth S. Vrba, Myrdene Anderson, Kevin Padian, Harry Shipman, Victor F. Weisskopf, Walter A. Hill, Patricia D. Moehlman, Melvin Kranzberg, Malak Kotb, Raymond Kurzweil, Marcia Mc-Nutt, Masakazu Konishi, Miriam Rothschild, Edward Teller, Alison Jolly, H. Jane Brockmann, Keith Stewart Thomson, Peter J. Denning, Benoît B. Mandelbrot, Abraham Pais, Paul Mac-Cready, Kip S. Thorne, Ruth Sager, Gerald J. Wasserburg, Neal E. Miller, Rita Levi-Montalcini, Stephen Jay Gould, Edwin H. Land, Michel Boudart, Anne Kernan, Douglas R. Hofstadter, Rosalyn S. Yalow, Bruce H. Tiffney, Mimi Koehl, Walter E. Massey, David P. Billington, John A. W. Kirsch, Abner Shimony, J. Donald Fernie, Brian J. Skinner, Lynn Margulis, Sheldon Lee Glashow, Michael LaBarbera, J. Tuzo Wilson, E. R. Ward Neale, Rudolf Peierls, Roald Hoffmann, Mary L. Good, Donald R. Griffin, Vaclav Smil, Michael S. Turner, Sarah Ann Woodin, Luis Alvarez, George A. Bartholomew, and George B. Schaller. Seventy-five reasons to become a scientist: American Scientist celebrates its seventyfifth anniversary. American Scientist, 76(5):450–463, September 1988. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786

(electronic). URL http://www.jstor.org/stable/27855384;
http://www.jstor.org/stable/pdfplus/27855384.pdf.

Ackland:1986:ANA

[AM86] Len Ackland and Steven McGuire. Assessing the Nuclear Age: Selections from the Bulletin of the Atomic Scientists. Educational Foundation for Nuclear Science, Chicago, IL, USA, 1986. ISBN 0-941682-07-2, 0-941682-08-0 (paperback). xvii + 382 pp. LCCN U264 .A79 1986.

Atalay:1973:PTP

[AMP73] B. Atalay, A. Mann, and Rudolf Peierls. Perturbation theory for projected states. II. Convergence criteria and a soluble model. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 335(1602):251–266, November 1973. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL

https://ui.adsabs.harvard.edu/#abs/1973RSPSA.335..251A.

Atalay:1997:PTP

[AMP97] B. Atalay, A. Mann, and Sir Rudolf Peierls. Perturbation theory for projected states II: Convergence criteria and a soluble model. In Dalitz and Peierls [DP97a], pages 604–619. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book).

LCCN QC21.2 .P42 1997. URL https://ui.adsabs.harvard.
edu/#abs/1997sspr.book..604A.

Anonymous:1938:PAS

[Ano38] Anonymous. Professor Arnold Sommerfeld. Journal of Applied Physics, 9(12):754–755, December 1938. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. Note for the 70th birthday of Arnold Sommerfeld.

Anonymous:1947:BRA

[Ano47a] Anonymous. Book review: Atomic Energy Number, by Rudolf Peierls and John Enogat. Science News (Washington, DC), 2(??): vii–viii, 168, ???? 1947. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic).

Anonymous:1947:SNA

[Ano47b] Anonymous. Science News. 2. Atomic Energy Number. Edited by R. E. Peierls, John Enogat. Penguin Books, Harmondsworth, Middlesex; and New York. 1s. (paper). viii + 168 pp. + 8 plates; text ill. 1947. Quarterly Review of Biology, 22(3):251, September 1947. CODEN QRBIAK. ISSN 0033-5770 (print), 1539-7718 (electronic).

Anonymous:1955:SRV

[Ano55] Anonymous. Some recent visa experiences. Bulletin of the Atomic Scientists, 11(10):368–371, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:HME

[Ano93] Anonymous. Honorary members elected. Europhysics News, 24(9): 229, ???? 1993. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-

1092 (electronic). URL https://www.europhysicsnews.org/
articles/epn/abs/1993/09/epn19932409p229/epn19932409p229.
html.

Anonymous:2024:HCM

[Ano24] Anonymous. Holders of the Copley Medal (1731–2020). In Oxford Dictionary of National Biography, page ?? Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2024.

Aitchison:1977:RPT

[AP77] Ian Johnston Rhind Aitchison and J. E. Paton, editors. Rudolf Peierls and theoretical physics: proceedings of the symposium held in Oxford on July 11th and 12th, 1974, to mark the occasion of the retirement of Professor Sir Rudolf E. Peierls, F.R.S., C.B.E., volume 13 of Progress in nuclear physics. Pergamon Press, New York, NY, USA, 1977. ISBN 0-08-020606-9. LCCN QC1 .P3 1974.

Allen:1937:HCL

[APU37] J. F. Allen, R. Peierls, and M. Zaki Uddin. Heat conduction in liquid helium. Nature, 140(3532):62–63, July 10, 1937. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

https://ui.adsabs.harvard.edu/#abs/1937Natur.140...62A.

Arms:1966:DP

[Arm66a] Nancy Arms. The diffusion project. In A Prophet in Two Countries: the Life of F. E. Simon [Arm66b], chapter 10, pages 100–111. ISBN 0-08-011562-4, 0-08-011563-2, 1-4831-6445-4, 1-4831-3915-8 (e-book). LCCN QC16.S5 A7 1966.

Arms:1966:PTC

[Arm66b] Nancy Arms. A Prophet in Two Countries: the Life of F. E. Simon. Pergamon Press, New York, NY, USA, 1966. ISBN 0-08-011562-4, 0-08-011563-2, 1-4831-6445-4, 1-4831-3915-8 (e-book). viii + 171 pp. LCCN QC16.S5 A7 1966.

Arnold:2000:RBK

[Arn00] Lorna Arnold. Recalling Britain’s key nuclear role. Physics World, 13(2):17–18, February 2000. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://stacks.iop.

org/2058-7058/13/i=2/a=18.

Arnold:2003:HNW

[Arn03] Lorna Arnold. The history of nuclear weapons: The Frisch– Peierls memorandum on the possible construction of atomic bombs of February 1940. Cold War History, 3(3):111–126, 2003. CODEN ???? ISSN 1468-2745 (print), 1743-7962 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
14682740312331391658.

Badash:1986:BRB

[Bad86] Lawrence Badash. Book review: Bird of Passage: Recollections of a Physicist by Rudolf Peierls. Isis, 77(2):332–333, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232666.

Barut:1992:BRM

[Bar92] A. O. Barut. Book review: More Surprises in Theoretical Physics, by Rudolf Peierls. American Journal of Physics, 60(10):957–958, October 1992. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1992AmJPh..60..957P.

Baskaran:2010:BPF

[Bas10] G. Baskaran. Bird of passage at four universities. Resonance, 15(5):411–427, May 2010. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic). URL https://link.springer.

com/article/10.1007/s12045-010-0068-7.

Blackett:1946:MAE

[BBD^{+}46] P. M. S. Blackett, M. Born, P. I. Dee, P. A. M. Dirac, N. Feather, E. A. Guggenheim, H. S. W. Massey, P. B. Moon, N. F. Mott, M. L. E. Oliphant, F. A. Paneth, R. E. Peierls, M. H. L. Pryce, F. E. Simon, Sir George Thompson, O. R. Frisch, and H. W. B. Skinner. Memo to the UN Atomic Energy Commission. Bulletin of the Atomic Scientists, 1(12):6–8, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bloch:1946:LSN

[BCC^{+}46] Felix Bloch, Robert F. Christy, C. Louis Critchfield, Edwin M. McMillan, Emilio Segrè, Edward Teller, Victor Frederick Weisskopf, John H. Williams, Hans A. Bethe, Robert E. Marshak, Lyman G. Parratt, Bruno Rossi, John W. Trischa, Rudolf Peierls, Klaus Fuchs, John Manley, Bernard T. Feld, Julian Schwinger, Joseph Keller, Julius Askin, Chaim Richman, George B. Kistiakowsky, Leonard Schiff, Eugene M. Baroody, and Emil J. Konopinski. Lecture series on nuclear physics, 1943–1944. Report LA-24, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1946.

329 pp. URL https://www.pbagalleries.com/view-auctions/
catalog/id/352/lot/106762/LA-Report-24-Lecture-Series-
on-Nuclear-Physics.

Bell:1980:BRS

[Bel80] J. S. Bell. Book review: Surprises in Theoretical Physics, by R. Peierls. Pp. 166. (Princeton University Press: Princeton, 1979). Nature, 284(5758):705, April 24, 1980. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1980Natur.284..705P.

Bernstein:2003:DWH

[Ber03] Jeremy Bernstein. The drawing, or why history is not mathematics. Physics in Perspective (PIP), 5(3):243–261, September 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/10.
1007/s00016-003-0177-8.

Bernstein:2009:SUM

[Ber09] Jeremy Bernstein. SWU for U and me. arXiv.org, ??(??):1–65,

June 13, 2009. URL https://arxiv.org/abs/0906.2505. This is

a detailed analysis of the proposed uranium isotope separation in [FP41, FP97b, Whi84a, Whi84b].

Bernstein:2011:MCW

[Ber11] Jeremy Bernstein. A memorandum that changed the world. American Journal of Physics, 79(5):440–446, May 2011. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Bethe:1997:SWH

[Bet97] Hans Albrecht Bethe. Selected works of Hans A. Bethe: with commentary, volume 18 of World Scientific series in 20th century physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ISBN 981-02-2876-7. viii + 605 pp. LCCN QC780 .B452 1996. URL http://www.

zentralblattmath.org/zmath/en/search/?an=0916.01027. En-

glish translations (when originals are in German) of 28 of the author’s papers, accompanied by a list of 290 of his publications.

Beyer:1959:BRS

[Bey59] R. T. Beyer. Book review: Statistical Physics, Volume 5 of Course of Theoretical Physics. L. D. Landau and E. M. Lifshitz. Translated from the Russian by E. Peierls and R. F. Peierls. Pp. 484. Pergamon Press, Inc., London. Addison-Wesley Publishing Company, Inc., Reading, 1958. Physics Today, 12(12):58, December 1959. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL https://physicstoday.scitation.org/doi/
abs/10.1063/1.3060614;   https://ui.adsabs.harvard.edu/
#abs/1959PhT....12l..58L.

Bethe:1947:BW

[BFH^{+}47] Hans A. Bethe, Klaus Fuchs, Joseph O. Hirschfelder, John L. Magee, Rudolf E. Peierls, and John von Neumann. Blast wave. U.S. Government Document LA-2000 (TID-4500, 13th edition, supplement), Los Alamos Scientific Laboratory, Los Alamos, NM, USA,

August 1947. 301 pp. URL https://fas.org/sgp/othergov/doe/

lanl/docs1/00320773.pdf. First distributed 27 March 1958.

Bethe:1944:LSH

[BFvN^{+}44] Hans A. Bethe, Klaus Fuchs, John von Neumann, Rudolf Peierls, and W. G. Penney. Lectures in Shock Hydrodynamics and Blast Waves. ????, October 28, 1944. Notes by J. O. Hirschfelder.

Bethe:1991:NTR

[BGM91] Hans A. Bethe, Kurt Gottfried, and Robert S. McNamara. The nuclear threat: A reply to Marvin M. Miller and Rudolf Peierls. New York Review of Books, 38(18):61, November 7, 1991. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.

com/articles/archives/1991/nov/07/the-nuclear-threat-2/

. See [Pei91g].

Bhaduri:2010:MTS

[Bha10] R. K. Bhaduri. My time with Sir Rudolf Peierls. Resonance, 15(5): 384–397, May 2010. CODEN RESOFE. ISSN 0971-8044 (print),

0973-712X (electronic). URL https://www.ias.ac.in/article/
fulltext/reso/015/05/0394-0397.

Bohr:2008:NBC

[BHK^{+}08] Niels Bohr, Ulrich Hoyer, Jørgen Kalckar, J. Rud Nielsen, Rudolf Ernst Peierls, Léon Rosenfeld, Erik Rüdinger, Klaus Stolzenburg, Jens Thorsen, Finn Aaserud, and David Favrholdt, editors. Niels Bohr: collected works. Elsevier, Amsterdam, The Netherlands, 2008. ISBN 0-444-53286-2 (set), 0-444-53283-8 (vol. 1), 0-444-53282-X (vol. 2), 0-444-53287-0 (vol. 3), 0-444-53288-9 (vol. 4), 0-444-53279-X (vol. 5), 0-444-53289-7 (vol. 6), 0-444-53290-0 (vol. 7), 0-444-53284-6 (vol. 8), 0-444-53277-3 (vol. 9), 0-444-53278-1 (vol. 10), 0-444-53280-3 (vol. 11), 0-444-53281-1 (vol. 12), 0-444-53291-9 (vol. 13). various pp. LCCN QC3

.B584; QC771 .B63 2008. URL http://www.elsevier.com/
wps/find/bookdescription.cws_home/BS_NB/description;
http://www.elsevierdirect.com/brochures/nielsbohr/;   http:
//www.sciencedirect.com/science/bookseries/18760503.

Bethe:1939:IBD

[BHP39] H. A. Bethe, F. Hoyle, and R. Peierls. Interpretation of betadisintegration data. Nature, 143(3614):200–201, February 4, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v143/
n3614/pdf/143200a0.pdf;   https://ui.adsabs.harvard.edu/
#abs/1939Natur.143..200B.

Bethe:1997:IBD

[BHP97] H. A. Bethe, F. Hoyle, and R. Peierls. Interpretation of betadisintegration data. In Dalitz and Peierls [DP97a], pages 270–272. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.

adsabs.harvard.edu/#abs/1997sspr.book..270B.

Badash:1986:NFR

[BHT86] Lawrence Badash, Elizabeth Hodes, and Adolph Tiddens. Nuclear fission: Reaction to the discovery in 1939. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 130(2):196–231, June 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.

jstor.org/stable/987181.

Bohr:1960:F

[Boh60] Niels Bohr. Foreword. In Markus Fierz and Victor F. (Victor Frederick) Weisskopf, editors, Theoretical physics in the twentieth century: a memorial volume to Wolfgang Pauli, pages 1–4. Interscience Publishers, New York, NY, USA, 1960. LCCN QC3F54T 1960.

Bethe:1934:Na

[BP34a] H. Bethe and R. Peierls. The “neutrino”. Nature, 133(3362): 532, April 7, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v133/n3362/abs/133532a0.html;   http://
www.nature.com/nature/journal/v133/n3362/pdf/133532a0.
pdf;   https://ui.adsabs.harvard.edu/#abs/1934Natur.133.
.532B.

Bethe:1934:Nb

[BP34b] H. Bethe and R. Peierls. The “neutrino”. Nature, 133(3366): 689–690, May 5, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v133/n3366/pdf/133689b0.pdf;   https://
ui.adsabs.harvard.edu/#abs/1934Natur.133..689B.

Bethe:1934:PDD

[BP34c] Hans A. Bethe and Rudolf E. Peierls. Photoelectric disintegration of the diplon. In International Conference on Physics, London, pages 93–94. Cambridge University Press, Cambridge, UK, 1934.

Bethe:1935:QTD

[BP35a] Hans A. Bethe and Rudolf E. Peierls. Quantum theory of the diplon. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 148(863):146–156, January 1, 1935. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (elec-

tronic). URL http://www.jstor.org/stable/96363;   https:/
/ui.adsabs.harvard.edu/#abs/1935RSPSA.148..146B.

Bethe:1935:SNP

[BP35b] Hans A. Bethe and Rudolf E. Peierls. The scattering of neutrons by protons. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 149(866):176–183, March 1, 1935. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (elec-

tronic). URL http://www.jstor.org/stable/96272;   https:/
/ui.adsabs.harvard.edu/#abs/1935RSPSA.149..176B.

Barker:1949:DER

[BP49] F. C. Barker and R. E. Peierls. On the definition of the “effective range” of nuclear forces. Physical Review, 75(2):312–313, January 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL https://link.aps.org/doi/10.1103/PhysRev.
75.312;   https://ui.adsabs.harvard.edu/#abs/1949PhRv..
.75..312B.

Byers:1958:TSM

[BP58] N. Byers and R. E. Peierls. Two-stage model of Fermi interactions. Il Nuovo Cimento, 10(3):520–524, November 1958. URL https://

ui.adsabs.harvard.edu/#abs/1958NCim...10..520B.

Brink:1967:CNR

[BP67a] D. M. Brink and Rudolf Peierls. Calculation of nuclear reaction matrix elements. Comments on Nuclear and Particle Physics, 1: 66–68, March 1967. CODEN CNPPAV. ISSN 0010-2709.

Brink:1967:CNS

[BP67b] D. M. Brink and Rudolf Peierls. Calculation of nuclear structure directly from two-nucleon phase shifts. Comments on Nuclear and Particle Physics, 1:146–150, July/August 1967. CODEN CNPPAV. ISSN 0010-2709.

Brink:1967:CPF

[BP67c] D. M. Brink and Rudolf Peierls. Calculations with phenomenological forces. Comments on Nuclear and Particle Physics, 1:206–210, November/December 1967. CODEN CNPPAV. ISSN 0010-2709.

Brink:1967:NSR

[BP67d] D. M. Brink and Rudolf Peierls. Nuclear structure and realistic forces. Comments on Nuclear and Particle Physics, 1:32–35, January 1967. CODEN CNPPAV. ISSN 0010-2709.

Brink:1968:ISD

[BP68a] D. M. Brink and Rudolf Peierls. Isobaric spin and dipole states. Comments on Nuclear and Particle Physics, 2:180–181, November/ December 1968. CODEN CNPPAV. ISSN 0010-2709.

Brink:1968:RTF

[BP68b] D. M. Brink and Rudolf Peierls. A realistic Thomas–Fermi approach to finite nuclei. Comments on Nuclear and Particle Physics, 2:28– 32, January/February 1968. CODEN CNPPAV. ISSN 0010-2709.

Brink:1968:TNB

[BP68c] D. M. Brink and Rudolf Peierls. The three-nucleon bound state. Comments on Nuclear and Particle Physics, 2:92–96, May/June 1968. CODEN CNPPAV. ISSN 0010-2709.

Burt:1973:MLW

[BP73] M. G. Burt and Rudolf Peierls. The momentum of a light wave in a refracting medium. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 333(1593):149– 156, May 1973. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-

9169 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1973RSPSA.333..149B.

Bethe:1997:Na

[BP97a] H. Bethe and R. Peierls. The “neutrino”. In Dalitz and Peierls [DP97a], pages 139–140. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42

1997. URL https://ui.adsabs.harvard.edu/#abs/1997sspr.
book..139B.

Bethe:1997:Nb

[BP97b] H. Bethe and R. Peierls. The neutrino. In Dalitz and Peierls [DP97a], page 141. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42

1997. URL https://ui.adsabs.harvard.edu/#abs/1997sspr.
book..141B.

Bethe:1997:SNP

[BP97c] H. A. Bethe and R. Peierls. The scattering of neutrons by protons. In Dalitz and Peierls [DP97a], pages 154–161. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.adsabs.

harvard.edu/#abs/1997sspr.book..154B.

Burt:1997:MLW

[BP97d] M. G. Burt and Rudolf Peierls. The momentum of a light wave in a refracting medium. In Dalitz and Peierls [DP97a], pages 571– 578. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://

ui.adsabs.harvard.edu/#abs/1997sspr.book..571B.

Byers:1997:TSM

[BP97e] N. Byers and R. E. Peierls. Two-stage model of Fermi interactions. In Dalitz and Peierls [DP97a], pages 449–453. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.adsabs.

harvard.edu/#abs/1997sspr.book..449B.

Bulmer:1984:BRP

[BPM84] Martin Bulmer, Rudolf Peierls, and Kenneth Mellanby. Book reviews: Philanthropic Foundations and the Development of a National Society in the United States (pp. 276–285). Private Power for the Public Good: A History of the Carnegie Foundation for the Advancement of Teaching by Ellen Lagemann. Review by: Martin Bulmer. Forty Years into the Atomic Age (pp. 285–291). The Nuclear Chain Reaction Forty Years Later by Robert G. Sachs. Review by: Rudolf Peierls. Scientific Ecology and the Protection of the Environment (pp. 291–293). Planning and Ecology by R. D. Roberts, T. M. Roberts; Lead Pollution: Causes and Control by R. M. Harrison, D. P. H. Laxen. Review by: Kenneth Mellanby. Minerva [Washington, DC], 22(2):276–293, Summer 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/accesspage/article/
10.1007/BF02207952;   http://www.jstor.org/stable/41820566;
http://www.jstor.org/stable/41820567;   http://www.jstor.
org/stable/41820568.

Bohr:1939:NRC

[BPP39] Niels Bohr, Rudolf Peierls, and George Placzek. Nuclear reactions in the continuous energy region. Nature, 144(3639): 200–201, July 29, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v144/n3639/pdf/144200a0.pdf;   https://
ui.adsabs.harvard.edu/#abs/1939Natur.144..200B.

Bohr:1997:NRC

[BPP97] N. Bohr, R. Peierls, and G. Placzek. Nuclear reactions in the continuous energy region. In Dalitz and Peierls [DP97a], pages 260– 263. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://

ui.adsabs.harvard.edu/#abs/1997sspr.book..260B.

Bohr:1986:NP

[BPR86] Niels Bohr, Rudolf E. (Rudolf Ernst) Peierls, and Erik Rüdinger. Nuclear Physics (1929–1952). North-Holland, Amsterdam, The Netherlands, 1986. ISBN 0-7204-1800-3 (set), 0-444-86929-8. xviii + 693 pp. LCCN QC776; SCI; M88.D00077.

Brown:1977:WLX

[BPS77] M. Brown, R. E. Peierls, and E. A. Stern. White lines in X-ray absorption. Physical Review B: Solid State, 15(2):738–744, January 15, 1977. CODEN PLRBAQ. ISSN 0556-2805. URL https://ui.

adsabs.harvard.edu/#abs/1977PhRvB..15..738B.

Bhaduri:1967:ESO

[BPT67] R. K. Bhaduri, R. E. Peierls, and E. L. Tomusiak. Evaluation of the second-order energy shift. Proceedings of the Physical Society, London, 90(4):945–951, April 1967. CODEN PPSOAU. ISSN 0959-5309 (print), 2051-2171 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1967PPS....90..945B.

Brown:1954:CSR

[BPW54] G. E. Brown, R. E. Peierls, and J. B. Woodward. The coherent scattering of γ-rays by K electrons in heavy atoms. I. Method. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 227(1168):51–58, December 1954. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL

https://ui.adsabs.harvard.edu/#abs/1954RSPSA.227...51B.

Brown:1997:CSE

[BPW97] G. E. Brown, R. E. Peierls, and J. B. Woodward. The coherent scattering of γ-rays by K electrons in heavy atoms I: Method. In Dalitz and Peierls [DP97a], pages 427–434. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN

QC21.2 .P42 1997. URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book..427B.

Bohr:1972:CW

[BRH^{+}99] Niels Bohr, L. (Léon) Rosenfeld, Ulrich Hoyer, Erik Rüdinger, Rudolf Peierls, Jørgen Kalckar, Jens Thorsen, Finn Aaserud, J. Rud Nielsen, Klaus Stolzenburg, and David Favrholdt. Collected works. North-Holland, Amsterdam, The Netherlands, 1972–1999. ISBN 0-7204-1800-3 (US: set). various pp. LCCN QC3 .B584.

Brown:1977:SAK

[Bro77] G. E. Brown. Some applications of Kapur–Peierls resonance theory. In Aitchison and Paton [AP77], pages 53–69. ISBN 0-08-020606-9. LCCN QC1 .P3 1974.

Bohr:1939:MNF

[BW39] Niels Bohr and John Archibald Wheeler. The mechanism of nuclear fission. Physical Review (2), 56(5):426–450, September 1939. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v56/i5/p426_1.

Casimir:1986:BRB

[Cas86a] Hendrik B. G. Casimir. Book review: Bird of Passage: Recollections of a Physicist, by Rudolf Peierls. Physics Today, 39(10):123–124, October 1986. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1986PhT....39j.123P.

Casimir:1986:BRR

[Cas86b] Hendrik B. G. Casimir. Book review: Rudolf Peierls, Bird of Passage: Recollections of a Physicist. Physics Today, 39(10):123–??, October 1, 1986. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL https://physicstoday.scitation.
org/doi/abs/10.1063/1.2815177.

Cathcart:1995:OSR

[Cat95] Brian Cathcart. Obituary: Sir Rudolf Peierls. The Independent, ?? (??):??, September 21, 1995. URL https://www.independent.

co.uk/news/people/obituary-sir-rudolf-peierls-1602308.
html.

Charles:1992:HPK

[Cha92] Dan Charles. Heisenberg’s principles kept bomb from Nazis. New Scientist, 135(1837):4–??, September 5, 1992. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL http://www.

newscientist.com/article/mg13518370.300-heisenbergs-principles-

kept-bomb-from-nazis.html. See rebuttals [Pei92d, Wel92].

Clark:1965:T

[Cla65] Ronald William Clark. Tizard. MIT Press, Cambridge, MA, USA, 1965. xvii + 458 + 8 pp. LCCN Q143.T5 C6. Forewords by Sir Solly Zuckerman and Vannevar Bush.

Clary:2023:FMR

[Cla23] David C. Clary. Foreign membership of the Royal Society: Schrödinger and Heisenberg? Notes and Records of the Royal Society of London, 77(3):513–536, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Crammer:1950:AE

[CP50] J. L. (John Lewis) Crammer and R. E. Peierls, editors. Atomic Energy. Penguin, London, UK and New York, NY, USA, 1950. 199 pp. LCCN TK9145 .C75. Revised and expanded version of Science News volume 2 (February 1947).

Chretien:1953:PFF

[CP53] M. Chrétien and R. E. Peierls. Properties of form factors in non-local theories. Il Nuovo Cimento (9), 10(5):668–676, May 1953. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (elec-

tronic). URL https://ui.adsabs.harvard.edu/#abs/1953NCim.
..10..668C.

Chretien:1954:SGI

[CP54] M. Chrétien and R. E. Peierls. A study of gauge-invariant nonlocal interactions. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 223(1155):468–481, May 20, 1954. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-

9169 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1954RSPSA.223..468C.

Cavanagh:1958:MEP

[CP58] P. E. Cavanagh and R. E. Peierls. Measurement of electron polarization [and discussion]. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 246(1247):466–470, August 1958. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Chretien:1997:PFF

[CP97a] M. Chrétien and R. E. Peierls. Properties of form factors in non-local theories. In Dalitz and Peierls [DP97a], pages 388–396. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.

adsabs.harvard.edu/#abs/1997sspr.book..388C.

Chretien:1997:SGI

[CP97b] M. Chrétien and R. E. Peierls. A study of gauge-invariant nonlocal interactions. In Dalitz and Peierls [DP97a], pages 397–410. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.

adsabs.harvard.edu/#abs/1997sspr.book..397C.

Crouse:1968:PRE

[Cro68] William H. Crouse. Peierls, Rudolf Ernst. In McGraw-Hill Modern Men of Science, volume 2, pages 409–410. McGraw-Hill, New York, NY, USA, 1968.

Dalitz:2008:PSR

[Dal08] R. H. Dalitz. Peierls, Sir Rudolf Ernst. In Oxford Dictionary of National Biography, page ?? Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008.

Dattagupta:2010:PED

[Dat10] Sushanta Dattagupta. Peierls’ elucidation of diamagnetism. Resonance, 15(5):428–433, May 2010. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic). URL https://link.

springer.com/article/10.1007/s12045-010-0069-6.

deBroglie:1929:EWG

[dB29] Louis de Broglie. Einführung in die Wellenmechanik. (German) [Introduction to wave mechanics]. Akademische Verlagsgesellschaft, Leipzig, Germany, 1929. iv + 221 + 14 pp. Translation to German by Rudolf Peierls. See [Pei85k, pages 40–41] for the story of the translation experience, and how de Broglie’s views later changed. The German translation appeared before the French original was in print, because French publishers viewed the material as too speculative. Their views changed after de Broglie received the Nobel Prize in Physics for 1929 “for his discovery of the wave nature of electrons”.

deBroglie:1932:FPR

[dB32] L. de Broglie. Sur une forme plus restrictive des relations d’incertitude d’après MM. Landau et Peierls. (French) [On a more restrictive form of uncertainty relations according to MM. Landau and Peierls]. Actualités scientifiques et industrielles (Exposés de physiques théorique I), 31, 1932.

Dirac:1943:AOS

[DFPP43] P. A. M. Dirac, Klaus Fuchs, Rudolf Peierls, and P. Preston. Application to the oblate spheroid hemisphere and oblate hemispheroid. Report number MSD 5, part II., 1943.

Dirac:1936:DCE

[Dir36] P. A. M. Dirac. Does conservation of energy hold in atomic processes? Nature, 137(3460):298–299, February 22, 1936. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v137/
n3460/pdf/137298a0.pdf. See also [Pei36f].

Domb:1963:CMM

[Dom63] Cyril Domb. Clerk Maxwell and modern science; six commemorative lectures. University of London, Athlone Press, London, UK, 1963. vii + 118 pp. LCCN QC71 .D63.

Peierls:1943:DDN

[DP43] B. Davison and Rudolf Peierls. Density distribution near a point source. Report MS-76 (PRO AB4/914), Directorate of Tube Alloys,

????, UK, 1943. URL http://discovery.nationalarchives.
gov.uk/details/r/C1978946.

Das:1973:FMC

[DP73] A. K. Das and R. Peierls. The force on a moving charge in an electron gas. Journal of Physics C: Solid state physics, 6(18):2811– 2821, September 1973. CODEN JPSOAW. ISSN 0022-3719. URL

https://ui.adsabs.harvard.edu/#abs/1973JPhC....6.2811D.

Das:1975:FE

[DP75] A. K. Das and R. Peierls. The force of electromigration. Journal of Physics C: Solid state physics, 8(20):3348–3352, October 1975. CODEN JPSOAW. ISSN 0022-3719. URL https://ui.adsabs.

harvard.edu/#abs/1975JPhC....8.3348D.

Dash:1982:CAF

[DP82] J. G. Dash and Rudolf Peierls. Characteristics of adsorbed films. Physical Review B: Condensed Matter and Materials Physics, 25 (8):5523–5526, April 15, 1982. CODEN PRBMDO. ISSN 1098-

0121. URL https://ui.adsabs.harvard.edu/#abs/1982PhRvB.
.25.5523D.

Dalitz:1986:OPA

[DP86] Richard Henry Dalitz and Sir Rudolf Ernst Peierls. Obituary: Paul Adrien Maurice Dirac. 8 August 1902–20 October 1984. elected F.R.S. 1930. Biographical Memoirs of Fellows of the Royal Society, 32(??):137–185, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL http://www.jstor.org/

stable/770111.

Dalitz:1997:SSP

[DP97a] Richard Henry Dalitz and Sir Rudolf Peierls, editors. Selected scientific papers of Sir Rudolf Peierls: with commentary, volume 19 of World Scientific series in 20th century physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (ebook). xxiii + 805 pp. LCCN QC21.2 .P42 1997. URL

https://ui.adsabs.harvard.edu/#abs/1997sspr.book.....D;
 https://ui.adsabs.harvard.edu/#abs/1997ssps.book.....
P;   https://ui.adsabs.harvard.edu/#abs/1997WSSP...19.
....D;   https://www.worldscientific.com/worldscibooks/
10.1142/3128.

Das:1997:FE

[DP97b] A. K. Das and Rudolf Peierls. The force in electromigration. In Dalitz and Peierls [DP97a], pages 655–659. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book).

LCCN QC21.2 .P42 1997. URL https://ui.adsabs.harvard.
edu/#abs/1997sspr.book..655D.

Das:1997:FMC

[DP97c] A. K. Das and Rudolf Peierls. The force on a moving charge in an electron gas. In Dalitz and Peierls [DP97a], pages 593–603. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.

adsabs.harvard.edu/#abs/1997sspr.book..593D.

Dirac:1942:LIQ

[DPP42] P. A. M. Dirac, Rudolf Peierls, and Maurice H. L. Pryce. On Lorentz invariance in the quantum theory. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 38(2):193–200, January 1942. CODEN PCPSA4. ISSN 0008-

1981. URL https://ui.adsabs.harvard.edu/#abs/1942PCPS..
.38..193D.

Dirac:1997:LIQ

[DPP97] P. A. M. Dirac, R. Peierls, and M. H. L. Pryce. On Lorentz invariance in the quantum theory. In Dalitz and Peierls [DP97a], pages 321–328. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL

https://ui.adsabs.harvard.edu/#abs/1997sspr.book..321D.

Dash:1970:DVP

[DPS70] J. G. Dash, R. E. Peierls, and G. A. Stewart. Desorption, vapor pressure, and absolute entropy of adsorbed films: Application to He^{3}. Physical Review A (Atomic, Molecular, and Optical Physics), 2(3):932–936, September 1970. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL https://ui.

adsabs.harvard.edu/#abs/1970PhRvA...2..932D.

Dash:1976:MEM

[DPS76] J. G. Dash, R. E. Peierls, and M. Schick. Mixing entropy of ^{3}He– 4He monolayers adsorbed on graphite. Journal of Low Temperature Physics, 23(3–4):491–494, May 1976. URL https://ui.adsabs.

harvard.edu/#abs/1976JLTP...23..491D.

Dash:1976:DMS

[DSSP76] J. G. Dash, J. Suzanne, H. Shechter, and R. E. Peierls. Direct measurement of spreading pressures of films adsorbed on crystalline substrates. Surface Science, 60(2):411–424, November 2, 1976. CODEN SUSCAS. ISSN 0039-6028 (print), 1879-

2758 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1976SurSc..60..411D.

Dash:1978:NSP

[DSSP78] J. G. Dash, J. Suzanne, H. Shechter, and R. E. Peierls. Note on the spreading pressure of an adsorbed film. Surface Science, 72(1):219– 222, March 1, 1978. CODEN SUSCAS. ISSN 0039-6028 (print),

1879-2758 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1978SurSc..72..219D.

Ellis:1937:DTN

[ECPR37] C. D. Ellis, J. D. Cockcroft, R. Peierls, and H. O. W. Richardson. A discussion on β-type of nuclear transmutations: 4 March 1937. Opening address. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 161(907):447– 460, August 20, 1937. CODEN PRLAAZ. ISSN 0080-4630 (print),

2053-9169 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1937RSPSA.161..447E.

Enz:2000:WPB

[Enz00] C. P. Enz. Wolfgang Pauli’s 100th birthday: a personal view by Prof. C. P. Enz, W. Pauli’s last assistant. Europhysics News, 31(4): 12, July/August 2000. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (electronic).

Edwards:1954:FEF

[EP54] S. F. Edwards and R. E. Peierls. Field equations in functional form. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 224(1156):24–33, June 1954. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL

https://ui.adsabs.harvard.edu/#abs/1954RSPSA.224...24E.

Edwards:1997:FEF

[EP97] S. F. Edwards and R. E. Peierls. Field equations in functional form. In Dalitz and Peierls [DP97a], pages 411–420. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.adsabs.

harvard.edu/#abs/1997sspr.book..411E.

Feld:1986:BRB

[Fel86] Bernard T. Feld. Book review: Bird of Passage, by Rudolf Peierls. Bulletin of the Atomic Scientists, 42(5):44, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

French:1985:NBC

[FK85] A. P. (Anthony Philip) French and P. J. Kennedy, editors. Niels Bohr: a centenary volume. Harvard University Press, Cambridge, MA, USA, 1985. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). xiv + 403 pp. LCCN QC16.B63 N49 1985.

Fuchs:1942:ESA

[FKP42] Klaus Fuchs, G. J. Kynch, and Rudolf Peierls. The equation of state of air at high temperatures. Report MS-61 (PRO AB4/897), Directorate of Tube Alloys, ????, UK, 1942.

Fuchs:1997:ESA

[FKP97] K. Fuchs, G. J. Kynch, and R. Peierls. The equation of state of air at high temperatures. In Dalitz and Peierls [DP97a], pages 329– 356. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://

ui.adsabs.harvard.edu/#abs/1997sspr.book..329F.

Frisch:1939:RSA

[FP39] O. R. Frisch and R. Peierls. Radioactivity and sub-atomic phenomena. Annual Reports on the Progress of Chemistry, 36: 7–32, 1939. CODEN ARPCAW. ISSN 0365-6217 (print),

1754-7512 (electronic). URL http://pubs.rsc.org/en/content/
articlelanding/1939/ar/ar9393600007.

Frisch:1940:FPM

[FP40a] Otto R. Frisch and Rudolf E. Peierls. The Frisch–Peierls memorandum, part 1. ALSOS Web site., March 1940. URL http://alsos.

wlu.edu/information.aspx?id=3171.

Frisch:1940:CSB

[FP40b] Otto Robert Frisch and Rudolf Peierls. On the construction of a ‘super-bomb’ based on a nuclear chain reaction in uranium. Technical report, University of Birmingham, Birmingham, UK,

March 1940. URL http://en.wikipedia.org/wiki/Frisch%
E2%80%93Peierls_memorandum;   http://www.atomicarchive.
com/Bios/Frisch.shtml;   http://www.atomicarchive.com/
Bios/Peierls.shtml;   http://www.atomicarchive.com/Docs/
Begin/FrischPeierls.shtml. Reprinted in [FP64].

Frisch:1940:PRS

[FP40c] Otto Robert Frisch and Rudolf Peierls. The properties of a radioactive “super-bomb”. Memorandum, University of Birmingham, Birmingham, UK, March 1940. URL http://en.

wikipedia.org/wiki/Frisch%E2%80%93Peierls_memorandum;
http://www.atomicarchive.com/Bios/Frisch.shtml;   http:
//www.atomicarchive.com/Bios/Peierls.shtml;   http://www.
atomicarchive.com/Docs/Begin/FrischPeierls.shtml;   http:
//www.atomicarchive.com/Docs/Begin/FrischPeierls2.shtml.

Reprinted in [Cla65].

Fuchs:1941:SI

[FP41] K. Fuchs and R. Peierls. Separation of isotopes. Report MS-12A (PRO AB4/838), Directorate of Tube Alloys, ????, UK, 1941. See [Ber09] for a detailed analysis of this paper.

Fuchs:1942:ETS

[FP42] Klaus Fuchs and Rudolf Peierls. Equilibrium time in a separation plant. Report MS-47A (PRO AB4/882), Directorate of Tube Alloys, ????, UK, May 1942. 1 + 8 pp.

Frisch:1943:LAT

[FP43] Otto Robert Frisch and Rudolf Peierls. Light absorption by two black hemispheres. Report MS-77 (PRO AB4/915), Directorate of Tube Alloys, ????, UK, 1943.

Ferretti:1947:RDT

[FP47] B. Ferretti and R. E. Peierls. Radiation damping theory and the propagation of light. Nature, 160(4068):531–532, October 18, 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1947Natur.160..531F.

Frisch:1964:CSB

[FP64] Otto Robert Frisch and Rudolf Peierls. On the construction of a ‘super-bomb’ based on a nuclear chain reaction in uranium [Part 1]. In Britain and Atomic Energy, 1939–1945 [Gow64], pages 389– 393. LCCN QC773.A1 G6 1964. With an introductory chapter by Kenneth Jay.

Frisch:1965:CSB

[FP65] Otto Robert Frisch and Rudolf Peierls. On the construction of a ‘super-bomb’ based on a nuclear chain reaction in uranium [Part 2]. In Tizard [Cla65], pages 215–217. LCCN Q143.T5 C6. Forewords by Sir Solly Zuckerman and Vannevar Bush.

Frisch:1997:FPM

[FP97a] O. R. Frisch and R. Peierls. The Frisch–Peierls memorandum of 1940. In Dalitz and Peierls [DP97a], pages 277–282. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.adsabs.

harvard.edu/#abs/1997sspr.book..277F.

Fuchs:1997:SI

[FP97b] K. Fuchs and R. Peierls. Separation of isotopes. In Dalitz and Peierls [DP97a], pages 303–320. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN

QC21.2 .P42 1997. URL https://ui.adsabs.harvard.edu/#abs/

1997sspr.book..303F. See [Ber09] for a detailed analysis of this paper.

Frisch:2004:MPR

[FP04] Otto R. Frisch and Rudolf Peierls. Memorandum on the properties of a radioactive super-bomb. In Kelly [Kel04], pages 177–180. ISBN 981-256-040-8. LCCN QC773.A1 R46 2004.

Fernandez:2013:UMA

[FR13] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (ebook). xviii + 522 pp. LCCN QC773 .F47 2013.

Freeman:1956:BRR

[Fre56] Ira M. Freeman. Book review: R. E. Peierls, The Laws of Nature. Physics Today, 9(10):28–29, October 1, 1956. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/abs/10.1063/1.3059792;
https://ui.adsabs.harvard.edu/#abs/1956PhT.....9j..28P.

Gowing:1974:IDB

[GA74] Margaret Gowing and Lorna Arnold. Independence and Deterrence: Britain and Atomic Energy, 1945–1952. MacMillan Publishing Company, New York, NY, USA, 1974. ISBN 0-333-15781-8 (vol. 1). xiv + 559 pp. LCCN QC773.3.G7 G68 1974b.

Garcia-Calderon:1976:RST

[GCP76] Gastón García-Calderón and Rudolf Peierls. Resonant states and their uses. Nuclear Physics A, 265(3):443–460, July 1976. CODEN NUPABL. ISSN 0375-9474 (print), 1873-1554 (electronic). URL

https://ui.adsabs.harvard.edu/#abs/1976NuPhA.265..443G.

Goldschmidt:1987:PAF

[Gol87] Bertrand Goldschmidt. Pionniers de l’atome. (French) [Pioneers of the atom]. Stock, Paris, France, 1987. ISBN 2-234-02070-0. 484 + 8 pp. LCCN QC773 .G64 1987. URL http://gallica.bnf.fr/ ark:/12148/bpt6k4803658x. See also English translation [Gol90].

Goldschmidt:1990:AR

[Gol90] Bertrand Goldschmidt. Atomic Rivals. Rutgers University Press, New Brunswick, NJ, USA, 1990. ISBN 0-8135-1518-1 (hardcover), 0-8135-1519-X (paperback). xvii + 372 + 8 pp. LCCN QC773 .G6413 1990. English translation of [Gol87].

Gowing:1964:BAEb

[Gow64] Margaret Gowing. Britain and Atomic Energy, 1939–1945. St. Martin’s Press, New York, NY, USA, 1964. xvi + 464 pp. LCCN QC773.A1 G6 1964. With an introductory chapter by Kenneth Jay.

Guth:1931:AFT

[GP31] Eugen Guth and Rudolf E. Peierls. Application of the Fermi– Thomas model to positive ions. Physical Review (2), 37(2):217, January 15, 1931. CODEN PHRVAO. ISSN 0031-899x (print),

1536-6065 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1931PhRv...37..217G.

Grace:1958:SES

[GP58] M. A. Grace and R. E. Peierls. Survey of the experimental situation in β-decay: Discussion. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 246(1247):454–462, August 1958. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-

9169 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1958RSPSA.246..462P.

Goldstein:1978:MLT

[GP78] F. S. Goldstein and R. Peierls. A model for low-temperature anharmonicity. Journal of Physics C: Solid state physics, 11(9):1921– 1934, May 1978. CODEN JPSOAW. ISSN 0022-3719. URL https:

//ui.adsabs.harvard.edu/#abs/1978JPhC...11.1921G.

Guinea:1986:SSC

[GPS86] F. Guinea, R. E. Peierls, and R. Schrieffer. Shape of solitons in classically forbidden states: “Lorentz expansion”. Physica Scripta, 33(3):282–283, March 1986. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1986PhyS...33..282G.

Guinea:1997:SSC

[GPS97] F. Guinea, R. E. Peierls, and R. Schrieffer. Shape of solitons in classically forbidden states: “Lorentz expansion”. In Dalitz and Peierls [DP97a], pages 775–777. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42

1997. URL https://ui.adsabs.harvard.edu/#abs/1997sspr.
book..775G.

Guinea:2002:SSC

[GPS02] F. Guinea, R. E. Peierls, and R. Schrieffer. Shape of solitons in classically forbidden states: “Lorentz expansion”. In Nicholas E. Bonesteel and Lev P. Gor’kov, editors, Selected Papers of J. Robert Schrieffer in Celebration of his 70th Birthday, pages 253–254. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2002. ISBN 981-277-704-0. URL https://ui.

adsabs.harvard.edu/#abs/2002sprs.book..253G.

Gregory:2000:BB

[Gre00] David Gregory. Brum and the bomb. BBC News Web story.,

February 17, 2000. URL http://news.bbc.co.uk/2/hi/uk_news/
643913.stm.

Gregersen:2011:SRE

[Gre11] Erik Gregersen. Sir Rudolf Ernst Peierls. Encyclopedia Britannica article., December 12, 2011. URL https://www.britannica.com/

biography/Rudolf-Ernst-Peierls.

Greenspan:2020:ASD

[Gre20] Nancy Thorndike Greenspan. Atomic Spy: The Dark Lives of Klaus Fuchs. Viking, New York, NY, USA, 2020. ISBN 0-593-08339-3 (hardcover), 0-593-08340-7 (paperback), 0-593-08341-5 (e-book), 2-01-905313-8. xii + 401 pp. LCCN UB271.R9 F8355 2020.

Halpern:2012:QHP

[Hal12] Paul Halpern. Quantum humor: The playful side of physics at Bohr’s Institute for Theoretical Physics. Physics in Perspective (PIP), 14(3):279–299, September 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-011-0071-8.

Hoddeson:1994:ASB

[HBP94] Lillian Hoddeson, Gordon Baym, and Rudolf Peierls. Atomic science. (book reviews: Critical Assembly. A Technical History of Los Alamos During the Oppenheimer Years, 1943–1945). Science, 263(5150):1162–1163, February 25, 1994. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1994Sci...263.1162H.

Hendry:1984:CPT

[Hen84] John Hendry, editor. Cambridge Physics in the Thirties. Adam Hilger Ltd., Bristol, UK, 1984. ISBN 0-85274-761-6. xi + 209 pp. LCCN QC51.G72 C352 1984.

Howes:1999:TDS

[HHW99] Ruth (Ruth Hege) Howes, Caroline L. Herzenberg, and Ellen C. Weaver. Their day in the sun: women of the Manhattan Project. Labor and social change. Temple University Press, Philadelphia, PA, USA, 1999. ISBN 1-56639-719-7 (hardcover), 1-59213-192-1 (paperback), 0-585-38881-4 (e-book). viii + 264 pp. LCCN QC773.3.U5 H68 1999.

Hepner:1942:NCF

[HP42] W. Hepner and R. Peierls. Non-central forces in the nuclear twobody problem. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 181(984):43–57, September 1942. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-

9169 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1942RSPSA.181...43H.

Hill:1995:DBNa

[HPPR95a] C. R. Hill, R. S. Pease, Rudolf Peierls, and J. Rotblat. Does Britain need nuclear weapons? In ????, pages vi, 78. British Pugwash Group, London, UK, 1995.

Hill:1995:DBNb

[HPPR95b] C. R. Hill, R. S. Pease, Rudolf Peierls, and J. Rotblat. Does Britain need nuclear weapons? In ????, pages 4–?? British Pugwash Group, London, UK, 1995.

Kelly:2004:RMP

[Kel04] Cynthia C. Kelly, editor. Remembering the Manhattan Project: Perspectives on the Making of the Atomic Bomb and Its Legacy. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2004. ISBN 981-256-040-8. LCCN QC773.A1 R46 2004.

Kelly:2007:MPB

[Kel07] Cynthia C. Kelly, editor. The Manhattan Project: the birth of the atomic bomb by its creators, eyewitnesses, and historians. Black Dog and Leventhal Publishers, New York, NY, USA, 2007. ISBN 1-57912-747-9, 1-57912-808-4 (paperback), 1-60376-206-X (e-book). xiv + 495 pp. LCCN QC773.3.U5 M27 2007. URL http://www.

loc.gov/catdir/toc/ecip0720/2007023984.html.

Kelly:2009:MPB

[Kel09] Cynthia C. Kelly, editor. The Manhattan Project: the birth of the atomic bomb in the words of its creators, eyewitnesses, and historians. Black Dog and Leventhal, New York, NY, USA, 2009. ISBN 1-57912-808-4 (paperback), 1-57912-747-9, 1-60376-206-X (e-book). xiv + 495 pp. LCCN QC773.3.U6 M27 2009.

Khalatnikov:1989:LPM

[Kha89] I. M. (Isaak Markovich) Khalatnikov, editor. Landau, the physicist and the man: recollections of L. D. Landau. Pergamon Press, New York, NY, USA, 1989. ISBN 0-08-036383-0. viii + 323 + 24 pp. LCCN QC16.L25 L38 1989.

Kurti:1988:CSA

[KK88] Nicholas Kurti and Giana Kurti, editors. But the crackling is superb: an anthology on food and drink by Fellows and Foreign Members of the Royal Society. Adam Hilger Ltd., Bristol, UK, 1988. ISBN 0-85274-301-7. xvii + 261 pp. LCCN TX651 .B89 1988. With a foreword by Sir George Porter.

Kurti:1997:CSA

[KK97] Nicholas Kurti and Giana Kurti, editors. But the crackling is superb: an anthology on food and drink by Fellows and Foreign Members of the Royal Society. IOP Publishing, Bristol, UK, 1997. ISBN 0-7503-0488-X (paperback). xviii + 265 pp. LCCN TX651 .B89

1998. URL http://www.loc.gov/catdir/enhancements/fy0745/

97041571-d.html. With a foreword by Lord Porter of Luddenham.

Knight:1995:RP

[Kni95] Professor P. L. Knight. Rudolf Peierls (1907–1995). Contemporary Physics, 36(6):375, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Koehler:1955:BRR

[Koe55] J. S. Koehler. Book reviews: R. E. Peierls, Quantum Theory of Solids. Intern. Ser. of Monogr. on Physics. Science, 122(3177):975– 976, November 18, 1955. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1955Sci...122..975P.

Kapur:1937:PPB

[KP37] P. L. Kapur and R. Peierls. Penetration into potential barriers in several dimensions. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 163(915):606–610, December 1937. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-

9169 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1937RSPSA.163..606K.

Kapur:1938:DFN

[KP38] P. L. Kapur and R. Peierls. The dispersion formula for nuclear reactions. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 166(925):277–295, May 1938. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-

9169 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1938RSPSA.166..277K.

Kapur:1997:DFN

[KP97a] P. L. Kapur and R. Peierls. The dispersion formula for nuclear reactions. In Dalitz and Peierls [DP97a], pages 239–257. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.adsabs.

harvard.edu/#abs/1997sspr.book..239K.

Kapur:1997:PPB

[KP97b] P. L. Kapur and R. Peierls. Penetration into potential barriers in several dimensions. In Dalitz and Peierls [DP97a], pages 230– 234. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://

ui.adsabs.harvard.edu/#abs/1997sspr.book..230K.

Kleint:2005:WHB

[KRW05] Christian Kleint, Helmut Rechenberg, and Gerald Wiemers, editors. Werner Heisenberg, 1901–1976: Beiträge, Berichte, Briefe: Festschrift zu seinem 100. Geburtstag. (German) [Werner Heisenberg, 1901–1976: Contributions, Reports, and Letters: Celebration of his 100th Birthday], volume 62 of Abhandlungen der Sächsischen Akademie der Wissenschaften zu Leipzig, Mathematisch-Naturwissenschaftliche Klasse; Aufsatzsammlung. Verlag der Sächsischen Akademie der Wissenschaften zu Leipzig, Leipzig, Germany, 2005. ISBN 3-7776-1402-5. ISSN 0365-6470. 424 pp. LCCN QC16.H35 W474 2005. URL http://www.loc.gov/

catdir/toc/fy0613/2006402966.html.

Kursunoglu:1987:RAG

[KW87] Behram Kurşunoğlu and Eugene Paul Wigner, editors. Reminiscences about a great physicist: Paul Adrien Maurice Dirac. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-34013-6. xviii + 297 pp. LCCN QC16.D57 R46 1987. URL http:

//adsabs.harvard.edu/abs/1987ragp.book.....D;   http:/
/www.loc.gov/catdir/description/cam023/86033409.html;
http://www.loc.gov/catdir/toc/cam028/86033409.html.

Kleint:1993:WHL

[KW93] Christian Kleint and Gerald Wiemers, editors. Werner Heisenberg in Leipzig, 1927–1942, volume 58(2) of Abhandlungen der Sächsischen Akademie der Wissenschaften zu Leipzig, Mathematisch-Naturwissenschaftliche Klasse, 0365-6470. Akademie-Verlag, Berlin, Germany, 1993. ISBN 3-05-501585-1. 263 pp. LCCN QC16.H35 W48 1933. EUR 607 (!).

Langsdorf:1957:BLN

[Lan57] Alexander Langsdorf, Jr. Books: The Laws of Nature, by R. E. Peierls. Bulletin of the Atomic Scientists, 13(2):72, February 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Laucht:2012:EGK

[Lau12] Christoph Laucht. Elemental Germans: Klaus Fuchs, Rudolf Peierls, and the making of British nuclear culture 1939–59. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire, UK, 2012. ISBN 0-230-35487-4 (hardcover). xiv + 274 pp. LCCN QC773.3.G7 L38 2012.

Lee:2007:BPC

[Lee07a] Sabine Lee. The Bethe–Peierls correspondence. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2007. ISBN 981-277-135-2. xi + 506 pp. LCCN

QC16.B46 A4 2007eb. URL http://site.ebrary.com/lib/
yale/Doc?id=10255467;   http://www.loc.gov/catdir/toc/
fy0805/2008274437.html;   http://www.worldscientific.com/
worldscibooks/10.1142/6595.

Lee:2007:REP

[Lee07b] Sabine Lee. Rudolf Ernst Peierls. 5 June 1907–19 September 1995. Biographical Memoirs of Fellows of the Royal Society, 53:265–284, 2007. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL http://rsbm.royalsocietypublishing.org/
content/53/265;   http://rsbm.royalsocietypublishing.org/
content/suppl/2009/05/01/53.0.265.DC1.

Lee:2007:SRP

[Lee09] Sabine Lee, editor. Sir Rudolf Peierls: selected private and scientific correspondence. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2007–2009. ISBN 981-256-503-5 (vol. 1), 981-279-706-8 (vol. 2). xiv + 866 (volume 1), xii + 1108 (volume 2) pp. LCCN QC16.P375 A4 2007. URL http:

//wspc-prod.literatumonline.com/worldscibooks/10.1142/
6790;   http://www.worldscientific.com/worldscibooks/10.
1142/5941.

Landau:1958:SP

[LL58] L. D. Landau and E. M. Lifshitz. Statistical Physics, volume 5 of Course of Theoretical Physics. Pergamon Press, New York, NY, USA, 1958. x + 484 pp. Translated by E. (Eugenia) Peierls and R. F. Peierls.

Landau:1930:QIK

[LP30] L. Landau and R. Peierls. Quantenelektrodynamik im Konfigurationsraum. (German) [Quantum electrodynamics in configuration space]. Zeitschrift für Physik, 62(3–4):188–200, March 1930. CODEN ZEPYAA. ISSN 0044-3328. URL https://ui.adsabs.

harvard.edu/#abs/1930ZPhy...62..188L.

Landau:1931:EUR

[LP31] L. Landau and R. Peierls. Erweiterung des Unbestimmtheitsprinzips für die relativistische Quantentheorie. (German) [Extension of the Uncertainty Principle for relativistic quantum theory]. Zeitschrift für Physik, 69(1–2):56–69, January 1931. CODEN

ZEPYAA. ISSN 0044-3328. URL https://ui.adsabs.harvard.
edu/#abs/1931ZPhy...69...56L.

Landau:1965:EUP

[LP65] Lev Landau and Rudolf Peierls. Extension of the Uncertainty Principle to relativistic quantum theory. In ter Haar [tH65], pages 40–51. LCCN QC16.L25 A2 1965.

McCauley:2024:HRT

[McC24] Joseph L. McCauley. Heisenberg’s 1939 reactor theory, Serber’s 1943 Los Alamos Primer, and Heisenberg’s 1945 Farm Hall critical mass calculation. American Journal of Physics, 92(10):765–774, October 2024. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). See comments and corrections [Ree24a].

McCauley:2025:PCR

[McC25] Joseph L. McCauley. Predictions of critical radii for reactors and bombs 1939–45 including the Frisch–Peierls memorandum. European Physical Journal H, 50(1):??, December 2025. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-024-00088-1.

McLennan:1935:DSO

[MCS^{+}35] J. C. McLennan, J. D. Cockcroft, D. Shoenberg, W. H. Keesom, W. Meissner, R. De L. Kronig, L. Brillouin, N. Kurti, F. Simon, R. Peierls, F. London, K. Mendelssohn, J. D. Bernal, N. F. Mott, and M. Blackman. A discussion on supraconductivity and other low temperature phenomena. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 152(875): 1–46, October 1935. CODEN PRLAAZ. ISSN 0080-4630 (print),

2053-9169 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1935RSPSA.152....1M.

Mehra:1973:PCN

[Meh73] Jagdish Mehra, editor. The physicist’s conception of nature: Symposium on the Development of the Physicist’s Conception of Nature in the 20th century. Held at the International Centre for Theoretical Physics, Miramare, Trieste, Italy, 18–25 September 1972. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1973. ISBN 90-277-0345-0, 90-277-2536-5. LCCN QC173.96 .S95 1972. URL http://www.springer.com/

us/book/9789027703453.

Metropolis:1987:NDP

[MKR87] N. (Nicholas) Metropolis, Donald M. Kerr, and Gian-Carlo Rota, editors. New Directions in Physics: The Los Alamos 40th Anniversary Volume. Academic Press, New York, NY, USA, 1987. ISBN 0-12-492155-8. xii + 292 pp. LCCN QC44 .N49

1987. URL http://www.osti.gov/energycitations/product.
biblio.jsp?osti_id=6120718&query_id=0.

Mott:1937:DPB

[MP37] N. F. Mott and R. Peierls. Discussion of the paper by de Boer and Verwey. Proceedings of the Physical Society, London, 49(45): 72–73, August 1937. CODEN PPSOAU. ISSN 0959-5309 (print),

2051-2171 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1937PPS....49...72M.

Mann:1974:TPP

[MP74] A. Mann and R. E. Peierls. Test of projected perturbation theory on a simplified model of H_{2}^{+}. Chemical Physics Letters, 24(1):136– 138, January 1, 1974. CODEN CHPLBC. ISSN 0009-2614 (print),

1873-4448 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1974CPL....24..136M.

Mott:1977:WHD

[MP77] Sir Nevill Mott and Sir Rudolf Peierls. Werner Heisenberg. 5 December 1901–1 February 1976. Biographical Memoirs of Fellows of the Royal Society, 23:212–251, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL http://www.

jstor.org/stable/769614.

Mann:1997:TPP

[MP97] A. Mann and R. E. Peierls. Test of projected perturbation theory on a simplified model of H_{2}^{+}. In Dalitz and Peierls [DP97a], pages 620– 621. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://

ui.adsabs.harvard.edu/#abs/1997sspr.book..620M.

Medawar:2001:HGT

[MP01] J. S. Medawar and David Pyke. Hitler’s Gift: the True Story of the Scientists Expelled by the Nazi Regime. Arcade Publishing, New York, NY, USA, 2001. ISBN 1-55970-564-7. xx + 268 pp. LCCN Q141 .M385 2001. Foreword by Max Perutz.

Nabarro:1947:DSC

[Nab47] F. R. N. Nabarro. Dislocations in a simple cubic lattice. Proceedings of the Physical Society, London, 59(2):256–272, March 1947. CODEN PPSOAU. ISSN 0959-5309 (print), 2051-2171 (elec-

tronic). URL https://iopscience.iop.org/article/10.1088/
0959-5309/59/2/309.

Nair:1988:IRP

[Nai88] Ranjit Nair. Interview with Rudolf Peierls. 2001 (Bombay), ??(??): 26–29, 36, 44, June 1988.

Navarro:2006:EAD

[Nav06] Jaume Navarro. Early attempts to detect the neutrino at the Cavendish Laboratory. Physics in Perspective (PIP), 8(1):64–82, March 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/10.
1007/s00016-005-0249-z.

Nier:1940:NFS

[NBDG40] Alfred O. Nier, E. T. Booth, J. R. Dunning, and A. V. Grosse. Nuclear fission of separated uranium isotopes. Physical Review, 57(6):546, March 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/doi/ 10.1103/PhysRev.57.546. This paper reports experiments that confirm the Bohr–Wheeler prediction [BW39] that U-235 is much more fissile than U-238. That information was critical for the production of both a nuclear reactor, and a nuclear bomb.

Nordheim:1934:TEM

[NBP^{+}34] L. Nordheim, H. A. Bethe, R. Peierls, A. Sommerfeld, L. Brillouin, A. W. Maue, R. H. Fowler, and H. Jones. La théorie des électrons dans les métaux. (French) [The theory of electrons in metals]. Helvetica Physica Acta, 7(Supplement 2):1–87, 1934. CODEN HPACAK. ISSN 0018-0238 (print), 2297-1971 (electronic).

Nickel:1967:BEC

[NTP67] B. G. Nickel, E. L. Tomusiak, and R. E. Peierls. Binding-energy calculations for finite nuclei. Nuclear Physics A, 102(3):655–666, October 1967. CODEN NUPABL. ISSN 0375-9474 (print), 1873-

1554 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1967NuPhA.102..655N.

Nabarro:1952:SHL

[NVP52] F. R. N. Nabarro, J. H. O. Varley, and R. E. Peierls. The stability of hexagonal lattices with a simple law of force. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 48(2):316–328, April 1952. CODEN PCPSA4. ISSN 0008-

1981. URL https://ui.adsabs.harvard.edu/#abs/1952PCPS..
.48..316N.

Oliphant:1947:SSA

[OJCP^{+}47] M. L. Oliphant, F. Joliot-Curie, R. E. Peierls, Albert Einstein, Federation of American Scientists, and General Leslie Groves. Statements on the second anniversary of Hiroshima. Bulletin of the Atomic Scientists, 3(9):235–236, 252, September 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oliphant:1966:TEa

[Oli66a] Mark L. Oliphant. The two Ernests — I. Physics Today, 19 (9):35–49, September 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://scitation.aip.org/

content/aip/magazine/physicstoday/article/19/9/10.1063/

1.3048466. Reprinted in [Oli85].

Oliphant:1966:TEb

[Oli66b] Mark L. Oliphant. The two Ernests — II. Physics Today, 19 (10):41–51, October 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://scitation.aip.

org/content/aip/magazine/physicstoday/article/19/10/10.

1063/1.3047765. Reprinted in [Oli85].

Oliphant:1985:TE

[Oli85] Mark L. Oliphant. The two Ernests. In Weart and Phillips [WP85], pages 173–193. ISBN 0-88318-468-0 (paperback). LCCN QC7 .H694 1985. Reprint of [Oli66a, Oli66b].

Oliphant:1938:RSA

[OPM38] M. L. Oliphant, R. Peierls, and P. B. Moon. Radioactivity and subatomic phenomena. Annual Reports on the Progress of Chemistry, 35:7–35, 1938. CODEN ARPCAW. ISSN 0365-6217 (print), 1754-7512 (electronic).

Paierls:1988:CSV

[Paĭ88] R. Paĭerls. Syurprizy v teoreticheskoĭ fizike. (Russian) [Surprises in theoretical physics], volume 6 of Biblioteka Teoreticheskoĭ Fiziki [Library of Theoretical Physics]. “Nauka”, Moscow, USSR, 1988. ISBN 5-02-013840-1. 176 pp. Translated from the English by I. M. Kaganova, Translation edited and with a preface by M. I. Kaganov.

Paierls:1990:STD

[Paĭ90] Rudol’f Paĭerls. Some thought Dirac odd. In Paul Dirac and 20th century physics (Russian), pages 112–114. “Nauka”, Moscow, Russia, 1990. Translated from the English by T. B. Romanovskaya.

Peierls:1944:ELD

[PD44a] Rudolf Peierls and B. Davison. Efficiency loss due to the lateral variation of the branching ratio for a nonrectangular membrane pair. Report MS-110 (PRO AB4/), Directorate of Tube Alloys, ????, UK, 1944.

Peierls:1944:VBR

[PD44b] Rudolf Peierls and B. Davison. Variation on branching ratio. Report MS-N11 (PRO AB4/949), Directorate of Tube Alloys, ????, UK, April 1944.

Peierls:1949:FS

[PD49] Rudolf Peierls and Sir Henry Dale. Freedom of science. Bulletin of the Atomic Scientists, 5(4):106–109, April 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1947:AEN

[PE47] Rudolf Peierls and John Enogat. Atomic Energy Number. Penguin Books, Harmondsworth, Middlesex, UK, 1947. viii + 168 + 8 pp.

Pei:2017:APS

[PE17] Zongrui Pei and Markus Eisenbach. Acceleration of the Particle Swarm Optimization for Peierls–Nabarro modeling of dislocations in conventional and high-entropy alloys. Computer Physics Communications, 215(??):7–12, June 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465517300346.

Pearson:2024:CFP

[Pea24] J. M. Pearson. Comments on the Frisch–Peierls estimate of the critical mass of a uranium fission bomb. Nuclear Technology, ??(??): 1–5, February 2024. CODEN NUTYBB. ISSN 1943-7471 (print), 1943-7471 (electronic).

Peierls:1929:AIO

[Pei29a] R. Peierls. Über die Abhängigkeit der Ionisierungsspannung von der Ordnungszahl. (German) [On the dependence of the ionization voltage on the atomic number]. Zeitschrift für Physik, 55(11–12):738– 743, November 1929. CODEN ZEPYAA. ISSN 0044-3328. URL

https://ui.adsabs.harvard.edu/#abs/1929ZPhy...55..738P.

Peierls:1929:ESZ

[Pei29b] R. Peierls. Über die Existenz stationärer Zustände. (German) [On the existence of stationary states]. Zeitschrift für Physik, 58(1–2): 59–62, January 1929. CODEN ZEPYAA. ISSN 0044-3328. URL

https://ui.adsabs.harvard.edu/#abs/1929ZPhy...58...59P.

Peierls:1929:KTWb

[Pei29c] R. Peierls. Zur kinetischen Theorie der Wärmeleitung in Kristallen. (German) [Kinetic theory of heat conduction in crystals]. Annalen der Physik (1900), 395(8):1055–1101, January 1929. ISSN 0003-3804 (print), 1521-3889 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1929AnP...395.1055P. This is Peierls’ doc-

toral dissertation.

Peierls:1929:THE

[Pei29d] R. Peierls. Zur Theorie des Hall-Effekts. (German) [The Hall Effect theory]. Physikalische Zeitschrift, 30:273–274, 1929. CODEN PHZTAO. ISSN 0369-982X.

Peierls:1929:TGE

[Pei29e] Rudolf Peierls. Zur Theorie der galvanomagnetischen Effekte. (German) [The theory of galvanomagnetic effects]. Zeitschrift für Physik, 53(3–4):255–266, March 1929. CODEN ZEPYAA.

ISSN 0044-3328. URL https://ui.adsabs.harvard.edu/#abs/
1929ZPhy...53..255P.

Peierls:1929:KTWa

[Pei29f] Rudolf Ernst Peierls. Zur kinetischen Theorie der Wärmeleitung. (German) [Kinetic theory of heat conduction]. D.Phil. dissertation, Universität Leipzig, Leipzig, Germany, 1929.

Peierls:1930:TET

[Pei30a] R. Peierls. Zur Theorie der elektrischen und thermischen Leitfähigkeit von Metallen. (German) [The theory of electrical and thermal conductivity of metals]. Annalen der Physik (1900), 396(2): 121–148, January 1930. ISSN 0003-3804 (print), 1521-3889 (elec-

tronic). URL https://ui.adsabs.harvard.edu/#abs/1930AnP.
..396..121P.

Peierls:1930:ZBT

[Pei30b] R. Peierls. Zwei Bemerkungen zur Theorie der Leitfähigkeit. (German) [Two remarks on the theory of conductivity]. Annalen der Physik (1900), 397(2):244–246, January 1930. ISSN 0003-3804 (print), 1521-3889 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1930AnP...397..244P.

Peierls:1930:VML

[Pei30c] Rudolf Peierls. Das Verhalten metallischer Leiter in starken Magnetfeldern. (German) [The behavior of metallic conductors in strong magnetic fields]. In Leipziger Vortrage: Elektronen-Interferenzen, pages 75–85. S. Hirzel, Leipzig, Germany, 1930.

Peierls:1931:TMW

[Pei31] R. Peierls. Zur Theorie der magnetischen Widerstandsänderung. (German) [The theory of magnetic resistance change]. Annalen der Physik (1900), 402(1):97–110, January 1931. ISSN 0003-3804 (print), 1521-3889 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1931AnP...402...97P.

Peierls:1932:EMG

[Pei32a] R. Peierls. Elektronentheorie der Metalle. (German) [Electron theory of metals]. Ergebnisse der exakten Naturwissenschaften, 11: 264–322, January 1932. CODEN EENAA3. ISSN 0367-0325. URL

https://ui.adsabs.harvard.edu/#abs/1932ErNW...11..264P.

Peierls:1932:AFK

[Pei32b] R. Peierls. Über die Absorptionsspektren fester Körper. (German) [On the absorption spectra of solid bodies]. Physikalische Zeitschrift der Sowjetunion, 1:297–298, 1932. CODEN PHZSAL. ISSN 0369-9811.

Peierls:1932:FEW

[Pei32c] R. Peierls. Zur Frage des elektrischen Widerstandsgesetzes für tiefe Temperaturen. (German) [On the question of the electrical resistance law for low temperatures]. Annalen der Physik (1900), 404(2): 154–168, January 1932. ISSN 0003-3804 (print), 1521-3889 (elec-

tronic). URL https://ui.adsabs.harvard.edu/#abs/1932AnP..
.404..154P.

Peierls:1932:TAF

[Pei32d] R. Peierls. Zur Theorie der Absorptionsspektren fester Körper. (German) [On the theory of absorption spectra of solid bodies]. Annalen der Physik (1900), 405(8):905–952, January 1932. ISSN 0003-3804 (print), 1521-3889 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1932AnP...405..905P.

Peierls:1932:TDL

[Pei32e] R. Peierls. Zur Theorie des Diamagnetismus von Leitungselektronen. (German) [The theory of diamagnetism of conduction electrons]. Physikalische Zeitschrift, 33:864–??, 1932. CODEN PHZ-TAO. ISSN 0369-982X.

Peierls:1933:TME

[Pei33a] R. Peierls. Zur Theorie der Metalle. Erwiderung auf eine Arbeit von A. H. Wilson. (German) [The theory of metals. Response to a work by A. H. Wilson]. Zeitschrift für Physik, 81(9–10):697–699, September 1933. CODEN ZEPYAA. ISSN 0044-3328. URL https:

//ui.adsabs.harvard.edu/#abs/1933ZPhy...81..697P.

Peierls:1933:TDLa

[Pei33b] R. Peierls. Zur Theorie des Diamagnetismus von Leitungselektronen. (German) [The theory of diamagnetism of conduction electrons]. Zeitschrift für Physik, 80(11–12):763–791, November 1933. CODEN ZEPYAA. ISSN 0044-3328. URL https://ui.adsabs.

harvard.edu/#abs/1933ZPhy...80..763P.

Peierls:1933:TDLb

[Pei33c] R. Peierls. Zur Theorie des Diamagnetismus von Leitungselektronen. II. Starke Magnetfelder. (German) [The theory of diamagnetism of conduction electrons. II. Strong magnetic fields]. Zeitschrift für Physik, 81(3–4):186–194, March 1933. CODEN

ZEPYAA. ISSN 0044-3328. URL https://ui.adsabs.harvard.
edu/#abs/1933ZPhy...81..186P.

Peierls:1934:BTM

[Pei34a] R. Peierls. Bemerkungen zur Theorie der Metalle. Antwort auf die Kritik von E. Kretschmann. (German) [Comments on the theory of metals. Answer to the criticism of E. Kretschmann]. Zeitschrift für Physik, 88(11–12):786–791, November 1934. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic).

Peierls:1934:ECL

[Pei34b] R. Peierls. Electric conductivity at low temperatures. Physikalische Zeitschrift der Sowjetunion, 6:516–519, 1934. CODEN PHZSAL. ISSN 0369-9811. Remark to a paper by S. V. Vonsovsky and A. A. Smirnov.

Peierls:1934:VDT

[Pei34c] R. Peierls. The vacuum in Dirac’s theory of the positive electron. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 146 (857):420–441, September 1, 1934. ISSN 0950-1207 (print), 2053-

9150 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1934RSPSA.146..420P.

Peierls:1934:BUG

[Pei34d] Rudolf Peierls. Bemerkungen über Umwandlungstemperaturen. (German) [Remarks on transition temperatures]. Helvetica Physica Acta, 7(Supplement 2):81–83, ???? 1934. CODEN HPACAK. ISSN 0018-0238.

Peierls:1934:SGE

[Pei34e] Rudolf Peierls. Über die statistischen Grundlagen der Elektronentheorie der Metalle. (German) [On the statistical foundations of the electron theory of metals]. Helvetica Physica Acta, 7 (Supplement 2):24–30, ???? 1934. CODEN HPACAK. ISSN

0018-0238. URL https://www.e-periodica.ch/digbib/view?
pid=hpa-001:1934:7::918#918.

Peierls:1935:DLL

[Pei35a] R. Peierls. Die diracsche Löchertheorie und die Lichtgeschwindigkeit im Vakuum. (German) [Dirac’s hole theory and the speed of light in a vacuum]. Physica, 2(1–12):399–402, January 1935. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic). URL

https://ui.adsabs.harvard.edu/#abs/1935Phy.....2..399P.

Peierls:1935:FPQ

[Pei35b] R. Peierls. The fundamental paradox of the quantum theory. Nature, 136(3436):395, September 7, 1935. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1935Natur.136..395P.

Peierls:1935:QTP

[Pei35c] R. Peierls. The quantum theory: Part 1. Quantum theory of atomic nuclei. Reports on Progress in Physics, 2(1):27–38, January 1935. CODEN RPPHAG. ISSN 0034-4885 (print), 1361-6633 (elec-

tronic). URL https://ui.adsabs.harvard.edu/#abs/1935RPPh.
...2...27P.

Peierls:1935:QPT

[Pei35d] R. Peierls. Quelques propriétés typiques des corps solides. (French) [Some typical properties of solid bodies]. Annales de l’Institut Henri Poincaré, 5(3):177–222, 1935. CODEN AIHPA2. ISSN 0365-320x (print), 2400-4855 (electronic). URL http://www.numdam.org/

item?id=AIHP_1935__5_3_177_0.

Peierls:1935:SEC

[Pei35e] R. Peierls. Statistical error in counting experiments. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 149(868):467–486, April 1935. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1935RSPSA.149..467P.

Peierls:1936:MTC

[Pei36a] R. Peierls. Magnetic transition curves of supraconductors. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 155(886):613–628, July 1936. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1936RSPSA.155..613P.

Peierls:1936:NDE

[Pei36b] R. Peierls. Note on the derivation of the equation of state for a degenerate relativistic gas. Monthly Notices of the Royal Astronomical Society, 96(8):780–784, June 1, 1936. CODEN MNRAA4. ISSN

0035-8711; 1365-2966/e. URL https://ui.adsabs.harvard.edu/
#abs/1936MNRAS..96..780P.

Peierls:1936:IMF

[Pei36c] R. Peierls. On Ising’s model of ferromagnetism. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 32(3):477–481, October 1936. CODEN PCPSA4. ISSN 0008-

1981. URL https://ui.adsabs.harvard.edu/#abs/1936PCPS..
.32..477P.

Peierls:1936:STA

[Pei36d] R. Peierls. Statistical theory of adsorption with interaction between the adsorbed atoms. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 32(3):471–476, October 1936. CODEN PCPSA4. ISSN 0008-1981. URL https://ui.

adsabs.harvard.edu/#abs/1936PCPS...32..471P.

Peierls:1936:STS

[Pei36e] R. Peierls. Statistical theory of superlattices with unequal concentrations of the components. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 154(881): 207–222, March 1936. CODEN PRLAAZ. ISSN 0080-4630 (print),

2053-9169 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1936RSPSA.154..207P.

Peierls:1936:ISE

[Pei36f] Rudolf Peierls. Interpretation of Shankland’s experiment. Nature, 137(3474):904, May 30, 1936. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1936Natur.137..904P.

Peierls:1937:CDT

[Pei37a] Rudolf Peierls. Contribution to ‘A discussion on β-type of nuclear transmutations’. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, A161(907):452–456, August 20, 1937. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). See also full paper [ECPR37].

Peierls:1937:PST

[Pei37b] Rudolf Peierls. The present state of the theory of β-decay. Izv. Akad. Nauk SSSR, ??(??):583–606, ???? 1937.

Peierls:1938:MPF

[Pei38] R. Peierls. On a minimum property of the free energy. Physical Review (2), 54(11):918–919, December 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1938PhRv...54..918P.

Peierls:1939:Ma

[Pei39a] R. Peierls. The meson. Reports on Progress in Physics, 6(1):78– 94, January 1939. CODEN RPPHAG. ISSN 0034-4885 (print),

1361-6633 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1939RPPh....6...78P.

Peierls:1939:CCN

[Pei39b] Rudolf Peierls. Critical conditions in neutron multiplication. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 35(4):610–615, November 1939. CODEN

PCPSA4. ISSN 0008-1981. URL https://ui.adsabs.harvard.
edu/#abs/1939PCPS...35..610P. This is the first paper with an

estimate of the critical mass of uranium needed to sustain a chain reaction. It contains a note added in proof to the fission model of Bohr and Wheeler [BW39] that appeared after this paper was submitted.

Peierls:1939:Mb

[Pei39c] Rudolf Peierls. The mesotron. Annual Reports on the Progress of Chemistry, 35(??):16–21, ???? 1939. CODEN ARPCAW. ISSN 0365-6217 (print), 1754-7512 (electronic).

Peierls:1940:SD

[Pei40a] R. Peierls. The size of a dislocation. Proceedings of the Physical Society, London, 52(1):34–37, January 1940. CODEN PPSOAU.

ISSN 0959-5309 (print), 2051-2171 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1940PPS....52...34P. See [Nab47,

PE17, ZLZ19].

Peierls:1940:TNFa

[Pei40b] R. Peierls. The theory of nuclear forces. Nature, 145(3679):687– 690, May 4, 1940. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1940Natur.145..687P.

Peierls:1940:ALS

[Pei40c] Rudolf Peierls. Admissible leaks in Simon plant. Report MS-23 (PRO AB4/852), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:AML

[Pei40d] Rudolf Peierls. Arrangement of membranes in layers. Report MS-18 (PRO AB4/846), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:BTN

[Pei40e] Rudolf Peierls. The Bohr theory of nuclear reactions. Reports on Progress in Physics, 7(1):87–106, January 1940. CODEN RP-PHAG. ISSN 0034-4885 (print), 1361-6633 (electronic). URL

http://iopscience.iop.org/0034-4885/7/1/306;   https://
ui.adsabs.harvard.edu/#abs/1940RPPh....7...87P.

Peierls:1940:CRCa

[Pei40f] Rudolf Peierls. Considerations regarding chain reactions. Report MS-1, Directorate of Tube Alloys, ????, UK, September 1940. ??

pp. URL https://pubarchive.lbl.gov/islandora/object/ir:
146363/datastream/PDF/download/citation.pdf.

Peierls:1940:CRCb

[Pei40g] Rudolf Peierls. Considerations regarding chain reactions. Report MS-2, Directorate of Tube Alloys, ????, UK, September 1940. ??

pp. URL https://pubarchive.lbl.gov/islandora/object/ir:
146363/datastream/PDF/download/citation.pdf.

Peierls:1940:CRCc

[Pei40h] Rudolf Peierls. Considerations regarding chain reactions. Report MS-3, Directorate of Tube Alloys, ????, UK, September 1940. ??

pp. URL https://pubarchive.lbl.gov/islandora/object/ir:
146363/datastream/PDF/download/citation.pdf.

Peierls:1940:CRCd

[Pei40i] Rudolf Peierls. Considerations regarding chain reactions. Report MS-4, Directorate of Tube Alloys, ????, UK, September 1940. ??

pp. URL https://pubarchive.lbl.gov/islandora/object/ir:
146363/datastream/PDF/download/citation.pdf.

Peierls:1940:DSDa

[Pei40j] Rudolf Peierls. Discussion of a specific design for diffusion apparatus. Report MS-14 (PRO AB4/841), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:DSDb

[Pei40k] Rudolf Peierls. Discussion of a specific design for diffusion apparatus. Report MS-14A (PRO AB4/842), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:ESL

[Pei40l] Rudolf Peierls. Effect of a stationary layer of thickness b in front of the diffusing membrane. Report MS-15 (PRO AB4/843), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:EES

[Pei40m] Rudolf Peierls. The effect of expansion of the sphere. Report MS-8 (PRO AB4/835B), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:EFNa

[Pei40n] Rudolf Peierls. The effect of fast neutrons in ordinary uranium. Report MS-4 (PRO AB4/831), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:EPCa

[Pei40o] Rudolf Peierls. Effect of packing on critical size. Report MS-7 (PRO AB4/834), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:EPCb

[Pei40p] Rudolf Peierls. Effect of packing on the critical radius of the sphere. Report MS-7A (PRO AB4/835A), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:EISb

[Pei40q] Rudolf Peierls. Efficiency of isotope separation. Report MS-12 (PRO AB4/837), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:EISa

[Pei40r] Rudolf Peierls. Efficiency of isotope separation by thermal diffusion. Report MS-11 (PRO AB4/836), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:EM

[Pei40s] Rudolf Peierls. Electrolytic method. Report MS-26 (PRO AB4/855), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:ETS

[Pei40t] Rudolf Peierls. Equilibrium time in a separation plant. Report MS-13 (PRO AB4/840), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:FAJ

[Pei40u] Rudolf Peierls. Finite amplitude in the “Jitterbug”. Report MS-27 (PRO AB4/856), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:FWM

[Pei40v] Rudolf Peierls. Finite width of membrane pair. Report MS-28 (PRO AB4/857), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:FWS

[Pei40w] Rudolf Peierls. Finite width of single membrane. Report MS-28A (PRO AB4/858), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:JPI

[Pei40x] Rudolf Peierls. The “Jitterbug” — part II. Report MS-19 (PRO AB4/848), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:MP

[Pei40y] Rudolf Peierls. Membrane pairs. Report MS-18A (PRO AB4/847), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:CSU

[Pei40z] Rudolf Peierls. On the critical size of a U235 sphere. Report MS-6 (PRO AB4/833), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:PSN

[Pei40-27] Rudolf Peierls. On the possibility of a slow-neutron chain reaction. Report MS-1 (PRO AB4/828), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:PMP

[Pei40-28] Rudolf Peierls. Packing of membrane pairs. Report MS-22 (PRO AB4/851), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:REU

[Pei40-29] Rudolf Peierls. Radioactive effects of uranium chain reaction. Report MS-3 (PRO AB4/830), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:SFD

[Pei40-30] Rudolf Peierls. A self-fractionating diffusion apparatus. Report MS-16 (PRO AB4/844), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:SEP

[Pei40-31] Rudolf Peierls. Separating effect of passage through capillary. Report MS-17 (PRO AB4/845), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:SMT

[Pei40-32] Rudolf Peierls. Separation of a mixture of three components. Report MS-20 (PRO AB4/849), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:SRC

[Pei40-33] Rudolf Peierls. Speed of reaction in a chain reaction with slow neutrons. Report MS-2 (PRO AB4/829), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:SFT

[Pei40-34] Rudolf Peierls. Streamline flow through a pipe of rectangular cross section. Report MS-24 (PRO AB4/853), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:SIU

[Pei40-35] Rudolf Peierls. Summary of information on uranium problem. Report MS-12B (PRO AB4/839), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1940:TNFb

[Pei40-36] Rudolf Peierls. Theory of nuclear forces. Annual Reports on the Progress of Chemistry, 36(??):24–32, ???? 1940. CODEN ARP-CAW. ISSN 0365-6217 (print), 1754-7512 (electronic).

Peierls:1940:VEB

[Pei40-37] Rudolf Peierls. Variation of efficiency with branching ratio. Report MS-25 (PRO AB4/854), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1941:ASP

[Pei41a] Rudolf Peierls. Assembly of a separation plant. Report MS-29A (PRO AB4/862), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:APSa

[Pei41b] Rudolf Peierls. Assembly problems in Simon plant. Report MS-29 (PRO AB4/860), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:APSb

[Pei41c] Rudolf Peierls. Assembly problems: size of recovery plant. Report MS-29 (PRO AB4/861), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:CCS

[Pei41d] Rudolf Peierls. Cascade of cascades in Simon plant: results. Report MS-30 (PRO AB4/863), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:CDD

[Pei41e] Rudolf Peierls. Clusius–Dickel device as a gland. Report MS-39 (PRO AB4/873), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:CCT

[Pei41f] Rudolf Peierls. Critical conditions and time constant in chain reaction. Report MS-41 (PRO AB4/875), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:EMP

[Pei41g] Rudolf Peierls. Efficiency of membrane pair for small pressure drop. Report MS-45 (PRO AB4/879), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:CSMa

[Pei41h] Rudolf Peierls. Cul de Sac method — II. Report MS-40 (PRO AB4/874), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:CSMb

[Pei41i] Rudolf Peierls. Cul de Sac method — III. Report MS-42 (PRO AB4/876), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:EFS

[Pei41j] Rudolf Peierls. Equation of flow in Simon plant and performance factor. Report MS-31 (PRO AB4/864), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:ETC

[Pei41k] Rudolf Peierls. Equilibrium time in composite Simon plant. Report MS-34 (PRO AB4/867), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:ETS

[Pei41l] Rudolf Peierls. Equilibrium time in separation plant. Report MS-47 (PRO AB4/841), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:IHS

[Pei41m] Rudolf Peierls. Influence of hole spacing in the thermo-molecular device. Report MS-33 (PRO AB4/866), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:IMP

[Pei41n] Rudolf Peierls. Influence of membrane properties. Report MS-37 (PRO AB4/871), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:ISM

[Pei41o] Rudolf Peierls. Isotope separation by means of the thermomolecular effect. Report MS-32 (PRO AB4/865), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:OSS

[Pei41p] Rudolf Peierls. Optimum size of separation plant. Report BR-98 (PRO AB4/99), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:SCS

[Pei41q] Rudolf Peierls. Separation by cul de sac method. Report MS-38 (PRO AB4/872), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:SME

[Pei41r] Rudolf Peierls. Separation by means of evaporation at low pressures. Report MS-46 (PRO AB4/880), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:SPSa

[Pei41s] Rudolf Peierls. Simon plant: separation of components. Report MS-36 (PRO AB4/869), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:SPSb

[Pei41t] Rudolf Peierls. Simon plant: separation of components. Report MS-36A (PRO AB4/870), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:SSS

[Pei41u] Rudolf Peierls. Small scale separation with membrane pairs. Report MS-43 (PRO AB4/877), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1941:TCC

[Pei41v] Rudolf Peierls. Theory of a certain class of self-fractionating apparatus. Report MS-35 (PRO AB4/868), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1942:AD

[Pei42a] Rudolf Peierls. The “accordion” device. Report MS-59 (PRO AB4/895), Directorate of Tube Alloys, ????, UK, 1942.

Peierls:1942:ASP

[Pei42b] Rudolf Peierls. Assembly of separation plant. Report MS-48 (PRO AB4/883), Directorate of Tube Alloys, ????, UK, 1942.

Peierls:1942:AAE

[Pei42c] Rudolf Peierls. The asymptotic approach to equilibrium conditions. Report MS-54 (PRO AB4/890), Directorate of Tube Alloys, ????, UK, 1942.

Peierls:1942:CC

[Pei42d] Rudolf Peierls. Cascade of cascades. Report MS-57 (PRO AB4/893), Directorate of Tube Alloys, ????, UK, 1942.

Peierls:1942:CSP

[Pei42e] Rudolf Peierls. Cost of separation plant as function of pressure, membrane porosity, and width of membrane pair. Report MS-49 (PRO AB4/884), Directorate of Tube Alloys, ????, UK, 1942.

Peierls:1942:EDF

[Pei42f] Rudolf Peierls. Energy distribution of fission neutrons. Report MS-65 (PRO AB4/902), Directorate of Tube Alloys, ????, UK, 1942.

Peierls:1942:NSS

[Pei42g] Rudolf Peierls. Note on single-stage vs. double-stage compression. Report MS-58 (PRO AB4/894), Directorate of Tube Alloys, ????, UK, 1942.

Peierls:1942:NAP

[Pei42h] Rudolf Peierls. Notes on the attached paper (AB4/42) by Dirac. Report BR42a (PRO AB4/42a), Directorate of Tube Alloys, ????, UK, June 1942.

Peierls:1942:FG

[Pei42i] Rudolf Peierls. On flushing glands. Report MS-50 (PRO AB4/885), Directorate of Tube Alloys, ????, UK, March 1942.

Peierls:1942:FGIa

[Pei42j] Rudolf Peierls. On flushing glands. II. Report MS-52 (PRO AB4/888), Directorate of Tube Alloys, ????, UK, March 1942.

Peierls:1942:FGIb

[Pei42k] Rudolf Peierls. On flushing glands. III. Report MS-53 (PRO AB4/889), Directorate of Tube Alloys, ????, UK, March 1942.

Peierls:1942:PDT

[Pei42l] Rudolf Peierls. Pressure distribution in top of plant. Report MS-62 (PRO AB4/898), Directorate of Tube Alloys, ????, UK, 1942.

Peierls:1942:VPMb

[Pei42m] Rudolf Peierls. Variation of pressure and membrane area — second supplement to MS. 56. Report MS-68 (PRO AB4/906), Directorate of Tube Alloys, ????, UK, 1942.

Peierls:1942:VPMa

[Pei42n] Rudolf Peierls. Variation of pressure and membrane area in diffusion plant. Report MS-56 (PRO AB4/892), Directorate of Tube Alloys, ????, UK, 1942.

Peierls:1943:CRC

[Pei43a] Rudolf Peierls. Comments on report A-387 (CF338) by Feld and Szilard. Report MS-75 (PRO AB4/913), Directorate of Tube Alloys, ????, UK, 1943.

Peierls:1943:ELD

[Pei43b] Rudolf Peierls. Efficiency loss due to the lateral variation of the branching ratio for a non-rectangular membrane pair. Report MS-110 (PRO AB4/949), Directorate of Tube Alloys, ????, UK, 1943.

Peierls:1943:GPP

[Pei43c] Rudolf Peierls. Gland problems — Part II. Report MS-70 (PRO AB4/908), Directorate of Tube Alloys, ????, UK, 1943.

Peierls:1943:MPC

[Pei43d] Rudolf Peierls. Membrane pair characteristics. Report MS-80 (PRO AB4/918), Directorate of Tube Alloys, ????, UK, 1943.

Peierls:1944:LDM

[Pei44a] Rudolf Peierls. Longitudinal diffusion in membrane pairs. Report MS-86 (PRO AB4/925), Directorate of Tube Alloys, ????, UK, 1944.

Peierls:1944:NCB

[Pei44b] Rudolf Peierls. Notes on composite barriers. Report MS-N7 (PRO AB4/970), Directorate of Tube Alloys, ????, UK, 1944.

Peierls:1944:OFI

[Pei44c] Rudolf Peierls. Output fluctuations in an idealised plant. Report MS-N6 (PRO AB4/969), Directorate of Tube Alloys, ????, UK, March 22, 1944.

Peierls:1945:VBP

[Pei45] Rudolf Peierls. Variation in barrier porosity. Report BI-85 (PRO AB4/802), Directorate of Tube Alloys, ????, UK, 1945.

Peierls:1946:LDA

[Pei46a] R. E. Peierls. Letter: Defence against the atom bomb. Nature, 158 (4011):379, September 14, 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1946Natur.158..379P.

Peierls:1946:UET

[Pei46b] R. E. Peierls. Use of the electrolytic tank for magnetic problems. Nature, 158(4023):831–832, December 7, 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1946Natur.158..831P.

Peierls:1946:FP

[Pei46c] Prof. R. E. Peierls, C.B.E., F.R.S. Fundamental particles. Nature, 158(4022):773–775, November 30, 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1946Natur.158..773P.

Peierls:1947:BRF

[Pei47a] R. E. Peierls. Book review: Fisica nucleare: La scissione nucleare dell’uranio. Nature, 159(4029):79–80, January 18, 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

https://ui.adsabs.harvard.edu/#abs/1947Natur.159...79P.

Peierls:1947:WEN

[Pei47b] R. E. Peierls. What experiments are needed in fundamental physics? Proceedings of the Physical Society, London, 59(3):412– 417, May 1947. CODEN PPSOAU. ISSN 0959-5309 (print), 2051-

2171 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1947PPS....59..412P.

Peierls:1947:AET

[Pei47c] Rudolf Peierls. Atomic energy — threat and promise. Endeavour, 6 (22):51–58, ???? 1947. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Peierls:1947:PAF

[Pei47d] Rudolf Peierls. Production of atomic fuel by isotope separation. Science News (Washington, DC), 2(??):37–62, ???? 1947. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic).

Peierls:1947:RAE

[Pei47e] Rudolf Peierls. The release of atomic energy. Science News (Washington, DC), 2(??):1–18, ???? 1947. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic).

Peierls:1948:ETS

[Pei48a] R. Peierls. Expansions in terms of sets of functions with complex eigenvalues. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 44(2):242–250, January 1948. CODEN PCPSA4. ISSN 0008-1981. URL https://ui.adsabs.

harvard.edu/#abs/1948PCPS...44..242P.

Peierls:1948:EMT

[Pei48b] R. E. Peierls. Electrons in modern theoretical physics. Journal of the Institution of Electrical Engineers — Part I: General, 95(87): 131–132, March 1948. CODEN IEJGAB. ISSN 0367-7559 (print), 2054-0582 (electronic).

Peierls:1948:LNA

[Pei48c] Rudolf Peierls. Living in the nuclear age. BBC Web site for schools.,

1948. URL http://www.bbc.co.uk/schoolradio/subjects/
history/britainsince1930s/technology/nuclear_age.

Peierls:1948:PNP

[Pei48d] Rudolf Peierls. Problems of nuclear physics. Science News (Washington, DC), 6(??):62–79, ???? 1948. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic).

Peierls:1950:P

[Pei50a] Rudolf Peierls. I am not perturbed. Sunday Express, ??(??):??, October 29, 1950.

Peierls:1950:STB

[Pei50b] Rudolf Peierls. The scientific and technical backgrounds. In Crammer and Peierls [CP50], pages 29–36. LCCN TK9145 .C75. Revised and expanded version of Science News volume 2 (February 1947).

Peierls:1950:BAS

[Pei50c] Rudolf E. Peierls. The British Atomic Scientists’ Association. Bulletin of the Atomic Scientists, 6(2):59, February 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1951:BRL

[Pei51] R. E. Peierls. Book review: Lectures on theoretical physics. Vorlesungen über theoretische Physik. Von Prof. Arnold Sommerfeld. Band 1. Mechanik. Vierte, neubearbeitete Auflage. Pp. xii + 276. Band 2: Mechanik der derformierbaren Medien. Pp. xv + 376 + 4 plates. Band 3: Elektrodynamik. Pp. xvi + 36. Band 6: Partielle Differentialgleichungen der Physik. Pp. xiii + 332 (Wiesbaden: Dieterich’sche Verlagsbuchhandlung, 1947–1949). Nature, 168(4282):887, November 24, 1951. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1951Natur.168Q.887P.

Peierls:1952:BRV

[Pei52a] R. E. Peierls. Book review: Vorlesungen über theoretische Physik. Band 4. Optik. Von Prof. Arnold Sommerfeld. Pp. xvi + 390 (Wiesbaden: Dieterich’sche Verlagsbuchhandlung Inh. W. Klemm, G.m.b.H., 1950). Nature, 170(4330):684, October 25, 1952. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

https://ui.adsabs.harvard.edu/#abs/1952Natur.170Q.684P.

Peierls:1952:CLRa

[Pei52b] R. E. Peierls. Commutation laws of relativistic field theory. In Report of an International Conference on Elementary Particles, Bombay, 1950, pages 11–14; discussion, 15–16. The International Union of Pure and Applied Physics, Bombay, India, 1952.

Peierls:1952:CLRb

[Pei52c] R. E. Peierls. The commutation laws of relativistic field theory. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 214(1117):143–157, August 21, 1952. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-

9169 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1952RSPSA.214..143P.

Peierls:1952:SBE

[Pei52d] R. E. Peierls. Some British experiences. Bulletin of the Atomic Scientists, 8(7):229–230, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders.

Peierls:1952:EAE

[Pei52e] Rudolf Peierls. An enquiry about entry to mathematics courses. Mathematical Gazette, 36(315):33–35, February 1952. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

http://www.jstor.org/stable/3610765.

Peierls:1953:RLA

[Pei53a] R. E. Peierls. 6th Rutherford Lecture: The atomic nucleus and its constituents. Proceedings of the Physical Society A, 66(4):313–

324, April 1953. URL https://ui.adsabs.harvard.edu/#abs/
1953PPSA...66..313P.

Peierls:1953:GEB

[Pei53b] R. E. Peierls. Guest editorial: Basic science and the Cold War. Bulletin of the Atomic Scientists, 9(3):66–67, April 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1953:FTN

[Pei53c] Rudolf Peierls. Field theories with non-local interactions. In Proc. Int. Conf. on Theor. Phys. at Kyoto and Tokyo 1953, pages 24–39. ????, ????, 1953.

Peierls:1953:DM

[Pei53d] Rudolf Peierls. On diamagnetism of metals. In Proc. Int. Conf. on Theor. Phys. at Kyoto and Tokyo 1953, pages 676–678. ????, ????, 1953.

Peierls:1953:RCB

[Pei53e] Rudolf Peierls. Report on the conference at Birmingham. In Proc. Int. Conf. on Theor. Phys. at Kyoto and Tokyo 1953, pages 301– 310. ????, ????, 1953.

Peierls:1954:SRM

[Pei54a] R. E. Peierls. Short review: Mechanics. (Lectures on Theoretical Physics, Vol. 1). By Arnold Sommerfeld. Translated from the fourth German edition by Martin O. Stern. Pp. xiv + 289. (New York: Academic Press, Inc.; London: Academic Books, Ltd., 1952). Nature, 173(4411):932, May 15, 1954. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1954Natur.173..932P.

Peierls:1954:NVS

[Pei54b] Rudolf Peierls. Note on the vibration spectrum of a crystal. Proceedings of the National Academy of Sciences, India, 20(??):121– 126, 1954.

Peierls:1954:FTN

[Pei54c] Rudolf E. Peierls. Field theories with non-local interactions. In Proceedings of the international conference of theoretical physics, Kyoto and Tokyo, 1953, pages 24–29; discussion, pp. 29–39. Science Council of Japan, Tokyo, 1954.

Peierls:1955:BRO

[Pei55a] R. E. Peierls. Book review: Optics, by Arnold Sommerfeld. (Lectures on Theoretical Physics, Volume 4.) Translated by Otto Laporte and Peter A. Moldauer. Pp. xiii + 383 (New York: Academic Press, Inc.; London: Academic Books, Ltd. 1954). Nature, 176 (4488):850, November 5, 1955. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1955Natur.176..850P.

Peierls:1955:LN

[Pei55b] R. E. Peierls. The Laws of Nature. George Allan and Unwin, Ltd., London, UK, 1955. 284 pp.

Peierls:1955:ODG

[Pei55c] R. E. Peierls. Obituary: Dr. G. Placzek. Nature, 176(4494): 1149–1150, December 17, 1955. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1955Natur.176R1149P.

Peierls:1955:QTS

[Pei55d] R. E. Peierls. Quantum Theory of Solids. International Series of Monographs on Physics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1955. ???? pp.

Peierls:1955:SFTb

[Pei55e] R. E. Peierls. A survey of field theory. I. General background. Reports on Progress in Physics, 18:424–432, January 1955. CODEN RPPHAG. ISSN 0034-4885 (print), 1361-6633 (electronic).

Peierls:1955:SFTc

[Pei55f] R. E. Peierls. A survey of field theory. VI. Final remarks. The new particles. Reports on Progress in Physics, 18:470–477, January 1955. CODEN RPPHAG. ISSN 0034-4885 (print), 1361-6633 (electronic).

Peierls:1955:AE

[Pei55g] Rudolf Peierls. Albert Einstein. In Physical Society Year Book 1955, pages 69–71. ????, ????, 1955.

Peierls:1955:FP

[Pei55h] Rudolf Peierls. The frontier of physics. In The Twentieth Century, volume 48, pages 204–213. ????, ????, 1955.

Peierls:1955:MJR

[Pei55i] Rudolf Peierls. In the matter of J. Robert Oppenheimer. Atomic Scientists’ Journal, 4(??):145–154, ???? 1955.

Peierls:1955:IPP

[Pei55j] Rudolf Peierls. Interpretation and properties of propagators. In E. H. Bellamy and R. G. Moorhouse, editors, Proceedings of the International Conference on Nuclear and Meson Physics at Glasgow 1954, pages 296–299. Pergamon Press, New York, NY, USA, 1955.

Peierls:1955:ABB

[Pei55k] Rudolf E. Peierls. The atom that Bohr built. Journal of Jocular Physics, III:47–48, October 7, 1955.

Peierls:1956:BRT

[Pei56a] R. E. Peierls. Book review: Thermodynamics and Statistical Mechanics. By Arnold Sommerfeld. (Lectures on Theoretical Physics, Vol. 5). Translated by J. Kestin. Edited by F. Bopp and J. Meixner. Pp. xviii + 401. (New York: Academic Press, Inc.; London: Academic Books, Ltd., 1956). Nature, 178(4528):288, August 11, 1956. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1956Natur.178..288P.

Peierls:1956:RDN

[Pei56b] R. E. Peierls. Recent developments in nuclear theory. British Journal of Applied Physics, 7(Supplement 5):S42, January 1956. CODEN BJAPAJ. ISSN 0508-3443 (print), 2057-7656 (elec-

tronic). URL http://stacks.iop.org/0508-3443/7/i=S5/a=
307;   https://ui.adsabs.harvard.edu/#abs/1956BJAP....
7S..42P.

Peierls:1956:SSP

[Pei56c] R. E. Peierls. Summer School in Physics in the University of Mexico. Nature, 178(4544):1218–1219, December 1, 1956. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

https://ui.adsabs.harvard.edu/#abs/1956Natur.178.1218P.

Peierls:1957:BRF

[Pei57a] R. E. Peierls. Book review: Field theory and meson practice. Mesons and Fields, Vol. 1. Fields. By Prof. Silvan Schweber, Prof. Hans. A. Bethe, and Frederic de Hoffmann. Pp. xvi + 449. Vol 2. Mesons. By Prof. Hans. A. Bethe and Frederic de Hoffmann. Pp. xiii + 446. (Evanston, Ill.: Row, Peterson and Company, 1955). Nature, 179(4562):711–712, April 6, 1957. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1957Natur.179..711P.

Peierls:1957:BRQ

[Pei57b] R. E. Peierls. Book review: Quantum theory of fields. Quantum Field Theory. By Prof. H. Umezawa. (Series in Physics). Pp. xv + 364 (Amsterdam: North-Holland Publishing Company; New York: Interscience Publishers, Inc. 1956). Nature, 179(4563):747– 748, April 13, 1957. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1957Natur.179..747P.

Peierls:1958:CR

[Pei58a] R. E. Peierls. Concluding remarks. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 246 (1247):490–494, August 1958. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1958RSPSA.246..490P.

Peierls:1958:IDP

[Pei58b] R. E. Peierls. Introduction: [to a discussion on parity nonconservation]. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 246(1247):441–443, August 1958. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-

9169 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1958RSPSA.246..441P.

Peierls:1958:MEP

[Pei58c] R. E. Peierls. Measurement of electron polarization: Discussion. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 246(1247):470, August 26, 1958. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL

https://ui.adsabs.harvard.edu/#abs/1958RSPSA.246..470P.

Peierls:1959:AN

[Pei59a] R. E. Peierls. The atomic nucleus. Scientific American, 200 (1):75–84, January 1959. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v200/n1/pdf/scientificamerican0159-
75.pdf;   https://ui.adsabs.harvard.edu/#abs/1959SciAm.
200a..75P.

Peierls:1959:CES

[Pei59b] R. E. Peierls. Complex eigenvalues in scattering theory. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 253(1272):16–36, November 1959. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1959RSPSA.253...16P.

Peierls:1959:GSM

[Pei59c] R. E. Peierls. The Graduate School in Mathematical Physics at Birmingham. Nature, 184(4680):96, July 11, 1959. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https:

//ui.adsabs.harvard.edu/#abs/1959Natur.184...96P.

Peierls:1959:LN

[Pei59d] R. E. Peierls. The Laws of Nature. Verlag Franz Deuticke, Vienna, Austria, 1959. viii + 334 + 67 pp.

Peierls:1959:STN

[Pei59e] R. E. Peierls. Selected topics in nuclear theory. In W. E. Brittin and L. G. Dunham, editors, Lectures in theoretical physics, Vol. I. Lectures delivered at the Summer Institute for Theoretical Physics, University of Colorado, Boulder, 1958, pages 238–267, 292–314. Interscience Publishers, New York, NY, USA, 1959.

Peierls:1960:QTS

[Pei60a] R. E. Peierls. Quantum theory of solids. In Theoretical physics in the twentieth century (Pauli memorial volume), pages 140–160. Interscience, New York, 1960.

Peierls:1960:WEP

[Pei60b] R. E. Peierls. Wolfgang Ernst Pauli: 1900–1958. Biographical Memoirs of Fellows of the Royal Society, 5:174–192, February 1960. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

URL http://www.jstor.org/stable/769285.

Peierls:1960:OP

[Pei60c] Rudolf Peierls. Open problems. In D. A. Bromley and E. W. Vogt, editors, Proceedings of the International Conference on Nuclear Structure at Kingston, Ontario, Canada 1960, pages 7–19. University of Toronto Press, Toronto, ON, Canada, 1960.

Peierls:1961:PHW

[Pei61a] R. E. Peierls. Professor H. W. B. Skinner, F.R.S. Nuclear Physics, 23:1–4, February/March 1961. CODEN NUPHA7. ISSN 0029-5582 (print), 1873-1546 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1961NucPh..23....1P.

Peierls:1961:AM

[Pei61b] Rudolf Peierls. Agonising misappraisal. Spectator, ??(??):592–593, April 28, 1961.

Peierls:1961:VDN

[Pei61c] Rudolf Peierls. Velocity-dependent nuclear forces. In J. Birks, editor, Proceedings of the Rutherford Jubilee Conference at Manchester 1961, pages 739–741. Heywood, London, UK, 1961.

Peierls:1962:PNP

[Pei62a] Rudolf Peierls. Praemia, Nobel Physics Prize to L. D. Landau. Experienta, 18(??):1–2, ???? 1962.

Peierls:1962:RAL

[Pei62b] Rudolf Peierls. Response to award of the Lorentz Medal. Proceedings of the Koninklijke Nederlandse Akademie van Wetenschappen, Series A: Mathematical Sciences, ??(??):6–8, November 24, 1962. CODEN KNWAA7. ISSN 0023-3358. URL https://www.lorentz.

leidenuniv.nl/lorentzmedal/lorentzmedal1962/1962.pdf.

Peierls:1963:ANB

[Pei63a] R. E. Peierls. An appreciation of Niels Bohr. Proceedings of the Physical Society, London, 81(5):793–799, May 1, 1963. CODEN PPSOAU. ISSN 0959-5309 (print), 2051-2171 (electronic). URL

http://stacks.iop.org/0370-1328/81/i=5/a=301;   https:/
/ui.adsabs.harvard.edu/#abs/1963PPS....81..793P.

Peierls:1963:FTS

[Pei63b] Rudolf Peierls. Field theory since Maxwell. In Clerk Maxwell and modern science; six commemorative lectures [Dom63], pages 26–42. LCCN QC71 .D63.

Peierls:1963:NM

[Pei63c] Rudolf Peierls. Nuclear matter. Endeavour, 22(??):146–150, ???? 1963. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Peierls:1963:WSW

[Pei63d] Rudolf Peierls. Wo stehen wir in der Kenntniss der Naturgesetze?. (German) [Where do we now stand in our knowledge of the Laws of Physics?]. Physikalische Blätter, 19(??):533–539, ???? 1963. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Peierls:1964:DQT

[Pei64a] R. E. Peierls. The development of quantum theory. Part I. Formulation and interpretation. Contemporary Physics, 6(2):129–139, December 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-

5812 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1964ConPh...6..129P.

Peierls:1964:IRS

[Pei64b] R. E. Peierls. Introductory remarks on solid-state physics. In E. R. Caianiello, editor, Lectures on the Many-Body Problem, volume 2, pages 1–?? Academic Press, New York, NY, USA, 1964. URL

https://ui.adsabs.harvard.edu/#abs/1964mbp..conf....1P.

Peierls:1964:QTS

[Pei64c] R. E. Peierls. Quantum Theory of Solids. International Series of Monographs on Physics. Clarendon Press, Oxford, UK, 1964. viii + 229 pp.

Peierls:1964:GC

[Pei64d] Rudolf Peierls. General conclusions. In P. Gugenberger, editor, C. R. Congrès Intl. phys. nucl. at Paris, 2–8 July 1964, pages 503–510. Éditions C.N.R.S., Paris, France, 1964.

Peierls:1964:PS

[Pei64e] Rudolf Peierls. A physicist sums up. In Analysis in function space, pages 209–213. MIT Press, Cambridge, MA, USA, 1964.

Peierls:1964:PCS

[Pei64f] Rudolf Peierls. Problems of collective security. In ????, editor, Proceedings of the 13th Pugwash Conference: Disarmament & peaceful collaboration among nations, at Karlovy Vary, Sept. 13–19, 1964, pages 259–264. Taylor & Francis, London, UK, 1964. URL https:

//pugwash.org/pugwash-international-conferences/.

Peierls:1964:WNK

[Pei64g] Rudolf Peierls. Warmeleitung in nichtmetallischen Kristallen. (German) [Thermal conductivity in nonmetallic crystals]. In E. Brueche, editor, Hauptvortrage v. d. Tagung Verbandes Dtscher Physik. Ges. 1963, pages 1–7. Mosbach, Baden, West Germany, 1964.

Peierls:1965:DQT

[Pei65a] R. E. Peierls. The development of quantum theory. Part II. Consolidation and extension. Contemporary Physics, 6(3):192–205, February 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-

5812 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1965ConPh...6..192P.

Peierls:1965:ERC

[Pei65b] R. E. Peierls. Essay review: The concept of the positron. History of Science (UK), 4(1):124–129, March 1965. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/4/1/124.full.pdf+html;   https:
//ui.adsabs.harvard.edu/#abs/1965HisSc...4..124P.

Peierls:1965:QTS

[Pei65c] R. E. Peierls. Quantum Theory of Solids. International Series of Monographs on Physics. Clarendon Press, Oxford, UK, 1965. viii + 229 pp.

Peierls:1966:KGN

[Pei66a] Rudolf Peierls. Die Kernmaterie. (German) [Nuclear matter]. Physikalische Blätter, 22(2):57–65, February 1966. CODEN PH-BLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Peierls:1966:PUB

[Pei66b] Rudolf Peierls. A possible use of “black boxes” in connection with a comprehensive test ban. In ????, editor, Proceedings of the 16th Pugwash Conference: Disarmament & world security, especially in Europe, at Sopot, Poland, September 1966, page 272. Taylor &

Francis, London, UK, 1966. URL https://pugwash.org/pugwash-
international-conferences/.

Peierls:1967:OIP

[Pei67a] R. E. Peierls. Obituary: Isaac Pomeranchuk. Nature, 213(5073): 238, January 21, 1967. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1967Natur.213..238P.

Peierls:1967:BEQ

[Pei67b] Rudolf E. Peierls. Books: An examination of the quantum jumps in the growth of quantum mechanics: The Conceptual Development of Quantum Mechanics, by Max Jammer. McGraw-Hill Book Company. Scientific American, 216 (1):137–140, 142, 144–145, January 1967. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:/

/www.nature.com/scientificamerican/journal/v216/n1/pdf/
scientificamerican0167-137.pdf;   https://ui.adsabs.harvard.
edu/#abs/1967SciAm.216a.137P.

Peierls:1968:LDL

[Pei68a] R. E. Peierls. Lev D. Landau. Physics Bulletin, 19(5):147, May 1968. CODEN PHSBB4. ISSN 0031-9112.

Peierls:1968:CPN

[Pei68b] Rudolf Peierls. Commentary on the “Peierls–Nabarro Force”. In A. R. Rosenfield, G. T. Hahn, A. L. Bement, Jr., and R. I. Jaffee, editors, Dislocation Dynamics, (Battelle Institute, Materials Science Colloquium at Seattle, May 1–6, 1967), pages xiii–xiv, xviii–xx. McGraw-Hill, New York, NY, USA, 1968.

Peierls:1969:PP

[Pei69a] Rudolf Peierls. Physicists and politics. The Times Literary Supplement, ??(??):1442–??, December 18, 1969. ISSN 0040-7895. An unsigned review.

Peierls:1969:SPA

[Pei69b] Sir Rudolf E. Peierls. The scientist in public affairs: Between the ivory tower and the arena. Bulletin of the Atomic Scientists, 25(9): 28–30, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1970:BDN

[Pei70a] Rudolf Peierls. A ban on the development of new weapons? In ????, editor, Proceedings of the 19th Pugwash Conference: World Security, disarmament & development, at Sochi, USSR, Oct. 1969, pages 280–283. Taylor & Francis, London, UK, 1970. URL https:

//pugwash.org/pugwash-international-conferences/.

Peierls:1970:CS

[Pei70b] Rudolf Peierls. Closing summary. In J. S. C. McKee and P. M. Rolph, editors, Proceedings of the 1st International Conference on three-body problem in nuclear and particle physics at Birmingham 8–10 July 1969, pages 481–488. Elsevier, Amsterdam, The Netherlands, 1970.

Peierls:1970:TRS

[Pei70c] Rudolf Peierls. Time reversal and the Second Law of Thermodynamics. In J. Bowcock, editor, Conference on Methods and Problems of Theoretical Physics at University of Birmingham 1967, pages 3–10. Elsevier, Amsterdam, The Netherlands, 1970.

Peierls:1970:BAA

[Pei70d] Sir Rudolf E. Peierls. Britain in the Atomic Age. Bulletin of the Atomic Scientists, 26(6):40–46, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1971:AGH

[Pei71a] Rudolf Peierls. Atomic Germans [Heisenberg’s recollections]. New York Review of Books, ??(??):23–24, July 1, 1971. ISSN 0028-7504 (print), 1944-7744 (electronic).

Peierls:1971:IAP

[Pei71b] Rudolf Peierls. International aspects of the pollution problem. In ????, editor, Proceedings of the 21st Pugwash Conference: Problems of world security, environment & development, at Sinaia, Romania, August 1971, pages 294–303. Taylor & Francis, London,

UK, 1971. URL https://pugwash.org/pugwash-international-
conferences/.

Peierls:1971:TCS

[Pei71c] Rudolf Peierls. Is there a crisis in science? Joseph Wunsch Lecture, Technion, Haifa, Israel., 1971.

Peierls:1971:SIB

[Pei71d] Rudolf Peierls. Strong interactions [Born–Einstein correspondence]. New Scientist, ??(??):339, May 6, 1971. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Peierls:1972:FDS

[Pei72] R. Peierls. Fermi–Dirac statistics. In Salam and Wigner [SW72], pages 117–127. ISBN 0-521-08600-0. LCCN QC174.1 .A85

1972. URL http://hooke.lib.cam.ac.uk/cgi-bin/bib_seek.
cgi?cat=ul&bib=1733506;   http://www.loc.gov/catdir/enhancements/
fy1001/72075298-d.html;   http://www.loc.gov/catdir/enhancements/
fy1001/72075298-t.html.

Peierls:1973:DQF

[Pei73a] Rudolf Peierls. The development of quantum field theory. In Mehra [Meh73], pages 370–379. ISBN 90-277-0345-0, 90-277-2536-5.

LCCN QC173.96 .S95 1972. URL http://www.springer.com/us/
book/9789027703453.

Peierls:1973:PTP

[Pei73b] Rudolf Peierls. Perturbation theory for projected states. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 333(1593):157–170, May 1973. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL

https://ui.adsabs.harvard.edu/#abs/1973RSPSA.333..157P.

Peierls:1974:QTS

[Pei74a] R. E. Peierls. Quantum Theory of Solids. International Series of Monographs on Physics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1974. ISBN 0-19-851240-6. viii + 229 pp.

Peierls:1974:JRO

[Pei74b] Rudolf Peierls. J. Robert Oppenheimer. In Charles Coulston Gillispie, editor, Dictionary of Scientific Biography: Volume 10: S. G. Navashin–W. Piso, pages 213–218. Charles Scribner’s Sons, New York, NY, USA, 1974. ISBN 0-684-10121-1. LCCN Q141D5 1970.

Peierls:1974:SSR

[Pei74c] Sir Rudolf Peierls. Some simple remarks on the basis of transport theory. In G. Kirczenow and J. Marro, editors, Transport Phenomena, Sitges International School of Statistical Mechanics, June 1984, volume 31 of Lecture Notes in Physics, pages 1–33. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. ISBN 3-540-06955-0. URL https://ui.

adsabs.harvard.edu/#abs/1974LNP....31....1P;   https://
ui.adsabs.harvard.edu/#abs/1974LNP....31....2P.

Peierls:1975:GDP

[Pei75] Rudolf Peierls. The glorious days of physics. In A. Zichichi, editor, Lepton and Hadron Structure, Proceedings of the 12th International School on Subnuclear Physics “Ettore Majorana”, Erice-Trapani, Sicily, 14–31 July 1974, pages 916–931. Academic Press, New York, NY, USA, 1975.

Peierls:1976:DS

[Pei76a] Rudolf Peierls. Decaying states. In John S. Risley and R. Geballe, editors, 9th International Conference on the Physics of Electronic and Atomic Collisions, IX ICPEAC. Invited lectures, review papers, and progress reports from the 9th International Conference, held in Seattle, Washington, July 24–30, 1975, pages 40–50. University of Washington Press, Seattle, WA, USA, January 1976. URL https:

//ui.adsabs.harvard.edu/#abs/1976peac.conf...40P.

Peierls:1976:TCP

[Pei76b] Rudolf Peierls. Twentieth-century physics. In Douglas Villiers, editor, New Year in Jerusalem: Jews in the Twentieth Century, pages 295–307. Douglas Villiers Publishing, London, UK, 1976.

Peierls:1976:MLR

[Pei76c] Sir Rudolf Peierls, F.R.S. The momentum of light in a refracting medium. I. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 347(1651):475–491, January 27, 1976. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-

9169 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1976RSPSA.347..475P.

Peierls:1977:I

[Pei77a] Rudolf Peierls. Introduction. In A Perspective of Physics, volume 1, pages ix–xxxv. Gordon & Breach, London, UK and New York, NY, USA, 1977.

Peierls:1977:MLR

[Pei77b] Rudolf Peierls. The momentum of light in a refracting medium. II. Generalization. Application to oblique reflexion. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 355(1681):141–151, June 1977. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1977RSPSA.355..141P.

Peierls:1978:PPV

[Pei78a] R. Peierls. A perspective of physics. Volume 2. Selections from 1977 Comments on Modern Physics. Gordon and Breach, London, UK,

1978. 34 + 259 pp. URL https://ui.adsabs.harvard.edu/#abs/
1978ppv..book.....P.

Peierls:1978:CAF

[Pei78b] Rudolf Peierls. Clustering in adsorbed films. Physical Review B: Condensed Matter and Materials Physics, 18(4):2013–2015, August 15, 1978. CODEN PRBMDO. ISSN 1098-0121. URL https://ui.

adsabs.harvard.edu/#abs/1978PhRvB..18.2013P.

Peierls:1978:I

[Pei78c] Rudolf Peierls. Introduction. In A Perspective of Physics, volume 2, pages ix–xxxiv. Gordon & Breach, London, UK and New York, NY, USA, 1978.

Peierls:1979:BRP

[Pei79a] Rudolf Peierls. Book review: A physicist who enjoys it. What Little I Remember, by Otto R. Frisch, pp. 219. (Cambridge University Press: Cambridge, 1979). Nature, 280(5719):257–259, July 19, 1979. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1979Natur.280..257P.

Peierls:1979:DOI

[Pei79b] Rudolf Peierls. The development of our ideas on the nuclear forces. In Stuewer [Stu79], pages 183–211. ISBN 0-8166-0869-5. LCCN QC773 .S95 1977.

Peierls:1979:STP

[Pei79c] Rudolf Peierls. Surprises in Theoretical Physics. Princeton Series in Physics. Princeton University Press, Princeton, NJ, USA, 1979. ISBN 0-691-08241-3 (hardcover), 0-691-08242-1 (paperback). viii +

167 pp. LCCN QC20 .P35. URL https://ui.adsabs.harvard.
edu/#abs/1979stp..book.....P.

Peierls:1980:RES

[Pei80a] R. E. Peierls. Recollections of early solid state physics. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 371(1744):28–38, June 1980. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1980RSPSA.371...28P.

Peierls:1980:M

[Pei80b] Rudolf Peierls. Misconceptions. In ????, editor, Proceedings of the 30th Pugwash Conference: Arms limitations, security & development, at Breukelen, The Netherlands, 20–25 August 1980, page 222. Taylor & Francis, London, UK, 1980.

URL https://catalog.hathitrust.org/Record/006302425;
https://pugwash.org/pugwash-international-conferences/.

Peierls:1980:NFD

[Pei80c] Rudolf Peierls. Nuclear family: Disturbing the Universe by Freeman Dyson Harper and Row, 283 pp, £6.95, November 1979, ISBN 0-06-011108-9 [from Winchester to Orion]. London Review of Books, 2(12):9–10, June 19, 1980. ISSN 0260-9592. URL https://www.

lrb.co.uk/v02/n12/rudolf-peierls/nuclear-family.

Peierls:1980:ORF

[Pei80d] Rudolf Peierls. Obituary: O. R. Frisch, 1904–1979. Nature, 284 (5752):196–197, March 13, 1980. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1980Natur.284..196P.

Peierls:1980:BRE

[Pei80e] Sir Rudolf Peierls. Book review: Energy from Heaven and Earth, by Edward Teller. Bulletin of the Atomic Scientists, 36(5):53–54, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1980:BRG

[Pei80f] Sir Rudolf Peierls. Book review: The growing pains of Robert Oppenheimer. Robert Oppenheimer: Letters and Recollections. Edited by Alice Kimball Smith and Charles Weiner. Pp. 250. (Harvard University Press: 1980). Nature, 286(5775):827–828, August 21, 1980. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1980Natur.286..827P.

Peierls:1980:MMP

[Pei80g] Sir Rudolf Peierls. Model-making in physics. Contemporary Physics, 21(1):3–17, January 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1980ConPh..21....3P.

Peierls:1980:PF

[Pei80h] Sir Rudolf Peierls. Particles and forces. In J. H. Mulvey, editor, The Nature of Matter, Wolfson College Lectures, 1980, pages 29–?? Clarendon Press, Oxford, UK, 1980. URL https://ui.adsabs.

harvard.edu/#abs/1980nama.book...29P.

Peierls:1981:AM

[Pei81a] Rudolf Peierls. Atom men. Financial Times, ??(??):8, September 5, 1981.

Peierls:1981:BRM

[Pei81b] Rudolf Peierls. Book review: Memories of the secret city. Reminiscences of Los Alamos, 1943–1945, Edited by Lawrence Badash, Joseph O. Hirschfelder, and Herbert P. Broida, pp. 188 (Reidel: 1980). Nature, 290(5806):529, April 9, 1981. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1981Natur.290..529P.

Peierls:1981:CWB

[Pei81c] Rudolf Peierls. Counting weapons: Britain and Nuclear Weapons, by Lawrence Freedman. Papermac, 160 pp, £3.25, September 1980, ISBN 0-333-30511-6. Countdown: Britain’s Strategic Forces, by Stewart Menual Hale, 188 pp, £8.25, October 1980, ISBN 0-7091-8592-8. The War Machine, by James Avery Joyce. Quartet, 210 pp, £6.95, October 1980, ISBN 0-7043-2254-4. Protest and Survive, edited by E.P. Thompson and Dan Smith. Penguin, 262 pp, £1.50, October 1980, ISBN 0-14-052341-3. London Review of Books, 3(4): 16–18, March 5, 1981. ISSN 0260-9592. URL https://www.lrb.

co.uk/v03/n04/rudolf-peierls/counting-weapons.

Peierls:1981:OC

[Pei81d] Rudolf Peierls. One culture. New York Review of Books, ??(??): 44–45, December 17, 1981. ISSN 0028-7504 (print), 1944-7744 (electronic).

Peierls:1981:ORF

[Pei81e] Sir Rudolf Peierls, F.R.S. Otto Robert Frisch, 1 October 1904–22 September 1979. Biographical Memoirs of Fellows of the Royal Society, 27:283–306, November 1981. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL http://www.jstor.org/stable/10.2307/769874;
 http://www.jstor.org/stable/769874;   https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1981.0012.

Peierls:1982:EDN

[Pei82a] R. Peierls. The early days of neutrino physics. In Ettore Fiorini, editor, Neutrino Physics and Astrophysics, Proceedings of the 10th International Conference held in Erice, Italy, 1982, pages 1– ?? Plenum Press, New York, NY, USA; London, UK, 1982. URL

https://ui.adsabs.harvard.edu/#abs/1982npa..conf....1P.

Peierls:1982:BRP

[Pei82b] Rudolf Peierls. Book review: Physicist extraordinary. Oliphant: The Life and Times of Sir Mark Oliphant. By Stewart Cockburn and David Ellyard. Pp. 369. ISBN 0-9594164-0-4. Nature, 296(5858):685, April 15, 1982. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1982Natur.296..685P.

Peierls:1982:KPA

[Pei82c] Rudolf Peierls. A key problem for arms control: is a first strike a realistic possibility? In ????, editor, Proceedings of the 32nd Pugwash Conference: The current danger of nuclear war: the relevance of the Russell–Einstein manifesto today, at Warsaw, Poland, August 1982, page 246. Taylor & Francis, London, UK, 1982. URL

https://pugwash.org/pugwash-international-conferences/.

Peierls:1982:OCT

[Pei82d] Rudolf Peierls. Odd couple [two mathematicians]. New York Review of Books, ??(??):16–18, February 18, 1982. ISSN 0028-7504 (print), 1944-7744 (electronic).

Peierls:1982:QSF

[Pei82e] Rudolf Peierls. Quantum squabbles [fact and fancy in physics]. New York Review of Books, ??(??):34–35, April 29, 1982. ISSN 0028-7504 (print), 1944-7744 (electronic).

Peierls:1982:LNW

[Pei82f] Sir Rudolf Peierls. Limited nuclear war? Bulletin of the Atomic Scientists, 38(5):2, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1983:BRP

[Pei83a] Rudolf Peierls. Book review: The physicists. Haphazard Reality: Half a Century of Science. By Hendrik B. G. Casimir. Harper and Row: 1983. Pp. 356. Nature, 303(5913):184–185, May 12, 1983. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1983Natur.303..184P.

Peierls:1983:FFP

[Pei83b] Rudolf Peierls. Far-flung physicist. Rend. Accad. Naz. Sci., 7(2): 235–246, ???? 1983.

Peierls:1983:MSP

[Pei83c] Rudolf Peierls. The momentum of a sound pulse in a slightly dispersive medium. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 390(1798):1–12, November 8, 1983. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-

9169 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1983RSPSA.390....1P.

Peierls:1983:WED

[Pei83d] Rudolf Peierls. What Einstein did: Subtle Is The Lord… the Science and the Life of Albert Einstein by Abraham Pais. New York Review of Books, 30(7):21–23, April 28, 1983. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.com/

articles/archives/1983/apr/28/what-einstein-did/.

Peierls:1983:EDN

[Pei83e] Sir Rudolf Peierls. The early days of neutrino physics. Contemporary Physics, 24(3):221–228, May 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1983ConPh..24..221P.

Peierls:1984:BRF

[Pei84a] Rudolf Peierls. Book reviews: Forty Years into the Atomic Age (pp. 285–291). The Nuclear Chain Reaction Forty Years Later by Robert G. Sachs. Minerva [Washington, DC], 22(2):281–291, Summer 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/accesspage/article/
10.1007/BF02207952;   http://www.jstor.org/stable/41820567.

Peierls:1984:FYA

[Pei84b] Rudolf Peierls. Forty years into the atomic age. Minerva [Washington, DC], 22(2):285–291, Summer 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02207952.

Peierls:1984:RCT

[Pei84c] Rudolf Peierls. Reminiscences of Cambridge in the Thirties. In Hendry [Hen84], pages 195–200. ISBN 0-85274-761-6. LCCN QC51.G72 C352 1984.

Peierls:1984:BND

[Pei84d] Sir Rudolf Peierls. Breaking the nuclear deadlock: First strike fiction. Bulletin of the Atomic Scientists, 40(1):50, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierlsn:1984:PAQ

[Pei84e] Rudolf Peierlsn. Prowling around quantum mechanics [struggling with quantum mechanics]. New Scientist, ??(??):45, December 13, 1984. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Peierls:1985:CRW

[Pei85a] Rudolf Peierls. Cambridge Russian who went home [Kapitza detained]. Financial Times, ??(??):xvi, May 11, 1985.

Peierls:1985:CW

[Pei85b] Rudolf Peierls. Closing words. In Symposium on the foundations of modern physics (Joensuu, 1985), page 713. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1985.

Peierls:1985:HTL

[Pei85c] Rudolf Peierls. Hiding their lights under a bushel. The Times [London, UK], ??(??):??, July 6, 1985. ISSN 0140-0460, 0956-1382.

Peierls:1985:MPL

[Pei85d] Rudolf Peierls. Momentum and pseudomomentum of light and sound. In F. Bassani, F. Fumi, and M. R. Tosi, editors, Highlights of condensed-matter theory, Proceedings of the International School of Physics “Enrico Fermi”, Course 89, pages 237–255. Elsevier, Amsterdam, The Netherlands, 1985.

Peierls:1985:OPA

[Pei85e] Rudolf Peierls. Obituary: Paul Adrien Maurice Dirac. Physics Today, 38(1):111, January 1, 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://physicstoday.

scitation.org/doi/abs/10.1063/1.2813724;   https://ui.
adsabs.harvard.edu/#abs/1985PhT....38a.111P.

Peierls:1985:OQM

[Pei85f] Rudolf Peierls. Observations in quantum mechanics and the “collapse of the wave function”. In Symposium on the foundations of modern physics (Joensuu, 1985), pages 187–196. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1985.

Peierls:1985:P

[Pei85g] Rudolf Peierls. A poem. In Symposium on the foundations of modern physics (Joensuu, 1985), page 1. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1985.

Peierls:1985:RBP

[Pei85h] Rudolf Peierls. Reflections of a British participant. Bulletin of the Atomic Scientists, 41(7):27–29, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1985:SRB

[Pei85i] Rudolf Peierls. Some recollections of Bohr. In French and Kennedy [FK85], pages 227–231. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985.

Peierls:1985:AP

[Pei85j] Sir Rudolf E. (Rudolf Ernst) Peierls. Assistant to Pauli. In Bird of Passage: Recollections of a Physicist [Pei85k], chapter 3, pages 46–81. ISBN 0-691-08390-8, 0-691-02416-2 (paperback), 0-691-60220-4, 1-4008-5461-X (e-book). LCCN QC16.P375 A32

1985. URL http://site.ebrary.com/id/10897402;   https:/
/www.jstor.org/stable/j.ctt7ztn4b.

Peierls:1985:BPR

[Pei85k] Sir Rudolf E. (Rudolf Ernst) Peierls. Bird of Passage: Recollections of a Physicist. Princeton legacy library. Princeton University Press, Princeton, NJ, USA, 1985. ISBN 0-691-08390-8, 0-691-02416-2 (paperback), 0-691-60220-4, 1-4008-5461-X (ebook). xii + 350 + 12 pp. LCCN QC16.P375 A32 1985.

URL http://site.ebrary.com/id/10897402;   https://www.
jstor.org/stable/j.ctt7ztn4b.

Peierls:1986:TC

[Pei86a] R. Peierls. Truth and clarity. In Jorrit de Boer, Erik Dal, and Ole Ulfbeck, editors, The Lesson of Quantum Theory, Niels Bohr Centenary Symposium held 3–7 October, 1985 in Copenhagen, Denmark, pages 379–380. North-Holland, Amsterdam, The Nether-

lands, January 1986. URL https://ui.adsabs.harvard.edu/#
abs/1986lqt..conf..379P.

Peierls:1986:CR

[Pei86b] Rudolf Peierls. Conservative revolutionary. New York Review of Books, ??(??):56–57, November 20, 1986. ISSN 0028-7504 (print), 1944-7744 (electronic).

Peierls:1986:FFW

[Pei86c] Rudolf Peierls. Fact or fission [wild speculation]. London Review of Books, 8(18):4, October 23, 1986. ISSN 0260-9592. URL https:/

/www.lrb.co.uk/v08/n18/letters#letter4.

Peierls:1986:F

[Pei86d] Rudolf Peierls. Foreword. In Niels Bohr: collected works, volume 9, pages 1–83. Elsevier, Amsterdam, The Netherlands, 1986.

Peierls:1986:I

[Pei86e] Rudolf Peierls. Introduction. In Niels Bohr: collected works, volume 9, pages 1–83. Elsevier, Amsterdam, The Netherlands, 1986.

Peierls:1986:TLN

[Pei86f] Rudolf Peierls. Is there any logic in the nuclear arms race? In A. Boserup, L. Christensen, and O. Nathan, editors, The Challenge of Nuclear Armaments, pages 265–269. Rhodos International for the University of Copenhagen, Copenhagen, Denmark, 1986.

Peierls:1986:PHB

[Pei86g] Rudolf Peierls. Physics and Homi Bhabha: Homi Jehangir Bhabha: Collected Scientific Papers. Edited by B. V. Sreekantan, Virendra Singh, and B. M. Udgaonkar. Tata Institute, Homi Bhabha Road, Colaba, Bombay 400 005, India: 1986. Pp. 1023. Nature, 323 (6085):212, September 18, 1986. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1986Natur.323..212P.

Peierls:1986:RBP

[Pei86h] Rudolf Peierls. Reflections of a British participant. In Assessing the Nuclear Age: Selections from the Bulletin of the Atomic Scientists [AM86], page ?? ISBN 0-941682-07-2, 0-941682-08-0 (paperback). LCCN U264 .A79 1986. Reprint of [Pei85h].

Peierls:1986:SDI

[Pei86i] Rudolf Peierls. Strategic Defence Initiative. Physics Bulletin, 37 (??):217, ???? 1986. CODEN PHSBB4. ISSN 0031-9112.

Peierls:1987:ADM

[Pei87a] Rudolf Peierls. Address to Dirac Memorial Meeting, Cambridge. In Taylor [Tay87], pages 35–37. ISBN 0-85274-480-3.

LCCN QC16.D57 T75 1987. URL http://www.loc.gov/catdir/
enhancements/fy0745/87153334-d.html. Based on the papers

presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Peierls:1987:BRC

[Pei87b] Rudolf Peierls. Book review: The case for the defence: Better a Shield Than a Sword: Perspectives on Defense and Technology, by Edward Teller. Free Press, New York/Collier Macmillan, London: 1987, pp. 257. US$19.95, £17.25. Nature, 328(6131):583, August 13, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1987Natur.328..583P.

Peierls:1987:DW

[Pei87c] Rudolf Peierls. Dirac’s way. In Kurşunoğlu and Wigner [KW87], pages 43–45. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987.

URL http://adsabs.harvard.edu/abs/1987ragp.book.....D;

 http://www.loc.gov/catdir/description/cam023/86033409.
html;   http://www.loc.gov/catdir/toc/cam028/86033409.
html.

Peierls:1987:HDD

[Pei87d] Rudolf Peierls. How deep is deep enough? In ????, editor, Proceedings of the 37th Pugwash Conference: From confrontation to rapprochement: 30th anniversary of Pugwash, at Gmunden, Austria, in Sept. 1987, pages 414–416. Taylor & Francis, London,

UK, 1987. URL https://pugwash.org/pugwash-international-
conferences/.

Peierls:1987:MIM

[Pei87e] Rudolf Peierls. Making it [the making of the atomic bomb]. New York Review of Books, ??(??):47–48, November 5, 1987. ISSN 0028-7504 (print), 1944-7744 (electronic).

Peierls:1987:UR

[Pei87f] Rudolf Peierls. Mit Umlaut: a reply. New York Review of Books, ??(??):41, February 12, 1987. ISSN 0028-7504 (print), 1944-7744 (electronic).

Peierls:1987:MHA

[Pei87g] Rudolf Peierls. Models, hypotheses and approximations. In Metropolis et al. [MKR87], pages 95–105. ISBN 0-12-492155-8. LCCN QC44 .N49 1987. URL http://www.osti.gov/

energycitations/product.biblio.jsp?osti_id=6120718&query_
id=0.

Peierls:1988:APT

[Pei88a] Rudolf Peierls. An appreciation of Professor Tony Skyrme. In A breadth of physics (Oxford, 1987), pages 203–204. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1988.

Peierls:1988:CMV

[Pei88b] Rudolf Peierls. The case of Mordechai Vanunu. New York Review of Books, ??(??):56, June 16, 1988. ISSN 0028-7504 (print), 1944-7744 (electronic).

Peierls:1988:FEE

[Pei88c] Rudolf Peierls. First encounter with English food. In Kurti and Kurti [KK88], pages 203–205. ISBN 0-85274-301-7. LCCN TX651 .B89 1988. With a foreword by Sir George Porter.

Peierls:1988:GSR

[Pei88d] Rudolf Peierls. Getting out of a spin: a reply. New York Review of Books, ??(??):45, March 3, 1988. ISSN 0028-7504 (print), 1944-7744 (electronic).

Peierls:1988:MLL

[Pei88e] Rudolf Peierls. Momentum of light and light pressure in a dispersive medium. In ????, editor, Interaction of radiation with matter, Gozzoni symposium at Scuola Normale, Pisa, 1988, pages 187–191. ????, ????, 1988.

Peierls:1988:RB

[Pei88f] Sir Rudolf Peierls, F.R.S. Rutherford and Bohr. Notes and Records of the Royal Society of London, 42(2):229–241, July 1, 1988. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

URL http://www.jstor.org/stable/531359.

Peierls:1989:BS

[Pei89a] Rudolf Peierls. Broken symmetries. In ????, editor, Symmetries in nature, Radicati symposium, Scuola Normale, Pisa, 1989, pages 673–676. ????, ????, 1989.

Peierls:1989:MCF

[Pei89b] Rudolf Peierls. Microwave cooking for foxes. Interdisciplinary Science Review, 14(2):187, ???? 1989.

Peierls:1989:MRL

[Pei89c] Rudolf Peierls. My recollections of Landau. In Khalatnikov [Kha89], pages 194–197. ISBN 0-08-036383-0. LCCN QC16.L25 L38 1989.

Peierls:1989:NWH

[Pei89d] Rudolf Peierls. Nuclear weapons. How did we get there and where are we going? Oxford Magazine, ??(51):10–11, 1989. 4th week, Michaelmas Term.

Peierls:1989:RDF

[Pei89e] Rudolf Peierls. Reflections on the discovery of fission. Nature, 342(6252):852–854, December 28, 1989. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1989Natur.342..852P.

Peierls:1989:TBV

[Pei89f] Rudolf Peierls. Test ban and verification. In Joseph Rotblat and John P. Holdred, editors, Proceedings of the 39th Pugwash Conference: Building global security through cooperation, at Cambridge, USA, 23–28 July 1989, pages 311–312. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-642-75845-2 (print), 3-642-75843-6 (e-book). LCCN QC176-176.9.

URL https://catalog.hathitrust.org/Record/010161896;
https://pugwash.org/pugwash-international-conferences/.

Peierls:1990:EC

[Pei90a] Rudolf Peierls. Energy conservation. In ????, editor, Proceedings of the 40th Pugwash Conference: Towards a secure world in the 21st century, at Egham, UK, in September 1990, pages 530–531. Taylor & Francis, London, UK, 1990. URL https://pugwash.

org/pugwash-international-conferences/.

Peierls:1990:EFQ

[Pei90b] Rudolf Peierls. Erinnerungen aus der Frühzeit der Quantenmechanik. (German) [Reminiscences of the early days of quantum mechanics]. Jahrb. Dtsch. Akad. Naturforscher Leopoldina, ??(??): 137–140, 1990.

Peierls:1990:L

[Pei90c] Rudolf Peierls. Landau in the 1930s. In E. A. Gotsman, Y. Ne’eman, and A. Voronel, editors, Frontiers of Physics, Proceedings of the Landau Memorial Conference held 6–10 June, 1988 in Tel Aviv, Israel, pages 23–26. Pergamon Press, New York, NY, USA, January

1990. URL https://ui.adsabs.harvard.edu/#abs/1990ftph.
conf...23P.

Peierls:1991:POD

[Pei91a] R. Peierls. Physics of our days: Momentum and pseudomomentum of light and sound. Soviet Physics. Uspekhi, 34(9):817, September 1991. CODEN SOPUAP. ISSN 0038-5670 (print), 2169-5296

(electronic). URL https://ufn.ru/en/articles/1991/9/d/;
https://ui.adsabs.harvard.edu/#abs/1991SvPhU..34..817P.

Peierls:1991:BEW

[Pei91b] Rudolf Peierls. Bell’s early work. Europhysics News, 22(4):69– 70, January 1991. CODEN EUPNAS. ISSN 0531-7479 (print),

1432-1092 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1991ENews..22...69P;   https://www.europhysicsnews.
org/articles/epn/pdf/1991/04/epn19912204p69.pdf.

Peierls:1991:DPE

[Pei91c] Rudolf Peierls. Does physics ever come to an end? In M. Suzuki and R. Kubo, editors, Evolutionary trends in the physical sciences, pages 35–38. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991.

Peierls:1991:DM

[Pei91d] Rudolf Peierls. In defence of “measurement”. Physics World, 4(1): 19–20, ???? 1991. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

Peierls:1991:MST

[Pei91e] Rudolf Peierls. More Surprises in Theoretical Physics. Princeton Series in Physics. Princeton University Press, Princeton, NJ, USA, 1991. ISBN 0-691-08576-5 (hardcover), 0-691-02522-3 (paperback). viii + 106 pp. LCCN QC20 .P345 1991. URL http://www.loc.

gov/catdir/description/prin021/90023189.html;   https:/
/ui.adsabs.harvard.edu/#abs/1991mstp.book.....P.

Peierls:1991:NT

[Pei91f] Rudolf Peierls. The nuclear threat. New York Review of Books, ??(??):61, November 7, 1991. ISSN 0028-7504 (print), 1944-7744 (electronic).

Peierls:1991:NTP

[Pei91g] Rudolf Peierls. The nuclear threat: A proposal. New York Review of Books, 38(18):??, November 7, 1991. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.com/

articles/archives/1991/nov/07/the-nuclear-threat-2/. See

reply [BGM91].

Peierls:1991:OWG

[Pei91h] Rudolf Peierls. Obituary: William George Penney. Physics Today, 44(10):138, 140, 142, October 1, 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/abs/10.1063/1.2810303;
https://ui.adsabs.harvard.edu/#abs/1991PhT....44j.138P.

Peierls:1991:RBS

[Pei91i] Rudolf Peierls. Reflections on broken symmetry. Physics Today, 44(2):13–14, February 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Peierls:1991:SBS

[Pei91j] Rudolf Peierls. Spontaneously broken symmetries. Journal of Physics A (Mathematical and General), 24(22):5273–5281, November 1991. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-

6447 (electronic). URL http://stacks.iop.org/0305-4470/
24/5273;   https://ui.adsabs.harvard.edu/#abs/1991JPhA.
..24.5273P.

Peierls:1991:WNW

[Pei91k] Rudolf Peierls. Why are nuclear weapons tests claimed to be necessary? In ????, editor, Proceedings of the 41st Pugwash Conference: Striving for peace, security, & development in the world, at Beijing, China, in September 1991, page 372. Taylor & Francis, London,

UK, 1991. URL https://pugwash.org/pugwash-international-
conferences/.

Peierls:1992:BSD

[Pei92a] R. Peierls. Broken symmetries: (Dirac memorial lecture, Cambridge, 15 June 1992). Contemporary Physics, 33(4):221–226, July 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-

5812 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1992ConPh..33..221P.

Peierls:1992:BRM

[Pei92b] Rudolf Peierls. Book reviews: The making of the atom bomb. British Scientists and the Manhattan Project: The Los Alamos years. By Ference Morton Szasz. Macmillan/St. Martin’s Press: 1992. Pp. 167. The Los Alamos Primer: The First Lectures on How to Build an Atom Bomb. By Robert Sperber. University of California Press: 1992. Pp. 98. Nature, 357(6376):290–291, May 28, 1992. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1992Natur.357..290P.

Peierls:1992:GEB

[Pei92c] Rudolf Peierls. German ethics and the bomb. The Independent, ?? (??):19, June 13, 1992.

Peierls:1992:LHB

[Pei92d] Rudolf Peierls. Letters: Heisenberg’s bomb. New Scientist, 135(1842):??, October 10, 1992. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL http:

//www.newscientist.com/article/mg13618426.100-letters-
heisenbergs-bomb.html. See [Cha92].

Peierls:1992:PED

[Pei92e] Rudolf Peierls. People in the early days of quantum mechanics. Fizika (Zagreb), A1(??):11–30, ???? 1992.

Peierls:1992:SCE

[Pei92f] Rudolf Peierls. Stone’s caricature of an education. The Daily Telegraph, ??(??):16, March 3, 1992.

Peierls:1992:USH

[Pei92g] Rudolf Peierls. The uncertain scientist [a Heisenberg biography]. New York Review of Books, ??(??):43–45, April 23, 1992. ISSN 0028-7504 (print), 1944-7744 (electronic).

Peierls:1992:WPM

[Pei92h] Rudolf Peierls. Where Pauli made his ‘wrong’ remark. Physics Today, 45(12):112, December 1992. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v45/i12/p112/s2;   https://ui.adsabs.harvard.
edu/#abs/1992PhT....45l.112P.

Peierls:1993:SHG

[Pei93a] Rudolf Peierls. Als Student bei Heisenberg. (German) [As a student of Heisenberg]. In Kleint and Wiemers [KW93], pages 104–107. ISBN 3-05-501585-1. LCCN QC16.H35 W48 1933. EUR 607 (!).

Peierls:1993:BNW

[Pei93b] Rudolf Peierls. The bomb that never was: [review of Heisenberg’s War: The Secret History of the German Bomb Project, by Thomas Powers (Knopf, New York, 1993)]. New York Review of Books, ?? (??):6–9, April 22, 1993. ISSN 0028-7504 (print), 1944-7744 (electronic).

Peierls:1993:EWS

[Pei93c] Rudolf Peierls. Early work on solids, mainly in the thirties. Reviews of Modern Physics, 65(2):251–254, April 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.65.
251;   http://rmp.aps.org/abstract/RMP/v65/i2/p251_1;
https://ui.adsabs.harvard.edu/#abs/1993RvMP...65..251P.

Peierls:1993:FEE

[Pei93d] Rudolf Peierls. First encounter with English food. In Kurti and Kurti [KK88], pages 203–205. ISBN 0-85274-301-7. LCCN TX651 .B89 1988. With a foreword by Sir George Porter.

Peierls:1993:P

[Pei93e] Rudolf Peierls. Physics 1920–45. Atti Convegni Lincei, 104(??): 23–33, ???? 1993. Proceedings of the Symposium “Enrico Fermi” (50th anniversary of the first reactor) held 10 December 1992 in Rome, Italy.

Peierls:1994:BRR

[Pei94a] Rudolf Peierls. Book review: Roger G. Newton, What Makes Nature Tick? Physics Today, 47(2):100, February 1994. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic).  URL https://physicstoday.scitation.org/doi/
abs/10.1063/1.2808407;   https://ui.adsabs.harvard.edu/
#abs/1994PhT....47b.100N.

Peierls:1994:LEB

[Pei94b] Rudolf Peierls. Letter to the Editor: Is Boltzmann entropy time’s arrow’s archer? Physics Today, 47(11):115, 117, November 1, 1994. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL https://physicstoday.scitation.org/doi/
abs/10.1063/1.2808690;   https://ui.adsabs.harvard.edu/
#abs/1994PhT....47k..11B.

Peierls:1994:YIF

[Pei94c] Rudolf Peierls. Yakov Il’ich Frenkel. Physics Today, 47(6):44–49, June 1, 1994. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL https://physicstoday.scitation.org/
doi/abs/10.1063/1.881435;   https://ui.adsabs.harvard.
edu/#abs/1994PhT....47f..44P.

Peierls:1994:RDR

[Pei94d] Sir Rudolf Peierls, F.R.S. Reminiscences and discoveries: Recollections of James Chadwick. Notes and Records of the Royal Society of London, 48(1):135–141, January 1994. CODEN NOREAY.

ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531426.

Peierls:1995:BRA

[Pei95a] Rudolf Peierls. Book review: Alan Cook, The Observational Foundations of Physics. Physics Today, 48(6):56, 58, June 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL https://physicstoday.scitation.org/doi/
abs/10.1063/1.2808066;   https://ui.adsabs.harvard.edu/
#abs/1995PhT....48f..56C.

Peierls:1995:CB

[Pei95b] Rudolf Peierls. The Communist bomb. New York Review of Books, ??(??):12–14, February 16, 1995. ISSN 0028-7504 (print), 1944-7744 (electronic).

Peierls:1996:QTS

[Pei96] R. E. Peierls. Quantum Theory of Solids. International Series of Monographs on Physics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1996. ISBN 0-19-267017-4, 0-19-851240-6. viii + 229 pp.

Peierls:1997:BTN

[Pei97a] R. Peierls. The Bohr theory of nuclear reactions. In Dalitz and Peierls [DP97a], pages 283–302. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN

QC21.2 .P42 1997. URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book..283P.

Peierls:1997:BS

[Pei97b] R. Peierls. Broken symmetries. In Dalitz and Peierls [DP97a], pages 780–785. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL

https://ui.adsabs.harvard.edu/#abs/1997sspr.book..780P.

Peierls:1997:CCN

[Pei97c] R. Peierls. Critical conditions in neutron multiplication. In Dalitz and Peierls [DP97a], pages 264–269. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN

QC21.2 .P42 1997. URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book..264P.

Peierls:1997:ETS

[Pei97d] R. Peierls. Expansions in terms of sets of functions with complex eigenvalues. In Dalitz and Peierls [DP97a], pages 357–366. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.

adsabs.harvard.edu/#abs/1997sspr.book..357P.

Peierls:1997:EUP

[Pei97e] R. Peierls. Extension of the Uncertainty Principle to relativistic quantum theory. In Dalitz and Peierls [DP97a], pages 83–94. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.

adsabs.harvard.edu/#abs/1997sspr.book...83P.

Peierls:1997:HCL

[Pei97f] R. Peierls. Heat conduction in liquid helium. In Dalitz and Peierls [DP97a], pages 235–238. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42

1997. URL https://ui.adsabs.harvard.edu/#abs/1997sspr.
book..235P.

Peierls:1997:MTC

[Pei97g] R. Peierls. Magnetic transition curves of supraconductors. In Dalitz and Peierls [DP97a], pages 203–218. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN

QC21.2 .P42 1997. URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book..203P.

Peierls:1997:MPL

[Pei97h] R. Peierls. Momentum and pseudomomentum of light and sound. In Dalitz and Peierls [DP97a], pages 745–764. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book).

LCCN QC21.2 .P42 1997. URL https://ui.adsabs.harvard.
edu/#abs/1997sspr.book..745P.

Peierls:1997:NDE

[Pei97i] R. Peierls. Note on the derivation of the equation of state for a degenerate relativistic gas. In Dalitz and Peierls [DP97a], pages 198–202. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL

https://ui.adsabs.harvard.edu/#abs/1997sspr.book..198P.

Peierls:1997:MPF

[Pei97j] R. Peierls. On a minimum property of the free energy. In Dalitz and Peierls [DP97a], pages 258–259. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN

QC21.2 .P42 1997. URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book..258P.

Peierls:1997:IMF

[Pei97k] R. Peierls. On Ising’s model of ferromagnetism. In Dalitz and Peierls [DP97a], pages 225–229. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN

QC21.2 .P42 1997. URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book..225P.

Peierls:1997:ASS

[Pei97l] R. Peierls. On the absorption spectra of solids. In Dalitz and Peierls [DP97a], pages 95–96. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42

1997. URL https://ui.adsabs.harvard.edu/#abs/1997sspr.
book...95P.

Peierls:1997:ESS

[Pei97m] R. Peierls. On the existence of stationary states. In Dalitz and Peierls [DP97a], pages 12–14. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN

QC21.2 .P42 1997. URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book...12P.

Peierls:1997:KTT

[Pei97n] R. Peierls. On the kinetic theory of thermal conduction in crystals. In Dalitz and Peierls [DP97a], pages 15–48. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book).

LCCN QC21.2 .P42 1997. URL https://ui.adsabs.harvard.
edu/#abs/1997sspr.book...15P.

Peierls:1997:SBE

[Pei97o] R. Peierls. On the statistical basis for the electron theory of metals. In Dalitz and Peierls [DP97a], pages 132–136. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.adsabs.

harvard.edu/#abs/1997sspr.book..132P.

Peierls:1997:TET

[Pei97p] R. Peierls. On the theory of electric and thermal conductivity of metals. In Dalitz and Peierls [DP97a], pages 49–68. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.adsabs.

harvard.edu/#abs/1997sspr.book...49P.

Peierls:1997:TGM

[Pei97q] R. Peierls. On the theory of galvano-magnetic effects. In Dalitz and Peierls [DP97a], pages 1–10. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN

QC21.2 .P42 1997. URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book....1P.

Peierls:1997:TDCb

[Pei97r] R. Peierls. On the theory of the diamagnetism of conduction electrons. In Dalitz and Peierls [DP97a], pages 97–120. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.adsabs.

harvard.edu/#abs/1997sspr.book...97P.

Peierls:1997:TDCa

[Pei97s] R. Peierls. On the theory of the diamagnetism of conduction electrons II: Strong magnetic fields. In Dalitz and Peierls [DP97a], pages 121–126. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL

https://ui.adsabs.harvard.edu/#abs/1997sspr.book..121P.

Peierls:1997:THE

[Pei97t] R. Peierls. On the theory of the Hall Effect. In Dalitz and Peierls [DP97a], page 11. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42

1997. URL https://ui.adsabs.harvard.edu/#abs/1997sspr.
book...11P.

Peierls:1997:PT

[Pei97u] R. Peierls. The Peierls transition. In Dalitz and Peierls [DP97a], pages 439–440. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL

https://ui.adsabs.harvard.edu/#abs/1997sspr.book..439P.

Peierls:1997:QEC

[Pei97v] R. Peierls. Quantum electrodynamics in configuration space. In Dalitz and Peierls [DP97a], pages 71–82. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN

QC21.2 .P42 1997. URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book...71P.

Peierls:1997:RTM

[Pei97w] R. Peierls. Remarks on the theory of metals. In Dalitz and Peierls [DP97a], pages 127–131. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42

1997. URL https://ui.adsabs.harvard.edu/#abs/1997sspr.
book..127P.

Peierls:1997:RTT

[Pei97x] R. Peierls. Remarks on transition temperatures. In Dalitz and Peierls [DP97a], pages 137–138. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN

QC21.2 .P42 1997. URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book..137P.

Peierls:1997:SD

[Pei97y] R. Peierls. The size of a dislocation. In Dalitz and Peierls [DP97a], pages 273–276. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL

https://ui.adsabs.harvard.edu/#abs/1997sspr.book..273P.

Peierls:1997:SEC

[Pei97z] R. Peierls. Statistical error in counting experiments. In Dalitz and Peierls [DP97a], pages 162–181. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN

QC21.2 .P42 1997. URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book..162P.

Peierls:1997:STA

[Pei97-27] R. Peierls. Statistical theory of adsorption with interaction between the adsorbed atoms. In Dalitz and Peierls [DP97a], pages 219–224. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.

adsabs.harvard.edu/#abs/1997sspr.book..219P.

Peierls:1997:STS

[Pei97-28] R. Peierls. Statistical theory of superlattices with unequal concentrations of the components. In Dalitz and Peierls [DP97a], pages 182–197. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL

https://ui.adsabs.harvard.edu/#abs/1997sspr.book..182P.

Peierls:1997:TRT

[Pei97-29] R. Peierls. Two remarks on the theory of conductivity. In Dalitz and Peierls [DP97a], pages 69–70. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN

QC21.2 .P42 1997. URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book...69P.

Peierls:1997:VDN

[Pei97-30] R. Peierls. Velocity-dependent nuclear forces. In Dalitz and Peierls [DP97a], pages 528–529. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42

1997. URL https://ui.adsabs.harvard.edu/#abs/1997sspr.
book..528P.

Peierls:1997:CLR

[Pei97-31] R. E. Peierls. The commutation laws of relativistic field theory. In Dalitz and Peierls [DP97a], pages 367–382. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book).

LCCN QC21.2 .P42 1997. URL https://ui.adsabs.harvard.
edu/#abs/1997sspr.book..367P.

Peierls:1997:CES

[Pei97-32] R. E. Peierls. Complex eigenvalues in scattering theory. In Dalitz and Peierls [DP97a], pages 454–474. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN

QC21.2 .P42 1997. URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book..454P.

Peierls:1997:IPP

[Pei97-33] R. E. Peierls. Interpretation and properties of propagators. In Dalitz and Peierls [DP97a], pages 435–438. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book).

LCCN QC21.2 .P42 1997. URL https://ui.adsabs.harvard.
edu/#abs/1997sspr.book..435P.

Peierls:1997:NVS

[Pei97-34] R. E. Peierls. Note on the vibration spectrum of a crystal. In Dalitz and Peierls [DP97a], pages 421–426. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN

QC21.2 .P42 1997. URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book..421P.

Peierls:1997:STN

[Pei97-35] R. E. Peierls. Selected topics in nuclear theory. In Dalitz and Peierls [DP97a], pages 475–527. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42

1997. URL https://ui.adsabs.harvard.edu/#abs/1997sspr.
book..475P.

Peierls:1997:TRS

[Pei97-36] R. E. Peierls. Time reversal and the Second Law of Thermodynamics. In Dalitz and Peierls [DP97a], pages 563–570. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.adsabs.

harvard.edu/#abs/1997sspr.book..563P.

Peierls:1997:DM

[Pei97-37] Rudolf Peierls. In defence of “measurement”. In Dalitz and Peierls [DP97a], pages 778–779. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42

1997. URL https://ui.adsabs.harvard.edu/#abs/1997sspr.
book..778P.

Peierls:1997:MMP

[Pei97-38] Rudolf Peierls. Model-making in physics. In Dalitz and Peierls [DP97a], pages 718–732. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42

1997. URL https://ui.adsabs.harvard.edu/#abs/1997sspr.
book..718P.

Peierls:1997:MSP

[Pei97-39] Rudolf Peierls. The momentum of a sound pulse in a slightly dispersive medium. In Dalitz and Peierls [DP97a], pages 733–744. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.

adsabs.harvard.edu/#abs/1997sspr.book..733P.

Peierls:1997:MLR

[Pei97-40] Rudolf Peierls. The momentum of light in a refracting medium. In Dalitz and Peierls [DP97a], pages 660–676. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book).

LCCN QC21.2 .P42 1997. URL https://ui.adsabs.harvard.
edu/#abs/1997sspr.book..660P.

Peierls:1997:OQM

[Pei97-41] Rudolf Peierls. Observations in quantum mechanics and the “collapse of the wave function”. In Dalitz and Peierls [DP97a], pages 765–774. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL

https://ui.adsabs.harvard.edu/#abs/1997sspr.book..765P.

Peierls:1997:PTP

[Pei97-42] Rudolf Peierls. Perturbation theory for projected states. In Dalitz and Peierls [DP97a], pages 579–592. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN

QC21.2 .P42 1997. URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book..579P.

Peierls:1997:RB

[Pei97-43] Rudolf Peierls. Rutherford and Bohr. Current Science, 73(8): 707–712, October 25, 1997. CODEN CUSCAM. ISSN 0011-

3891. URL http://www.currentscience.ac.in/Downloads/
download_pdf.php?titleid=id_073_08_0707_0712_0. Text of

Rutherford Memorial Lecture, given at National Physical Laboratory, New Delhi, India, November 10, 1987.

Peierls:1997:SSR

[Pei97-44] Sir Rudolf Peierls. Some simple remarks on the basis of transport theory. In Dalitz and Peierls [DP97a], pages 622–654. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.adsabs.

harvard.edu/#abs/1997sspr.book..622P.

Peierls:1997:AH

[Pei97-45] Sir Rudolf Ernst Peierls. Atomic Histories, volume 18 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1997. ISBN 1-56396-243-8 (hardcover). xvii + 378 pp. LCCN QC71 .P38 1997.

Peierls:2001:QTS

[Pei01] R. E. Peierls. Quantum Theory of Solids. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-850781-X, 0-19-267017-4, 0-19-151648-1 (e-book). viii + 229 pp. LCCN QC176 .P42 2001.

Peierls:2007:LFC

[Pei07] Rudolf Peierls. Lesson of the Fuchs case. In Lee [Lee09], chapter 6, page ?? ISBN 981-256-503-5 (vol. 1), 981-279-706-8 (vol. 2). LCCN QC16.P375 A4 2007. URL http://wspc-prod.

literatumonline.com/worldscibooks/10.1142/6790;   http:
//www.worldscientific.com/worldscibooks/10.1142/5941.

Peierls:2008:NBC

[Pei08] Sir Rudolf Peierls, editor. Niels Bohr collected works. Volume 9, Nuclear physics (1929–1952). Elsevier, Amsterdam, The Netherlands, 2008. ISBN 0-444-53277-3. xviii + 693 pp. LCCN ????

Peierls:2009:DQT

[Pei09] R. E. Peierls. The development of quantum theory. Contemporary Physics, 50(1):25–35, January 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/2009ConPh..50...25P.

Peierls:2010:RB

[Pei10] Rudolf Peierls. Rutherford and Bohr. Resonance, 15(5):476–487, May 2010. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic). Text of the Rutherford Memorial Lecture delivered by Sir Rudolf Peierls, FRS, on 10 November 1987 at the National Physical Laboratory, New Delhi. Reproduced from Current Science, Vol.73, pp.707–712, 25 October, 1997.

Perrin:1939:CRC

[Per39] Francis Perrin. Calcul relatif aux conditions éventuelles de transmutation en chaîne de l’uranium. (French) [Calculation of possible conditions on the uranium chain transmutation]. Comptes rendus de l’Académie des sciences, Paris, 208(18):1394–

1396, May 1, 1939. URL http://gallica.bnf.fr/ark:/12148/
bpt6k6238836c/f26.image.r=perrin%20calcul%20relative%20aux%
20conditions#.

Peierls:1940:EFNb

[PF40] Rudolf Peierls and Otto R. Frisch. The effects of fast neutrons in ordinary uranium. Report MS-4A (PRO AB4/832), Directorate of Tube Alloys, ????, UK, 1940.

Peierls:1941:ESI

[PF41] Rudolf Peierls and Klaus Fuchs. Effect on separation of isotopes of compound molecules. Report MS-44 (PRO AB4/878), Directorate of Tube Alloys, ????, UK, 1941.

Peierls:1942:GP

[PF42] Rudolf Peierls and Klaus Fuchs. Gland problems. Report MS-55 (PRO AB4/891), Directorate of Tube Alloys, ????, UK, 1942.

Peierls:1943:CCP

[PF43a] Rudolf Peierls and Klaus Fuchs. Cascades of cascades — Part II: sub-cascades composed of two compressor sizes. Report MS-90 (PRO AB4/939), Directorate of Tube Alloys, ????, UK, 1943.

Peierls:1943:CCD

[PF43b] Rudolf Peierls and Klaus Fuchs. Cascades of cascades in the D. S. system. Report MS-83 (PRO AB4/921), Directorate of Tube Alloys, ????, UK, 1943.

Peierls:1943:ENU

[PF43c] Rudolf Peierls and Klaus Fuchs. Effect of non-uniform hole size on separation. Report MS-101 (PRO AB4/940), Directorate of Tube Alloys, ????, UK, 1943.

Peierls:1943:RM

[PF43d] Rudolf Peierls and Klaus Fuchs. The rabbit machine. Report MS-81 (PRO AB4/919), Directorate of Tube Alloys, ????, UK, 1943.

Peierls:1943:ARN

[PFD43] Rudolf Peierls, Klaus Fuchs, and P. A. M. Dirac. Approximate rate of neutron multiplication for a solid of arbitrary shape and uniform density: Part II, Application to the oblate spheroid, hemisphere and oblate hemispheroid. Report MS-D#.5 (PRO AB4/963), Directorate of Tube Alloys, ????, UK, 1943.

Peierls:1991:LER

[PKA91] Rudolf Peierls, Thomas A. Kaplan, and Philip W. Anderson. Letter to the Editor: Reflections on broken symmetry. Physics Today, 44(2):13, February 1, 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/abs/10.1063/1.2809982;
https://ui.adsabs.harvard.edu/#abs/1991PhT....44b..13P.

Peierls:1980:LAN

[PM80] Rudolf Peierls and Nicole Vinh Mau. Local approximation to a nonlocal potential. Nuclear Physics A, 343:1–23, July 1980. CODEN NUPABL. ISSN 0375-9474 (print), 1873-1554 (electronic). URL

https://ui.adsabs.harvard.edu/#abs/1980NuPhA.343....1P.

Peierls:1997:LAN

[PM97] Rudolf Peierls and Nicole Vinh Mau. Local approximation to a non-local potential. In Dalitz and Peierls [DP97a], pages 695–717. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.

adsabs.harvard.edu/#abs/1997sspr.book..695P.

Peierls:1958:MEN

[PMB^{+}58] R. E. Peierls, P. T. Matthews, G. H. Burkhardt, B. W. Ridley, A. Salam, D. L. Pursey, P. B. Moon, V. Knapp, J. M. Cassels, P. Farago, C. F. Coleman, G. M. Lewis, and A. Lundby. The measurement of electron–neutrino angular correlations: Discussion. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 246:475–481, August 1958. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL

https://ui.adsabs.harvard.edu/#abs/1958RSPSA.246..475P.

Peierls:1977:PPV

[POZ^{+}77] R. Peierls, L. B. Okun, Ya. B. Zel’dovich, B. W. Lee, C. Quigg, H. Harari, M. Jacob, D. Cline, A. K. Mann, Q. Lewin Keller, J. W. Negele, F. S. Stephens, A. J. Leggett, R. W. Keyes, J. Tauc, S. R. Nagel, S. W. Lovesey, A. H. Thompson, B. Paczynski, P. Goldreich, W. Sargent, R. B. Larson, J. E. Grindlay, L. Spitzer, S. J. Brodsky, G. Karl, J. L. Picqué, H. H. Stroke, V. Dose, R. V. Ambartzumian, V. S. Letokhov, F. M. Harris, G. Bekefi, C. Deutsch, H. P. Furth, V. N. Tsytovich, and J. D. Callen. A perspective of physics. Volume 1. Selections from 1976 Comments on Modern Physics. Gordon and Breach, London, UK, 1977. 35 + 243 pp. URL https://ui.

adsabs.harvard.edu/#abs/1977ppv..book.....P.

Peierls:1947:PPS

[PP47] R. E. Peierls and M. A. Preston. Proton–proton scattering at 10 MeV. Physical Review, 72(3):250, August 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1947PhRv...72..250P.

Peierls:1968:PW

[PP68] Rudolf Peierls and C. F. Powell. Pugwash & the world in 1968. In ????, editor, Proceedings of the 18th Pugwash Conference: Current problems of peace, security & development, at Nice, France, Sept. 1968, pages 209–212. Taylor & Francis, London, UK, 1968. URL

https://pugwash.org/pugwash-international-conferences/.

Peierls:1983:DMP

[PPG83] Rudolf Peierls, W. H. Press, and T. Gold. Dependence of macrophysical phenomena on the values of the fundamental constants: Discussion. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 310:334–336, December 1983. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1983RSPTA.310..334P.

Penney:1944:SHB

[PPvNF44] W. G. Penney, Rudolf E. Peierls, John von Neumann, and Klaus Fuchs. Shock hydrodynamics and blast waves. Report AECD 2860, U.S. Atomic Energy Commission, Technical Information Division, Oak Ridge, TN, USA, October 28, 1944. 60 pp. URL http://

hdl.handle.net/2027/mdp.39015086444083. Notes written by J.

O. Hirschfelder. Declassified 22 April 1949. This report is based on LADC-549.

Peierls:2004:FOR

[PR04] Rudolf Peierls and Xavier Roqué. Frisch, Otto Robert (1904–1979), physicist. In Oxford Dictionary of National Biography, page ?? Oxford University Press, Walton Street, Oxford OX2 6DP, UK,

2004. URL https://www.oxforddnb.com/search?q=Frisch%2C+
Otto+Robert+%281904%E2%80%931979%29%2C+physicist.

Peierls:1949:TMM

[PS49] R. E. Peierls and T. H. R. Skyrme. Tank model for magnetic problems of axial symmetry. Philosophical Magazine Series 7 (1926– 1955), 40:269–273, 1949. CODEN PHMAA4. ISSN 0031-8086.

Peierls:1955:SFTa

[PSMF55] R. E. Peierls, A. Salam, P. T. Matthews, and G. Feldman. A survey on field theory. Reports on Progress in Physics, 18:423–477, January 1955. CODEN RPPHAG. ISSN 0034-4885 (print), 1361-

6633 (electronic). URL https://ui.adsabs.harvard.edu/#abs/
1955RPPh...18..423P.

Peierls:1958:PVS

[PSP^{+}58] R. E. Peierls, A. Salam, D. L. Pursey, P. K. Kabir, D. Sciama, T. W. B. Kibble, J. S. Bell, and Y. Yamaguchi. Parity violation in strange-particle decays: Discussion. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 246: 487–490, August 1958. CODEN PRLAAZ. ISSN 0080-4630 (print),

2053-9169 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1958RSPSA.246..487P.

Peierls:1952:PTO

[PSW52] R. E. Peierls, K. S. Singwi, and D. Wroe. The polyneutron theory of the origin of the elements. Physical Review (2), 87(1):46–50, July 1952. CODEN PHRVAO. ISSN 0031-899x (print), 1536-6065 (elec-

tronic). URL https://link.aps.org/doi/10.1103/PhysRev.87.
46;   https://ui.adsabs.harvard.edu/#abs/1952PhRv...87.
..46P.

Peierls:1997:PTO

[PSW97] R. E. Peierls, K. S. Singwi, and D. Wroe. The polyneutron theory of the origin of the elements. In Dalitz and Peierls [DP97a], pages 383– 387. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://

ui.adsabs.harvard.edu/#abs/1997sspr.book..383P.

Peierls:1962:VAC

[PT62] R. E. Peierls and D. J. Thouless. Variational approach to collective motion. Nuclear Physics, 38:154–176, October/November 1962. CODEN NUPHA7. ISSN 0029-5582 (print), 1873-1546 (electronic).

URL https://ui.adsabs.harvard.edu/#abs/1962NucPh..38..
154P.

Peierls:1997:VAC

[PT97] R. E. Peierls and D. J. Thouless. Variational approach to collective motion. In Dalitz and Peierls [DP97a], pages 530–552. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.adsabs.

harvard.edu/#abs/1997sspr.book..530P.

Peierls:1968:VFNa

[PU68] R. E. Peierls and J. N. Urbano. The Villars formalism for nuclear rotation. Journal of Physics A (Mathematical and General), 1(1):1– 10, January 1968. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-

6447 (electronic). URL http://stacks.iop.org/0022-3689/1/i=
1/a=301;   https://ui.adsabs.harvard.edu/#abs/1968JPhA.
...1....1P.

Peierls:1997:VFN

[PU97] R. E. Peierls and J. N. Urbano. The Villars formalism for nuclear rotation. In Dalitz and Peierls [DP97a], pages 553–562. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.adsabs.

harvard.edu/#abs/1997sspr.book..553P.

Peierls:1957:CMN

[PY57] R. E. Peierls and Jean Yoccoz. The collective model of nuclear motion. Proceedings of the Physical Society, London, Section A, 70(5): 381–387, May 1957. CODEN PPSAAM. ISSN 0370-1298 (print),

1747-3829 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1957PPSA...70..381P.

Peierls:1997:CMN

[PY97] R. E. Peierls and J. Yoccoz. The collective model of nuclear motion. In Dalitz and Peierls [DP97a], pages 441–448. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.adsabs.

harvard.edu/#abs/1997sspr.book..441P.

Reed:2022:CPF

[Ree22] B. Cameron Reed. Comments on the physics of the Frisch–Peierls memorandum. Nuclear Technology, 208(12):1890–1893, August 2022. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Reed:2024:HB

[Ree24a] B. Cameron Reed. On Heisenberg and the bomb. American Journal of Physics, 92(12):908–909, December 2024. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). This short paper corrects some errors in [McC24] and supplies additional historical material.

Reed:2024:RFP

[Ree24b] B. Cameron Reed. Revisiting the Frisch–Peierls memorandum. European Physical Journal H, 49(1):??, December 2024. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-024-00070-x.

Reynolds:2003:HAS

[Rey03] Paul Reynolds. How atom spy slipped security net. BBC News

Web story., May 21, 2003. URL http://news.bbc.co.uk/1/hi/
uk/3046255.stm.

Roberts:1956:BRR

[Rob56] Louis D. Roberts. Book review: R. E. Peierls, Quantum Theory of Solids. Physics Today, 9(5):29–??, May 1956. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/abs/10.1063/1.3059963;
https://ui.adsabs.harvard.edu/#abs/1956PhT.....9e..29P.

Salpeter:1992:BRR

[Sal92] Edwin E. Salpeter. Book review: Rudolf Peierls, More Surprises in Theoretical Physics. Physics Today, 45(8):68, August 1, 1992. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL https://physicstoday.scitation.org/doi/
abs/10.1063/1.2809773;   https://ui.adsabs.harvard.edu/
#abs/1992PhT....45h..68P.

Salpeter:2002:GLB

[Sal02] Edwin E. Salpeter. A generalist looks back. Annual Review of Astronomy and Astrophysics, 40(1):1–25, September 2002. CODEN ARAAAJ. ISSN 0066-4146 (print), 1545-4282 (electronic).

Schick:2017:HNP

[Sch17] Michael Schick. Historical note on 2016 Physics Nobel. Physics Today, 70(8):13, August 2017. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Segre:1985:HPR

[Seg85] Emilio Segrè. Historical perspective: Refugee scientists and nuclear energy. Annals of the New York Academy of Sciences, 452(1):xv– xix, 1985. CODEN ANYAA9. ISBN 0-89766-298-9, 0-89766-299-7 (paperback). ISSN 0077-8923 (print), 1749-6632 (electronic). Sixth International Conference on Collective Phenomena: reports from the Moscow Refusnik Seminar / edited by Inga Fischer-Hjalmars and Joel L. Lebowitz. Contributions from the Moscow Refusnik Seminar and from two International Conferences on Collective Phenomena, one held in Stockholm, Sweden, 1–2 December 1983, and the other in Tel Aviv, Israel, 31 May–1 June 1984.

Segre:1993:MAM

[Seg93] Emilio Segrè. A mind always in motion: the autobiography of Emilio Segrè. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1993. ISBN 0-520-07627-3. xii + 332 pp. LCCN QC16.S35 A3 1993. URL http://ark.cdlib.

org/ark:/13030/ft700007rb;   http://www.loc.gov/catdir/
description/ucal041/92010722.html.

Shapiro:1950:BRA

[Sha50] Maurice M. Shapiro. Book review: Atomic Energy, edited by R. E. Peierls and J. L. Crammer. Bulletin of the Atomic Scientists, 6(11): 347, November 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shoenberg:1939:MPB

[Sho39] D. Shoenberg. The magnetic properties of bismuth III. Further measurements on the de Haas–van Alphen effect. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, A170(942):341–364, April 3, 1939. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Stuewer:1979:NPR

[Stu79] Roger H. Stuewer, editor. Nuclear physics in retrospect: proceedings of a symposium on the 1930s. University of Minnesota Press, Minneapolis, MN, USA, 1979. ISBN 0-8166-0869-5. LCCN QC773 .S95 1977.

Stuewer:2018:AIN

[Stu18] Roger H. Stuewer. The Age of Innocence: Nuclear Physics Between the First and Second World Wars. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-186658-X, 0-19-882787-3 (hardback), 0-19-256290-8 (e-book). xv + 484 pp. LCCN QC773 .S78 2018.

Salam:1972:AQT

[SW72] Abdus Salam and Eugene Paul Wigner, editors. Aspects of quantum theory. Cambridge University Press, Cambridge, UK, 1972. ISBN 0-521-08600-0. xvi + 268 pp. LCCN

QC174.1 .A85 1972. URL http://hooke.lib.cam.ac.uk/cgi-
bin/bib_seek.cgi?cat=ul&bib=1733506;   http://www.loc.
gov/catdir/enhancements/fy1001/72075298-d.html;   http:
//www.loc.gov/catdir/enhancements/fy1001/72075298-t.html.

Taylor:1987:TPD

[Tay87] John Gerald Taylor, editor. Tributes to Paul Dirac. Adam Hilger Ltd., Bristol, UK, 1987. ISBN 0-85274-480-3. LCCN QC16.D57 T75

1987. URL http://www.loc.gov/catdir/enhancements/fy0745/

87153334-d.html. Based on the papers presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

terHaar:1965:CPD

[tH65] Dirk ter Haar, editor. Collected papers of L. D. Landau. Pergamon Press, New York, NY, USA, 1965. xx + 836 pp. LCCN QC16.L25 A2 1965.

Theaker:2012:BRE

[The12] Martin Theaker. Book review: Elemental Germans: Klaus Fuchs, Rudolf Peierls and the Making of British Nuclear Culture 1939–59, Christoph Laucht, Basingstoke, Palgrave Macmillan, 2012, xiv + 274 pp., ISBN 978-0-230-35487-6 (hbk) (£55.00). Contemporary British History, 26(4):553–555, December 2012. ISSN 1361-9462 (print), 1743-7997 (electronic).

Thouless:1981:BRS

[Tho81] David Thouless. Book review: Surprises in Theoretical Physics, by Rudolf Peierls. American Journal of Physics, 49:382, April 1981. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL https://ui.adsabs.harvard.edu/#abs/1981AmJPh..49..
382P.

Thomas:1995:RPA

[Tho95] Robert McG. Thomas, Jr. Rudolf Peierls, 88, atomic physicist, dies in England. New York Times, ??(??):??, September 22, 1995. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-

667X, 1553-8095. URL https://www.nytimes.com/1995/09/22/
obituaries/rudolph-peierls-88-atomic-physicist-dies-in-
england.html.

Thouless:1999:BRT

[Tho99] David Thouless, F.R.S. Book review: Two pioneers of modern theoretical physics: H. A. Bethe, Selected works of Hans A. Bethe, with commentary. World Scientific Publishing Company, 1997. Pp. viii + 605, £47.00 (hardback). ISBN 981-02-2876-7. R. H. Dalitz and Sir Rudolf Peierls, Selected scientific papers of Sir Rudolf Peierls, with commentary. World Scientific Publishing Company and Imperial College Press, 1997. Pp. xxiii + 805, £67.00 (hardback). ISBN 981-02-2692-6. Notes and Records of the Royal Society of London, 53(2):282–288, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532214.

Turner:1940:NF

[Tur40] Louis A. Turner. Nuclear fission. Reviews of Modern Physics, 12(1):1–29, January 1940. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.12.1;   http://rmp.
aps.org/abstract/RMP/v12/i1/p1_1. This paper surveys the

early work from 1934 to 1939 on nuclear fission, including more than 100 from 1939 alone following the news of the Hahn and Strassmann work that Niels Bohr brought to the USA in January 1939.

Turner:1956:BRR

[Tur56] Joseph Turner. Book reviews: R. E. Peierls, The Laws of Nature. The Scientific Monthly, 83(6):310, December 1956. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (elec-

tronic). URL http://www.jstor.org/stable/21669;   https:
//ui.adsabs.harvard.edu/#abs/1956SciMo..83..310P.

Uhlenbeck:1959:BRS

[Uhl59] George E. Uhlenbeck. Book review: Statistical Physics, Volume 5 of Course of Theoretical Physics. L. D. Landau and E. M. Lifshitz. Translated from the Russian by E. Peierls and R. F. Peierls. Pp. 484. Pergamon Press, Inc., London. Addison-Wesley Publishing Company, Inc., Reading, 1958. American Journal of Physics, 27 (5):371–372, May 1959. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Walker:2015:BRE

[Wal15] Mark Walker. Book review: Émigré Biographies: Anja Skaar Jacobsen, Léon Rosenfeld: Physics, Philosophy, and Politics in the Twentieth Century. Hackinsack: World Scientific, 2012. xii + 354 pp., illus., index. ISBN 978-981-4307-81-9. $87.00 (hardcover).; Christoph Laucht, Elemental Germans: Klaus Fuchs, Rudolf Peierls and the Making of British Nuclear Culture 1939– 1959. London: Palgrave Macmillan, 2012. xiv + 274 pp., index. ISBN 978-0-230-35487-6. $75.00 (hardcover).; Mary Jo Nye, Michael Polanyi and His Generation: Origins of the Social Construction of Science. Chicago: University of Chicago Press, 2011. xi + 405 pp., illus., index. ISBN 978-0-226-61063-4. $50.00 (hardcover).; Silvan S. Schweber, Nuclear Forces: The Making of the Physicist Hans Bethe. Cambridge, MA: Harvard University Press, 2012. viii + 579 pp., illus., index. ISBN 978-0-674-06587-1. $35.00 (hardcover). Historical Studies in the Natural Sciences, 45(1): 206–216, February 2015. CODEN ???? ISSN 1939-1811 (print),

1939-182X (electronic). URL http://www.jstor.org/stable/10.
1525/hsns.2015.45.1.206.

Weisskopf:1981:BRS

[Wei81] Victor F. Weisskopf. Book review: Surprises in Theoretical Physics, by Rudolf Peierls. Physics Today, 34(8):64, August 1981. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL https://physicstoday.scitation.org/doi/
abs/10.1063/1.2914702;   https://ui.adsabs.harvard.edu/
#abs/1981PhT....34h..64P.

Weisskopf:1986:BRR

[Wei86] Victor F. Weisskopf. Book review: Rudolf Peierls, Bird of Passage — Recollections of a Physicist. American Journal of Physics, 54:856–857, September 1986. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1986AmJPh..54..856P.

Wells:1992:LHB

[Wel92] Alan Wells. Letters: Heisenberg’s bomb. New Scientist, 135(1842): 50, October 10, 1992. CODEN NWSCAL. ISSN 0262-4079 (print),

1364-8500 (electronic). URL http://www.newscientist.com/
article/mg13618426.200-letters-heisenbergs-bomb.html. See

[Cha92].

Whitley:1984:RGCa

[Whi84a] Stanley Whitley. Review of the gas centrifuge until 1962. Part I: Principles of separation physics. Reviews of Modern Physics, 56 (1):41–66, January 1984. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.56.41;   http://rmp.aps.
org/abstract/RMP/v56/i1/p41_1.

Whitley:1984:RGCb

[Whi84b] Stanley Whitley. Review of the gas centrifuge until 1962. Part II: Principles of high-speed rotation. Reviews of Modern Physics, 56(1):67–97, January 1984. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.56.67;   http://rmp.aps.
org/abstract/RMP/v56/i1/p67_1.

Whitaker:1996:EBQ

[Whi96] Andrew Whitaker. Einstein, Bohr, and the quantum dilemma. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-48220-8 (hardcover), 0-521-48428-6 (paperback). xvii + 349 pp. LCCN QC174.12 .W48 1996. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.cambridge.
org/uk/catalogue/catalogue.asp?isbn=0521484286;   http:
//www.loc.gov/catdir/description/cam027/95018270.html;
 http://www.loc.gov/catdir/toc/cam021/95018270.html.

Wheater:1983:CPD

[WP83] John Wheater and Rudolf Peierls. Comment on “Is Proton Decay Measurable?”. Physical Review Letters, 51(17):1601, October 24, 1983. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114

(electronic), 1092-0145. URL https://ui.adsabs.harvard.edu/
#abs/1983PhRvL..51.1601W.

Weart:1985:HP

[WP85] Spencer R. Weart and Melba Phillips, editors. History of Physics, volume 2 of Readings from Physics Today. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-468-0 (paperback). x + 375 pp. LCCN QC7 .H694 1985.

Yamaguchi:1958:PMD

[YBL^{+}58] Y. Yamaguchi, G. H. Burkhardt, A. Lundby, J. M. Cassels, P. E. Cavanagh, and R. E. Peierls. Pion and muon decay: Discussion.

Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 246:465–466, August 1958. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL

https://ui.adsabs.harvard.edu/#abs/1958RSPSA.246..465Y.

Zhang:2019:PAP

[ZLZ19] S. H. Zhang, D. Legut, and R. F. Zhang. PNADIS: an automated Peierls–Nabarro analyzer for dislocation core structure and slip resistance. Computer Physics Communications, 240(??): 60–73, July 2019. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465519300839.